A Complete Bibliography of Computers and Security, 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

23 September 2019 Version 1.00

Title word cross-reference

$12.50 [Ano83j]. $12.95 [Hig83e]. $13.95 [Hig83d]. $14.95 [Ano83g, Ano84h, Ano83h]. $15.00 [Hig83f]. $15.95 [Hig82c]. $16.95 [Ano84j, RFB82b]. $18.95 [HJH85c]. $19.50 [Woo85a]. $19.95 [Ano83g, HJH85b]. $20.00 [Hig83f]. $22.95 [Ano82d]. $26.95 [Ano84i]. $28.75 [EHH85a]. $29.50 [Hig83b]. $29.95 [Ano84h, Hig82b, RFB82a]. $34.95 [Ano82e]. $39.95 [Ano84f]. $46.25 [HJH85a]. $500 [Bar88]. $7.50 [Hig82a]. x [Hig83e].

1 [Hig88q, JO89a, Ano84e]. 1-2-3 [Ano88a]. 101/1027 [EHH85b]. 11 [Ano83s, Ano84-32, Ano88-66]. 112 [Hig88-89]. 12 [Paa84b]. 120.00/US [HJH85a]. 121 [Hig88-42]. 13 [Hig88-94]. 14 [Hig88-121, Paa87a]. 16 [Hig88-55]. 17 [Hig88-103, Hig88-112]. 170 [Hig88-135]. 18 [Hig88-73]. 1974 [Ano88-100]. 1975 [Ano88x, Ano88-27, Ano88-43, Ano88-98]. 1978-1986 [Ano88-50]. 1982 [Ano82n, Ano82p, Ano82o]. 1983 [Ano83u, Ano83w, Ano83v]. 1984 [Ano84z, Ano84y, Ano84-28, Ano84-27]. 1985 [Ano85t, Ano85r, Ano85q, Ano85s]. 1986 [Ano86-35, Ano86-36, Ano86-33, Ano86-34, Ano88v, Ano88t, Ano88x, Ano88-39, Ano88-40, Ano88-94, Ano88-142, Ano88-144]. 1986-September [Ano88u]. 1987 [Ano87-30, Ano87-27, Ano87-29, Ano87y, Ano87z, Ano87-28, Ano88-38, Ano88-37, Ano88w, Ano88u, Ano88y, Ano88z, Ano88-27, Ano88-30, Ano88-31, Ano88-33, Ano88-41, Ano88-43, Ano88-53, Ano88-54, Ano88-63, Ano88-100, Ano88-98, Ano88-141, Ano88-143]. 1988 [Ano88-28, Ano88-99, Ano88-116, Ano88-118, Ano88-120, Ano88-115, Ano88-117, Ano88-119]. 1989 [Ano89-38, Ano89-34, Ano89-37, Ano89-32, Ano89-33, Ano89-31, Ano89-36, Ano89-35].

2 [JO89b, Ano84f]. 2.7 [Kug86]. 21 [Hig88-78]. 26 [Hig88m]. 29 [Hig88-97].

3 [Hig88-63]. 30 [Hig88-53]. 31 [Hig88l, Hig88-137]. 33 [Hig88-93]. 35 [Hig88u, Hig88-48, Hig88-92]. 36 [Hig88k]. 37 [Hig88-68]. 38 [Ano88-97]. 39 [Hig88-82].

4 [Fåk87b, Hig88-62, Hig88-98]. 40 [Hig88-100]. 43 [Hig88-30]. 44 [Hig88-79]. 48 [Hig88-69, Hig88-115]. 497 [Hig88-122].

5 [Hig88-71, Hig88-95]. 512 [Hig88-52]. 53 [Hig88-50, Hig88-113]. 54 [Hig88-35, Hig88-76].

6 [Hig88-99, Hig88-120].

72 [Hig88-105]. 75.00/US [EHH85a]. 776-101 [EHH85b]. 776-101/1027 [EHH85b]. 78 [Hig88-101]. 7th [Ano85a].

8 [Hig88o, Hig88-64]. 81 [Hig88-96]. ’83 [Ano83p, Ano83q, Hig88-126, Ano83i]. ’84 [Ano84p, Ano85n, Woo84, Hig88-58, Hig88-114]. ’85 [Ano85w, Ano86n]. ’86 [Ano86u]. ’88 [Hig88-162, Hig88-72, Hig88-118]. 89 [Hig88-41].

9 [Hig88-104]. 92 [Hig88-66]. 93 [Hig88y, Hig88-124].

A1 [Hig88-91, LDS^{+}88]. A19 [Hig88-91]. Abstract [Ano85b, Hug88, Ano89-39]. Abstracts [Ano82a, Ano82b, Ano82i, Ano83a, Ano83b, Ano83c, Ano84a, Ano84b, Ano84c, Ano85c, Ano85d, Ano85e, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87e, Ano87c, Ano87f, Ano87d, Ano89a, Ano89b, Hig88j, Hig89a, Hig89b, Hig89c, Hig89d, Hig89e, Hig89f]. abuse [Hig87c, Mou82, Woo82b]. Access [Hov83, Lun89, Ams88, Ano86y, Ano87-40, Ano88-95, Boz88b, EHH85a, Hig88-109, Hig88-118, Hig88-131, Kar86, Men89e, San88, Yeo85, dC85]. accountant [Har86]. accounting [GM85]. accurate [Har88]. achieve [Woo89]. ACM [Hig88-52, Hig88-122, LS87]. acquisition [All88]. acquisitions [Ano88-134]. act [Hig88-75, Hig88-110, Mil88d]. ACTEN [FM84]. activities [Hov83, TR86a, TR86b]. actors [Hof89b]. Adaptable [GPR^{+}88a, GPR^{+}88b]. adaptive [WC88]. added [Hig86b]. address [Ano88f, Men88e]. Addresses [Ano88b, Ano88c, Owe86]. addressing [BvS89]. adds [Ano88-55]. administration [Coh87a, Coh87d]. administrator [Hig89w]. ADP [SK82]. Advanced [Ano86e, vdBL84]. Advances [Ano84e, Ano84f, Hig83b]. advice [Hig88-112]. advocates [Hig88-78]. affordable [Hig88-126]. afforded [WC88]. After [CM88, Mil85]. Aftershocks [Ano88d]. against [ALN88, Coh89b, Hig88-126, Sop88]. Age [Hig88y, Hig88-101, Hig88-105, Hig88-118, Hig88-114, Ano88-81, Ano88-82, Hof89b, Hig82c]. Agencies [Hig88k, Hig88l, Hig88-35, Hig88-75, Men88e]. agency [Ano84j]. ahead [PH87]. Ahituv [Ano88-123]. aid [CMRW88]. aided [HJH85c]. aids [Hig88-167]. aimed [Ano88-130]. Alan [Ano82e, Hig88-101, Hig88-104, Ano88-104, Hig88-137]. alert [Ano88-150]. Alex [Ano88-35]. algebraic [She88]. algorithm [Ben88, Hen82, HC88]. algorithms [SB84, She88]. Allen [Hig83e]. Alper [Ano88-104]. alternatives [LS87]. Altman [Ano88-103]. America [Ano84j]. Amsterdam [Ano83i, EHH85a, HJH85a]. Analog [Jay82]. Analysis [Ben88, TTA^{+}88, All88, Ano86-41, BR83, CW83, Car83, Gua87, Hof89a, Hof89b, Lon87, NCCW88, Sto88]. analysts [Ano82d]. analyzing [CI85]. Anatomy [Ano88e, Hig88-143, Hig88-142]. and/or [Ano86y]. Anderson [Ano88-61]. Andrew [Hig88u]. Angel [Hig88-30]. Annotated [OR88, Kur82]. annoying [Per88]. ANSI [Ano88f]. anti [Hig88o, Hig89g]. anti-viral [Hig88o]. anti-virus [Hig89g]. antiviral [Par88b]. any [Fåk86]. Apollo [Ano88g]. Apple [Hig88-96, Ano88h, Gof88a, Hig88n, Hig88r, Hig88q, Hig88m, Hig88o, Hig88p, Hig88-43, Hig88-71, Jon88, Ric88]. Apple/Microsoft [Hig88q]. application [Mur88]. applications [Ano88-39, Ano88-40, Ano88-41, Ano88-53, SK88, TC88, Woo82a]. applied [Lon87]. Applying [Sch82]. appreciate [Hig88k]. approach [Ala86, Ano88-102, Bac83, BM83, Cou82, Gua87, Har89, Hig88-27, Hig88-117, Hug88, Lun88, Mil88b, Mou82, Mou84b, PP88, PG87, Woo83b]. Approaches [LAN86]. approved [Ano88-29]. approves [Ano88-139]. April [Ano88y, Ano88-28, Hig88k, Hig88y, Hig88-30, Hig88-35, Hig88-41, Hig88-46, Hig88-55, Hig88-58, Hig88-72, Hig88-69, Hig88-63, Hig88-64, Hig88-73, Hig88-76, Hig88-79, Hig88-82, Hig88-91, Hig88-93, Hig88-94, Hig88-96, Hig88-97, Hig88-100, Hig88-89, Hig88-101, Hig88-106, Hig88-105, Hig88-118, Hig88-120, Hig88-112, Hig88-114, Hig88-115, Hig88-130, Hig88-129, Ano87-30, Ano88-116, Ano89-38, Hig88-47]. architectural [Bon83]. architecture [Ram89a, Ram89b]. area [Ano88-72, Ano88-94, Har86, Hig88-102, JL89, Sch82, TC88]. arm [Hig88-60]. Arnett [Hig88o, Hig88-63, Hig88-64, Hig88-120]. arriving [Per88]. art [Lob83]. Artech [Hig83b]. Articles [Ano85e, Ano82a, Ano82b, Ano82i, Ano83a, Ano83b, Ano83c, Ano84a, Ano84b, Ano84c, Ano85c, Ano85d, Ano85b, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87e, Ano87c, Ano87f, Ano87d, Ano89a, Ano89b, Hig88j, Hig89a, Hig89b, Hig89c, Hig89d, Hig89e, Hig89f, Hug88]. aspects [Ano88-33, Buu84, Coh89a, Hig88-51, LK84, Paa84b]. Assessment [Ano83t, PES88, Ano87-33, For82, Gar89a, PHEW88]. assets [FM83, Men89c]. Association [Hig83c, Paa84b, Smi87]. assumes [Ano88-102]. assurance [BPR88, LDS^{+}88, TR86a, TR86b]. attack [Fåk88, Hig88-143, Hig88-146, Hig88-163, Hig88-130, Hig88-144, Whi89, vdAvE86]. attacks [Hig88-151, Hig88-41, Hig88-80, Sto88]. attitude [Hig88-121]. attitudes [HR88]. Audit [Ano87-34, KD86, Paa87b, Ano86z, Bea86a, Bea86b, Boz88b, Dos86, Koc83, Paa88, Top84, Web88, vBH88, vdBL84]. Auditing [Koc84, Ano88-109, Hig88-99, Paa84a, Ten86a]. auditor [Bol89b, Men85a]. Auditors [Mou87, Hig88-132]. augmented [Thu87]. August [Ano88-30, Ano88-31, Ano88-63, Ano88-141, Ano84z, Ano87-27, Ano88-118, Ano89-34]. Authenticating [Smi87]. Authentication [Kar86, PvL86, PW86, PW87b, ALN87, Ano86-28, Ano87u, Ano87-41, Ano88-126, Chr88, MRW89, Spe87]. authenticators [FVTS87]. Author [Ano82c, Ano83d, Ano84d, Ano85f, Ano86g, Ano87g, Ano88i, Ano89c]. authorities [Hig87-32]. authority [Boz88b]. authorization [Plu88]. Autodesk [Ano88-67]. automated [CR88]. automatic [Hig88-169]. Automating [FVTS87]. automation [Hig85d, Hig87w, Hig88-167, Kee89, Sch84, Won88a]. averts [Hig88-98]. awareness [CMRW88].

B [Ano84i, Ano88-122]. Back [Ano86h]. backed [Hig88-110]. backfire [Ric88]. backup [Ano87r, Hig88v]. bad [Hof89b]. Ball [Ano88-148]. Bamford [Ano84j]. band [Bur88]. bandwagon [Par88b]. bank [Woo86b]. banking [Sei84]. Barbara [Hig88-76, Hig88-115]. bark [Hig88-50]. Barney [Hig88-95, Ano88a]. Barton [Ano88-106]. base [CJ86, EHH85a]. based [BS82, Car82, Hig88-109, Sal88, SK88]. baselines [PHEW88]. basic [Hig88-67, vdAvE86]. basics [Ano88-132]. BASIS [Ano88-52]. Bass [Hig88-58]. Bates [Hig88-76, Hig88-115]. Battling [Ano88j]. be [Hig88-50, Zaj88a]. Beatson [Ano88-51]. Becker [Ano83e, Ano84i]. Bedford [Ano83f]. been [Mil85, Zaj86]. before [Ano88k]. behind [Hig88-50]. Beijing [Hig87-27]. being [Hig85d]. belatedly [Men88e]. Belden [Hig88k]. Belgian [vH87]. belief [TTA^{+}88]. believe [Zaj86]. Bell [Men85b]. Benutzer [MB88]. Berkowitz [Ano88-45]. Bernard [Ano88k]. Betts [Ano88f]. Beyond [Hig88-90]. bibliography [Hig83c, Kur82, OR88]. big [Hig88-90]. Bill [Gri88a, Hig88-110, Hig88-83, Hig88-45]. biometric [Hig88-83]. biometrics [Hig88-139]. bits [Ano88-125, Hig88-168, Hig89n, Hig89o, Hig89p, Hig89q, Hig89r, Hig89s, Hig89t, Hig89u]. Blitzed [Hig88a]. blunders [Hig88-116]. blurs [Lew88a]. Board [Ano83r, Ano83-27, Ano84x, Ano85o, Ano87v, Ess86, Hig87v, Ano82l, Ano82m, Ano85p, Ano86-29, Ano88-83, Ano88-84, Ano88-85, Ano88-86, Ano88-87, Ano88-88, Ano89v, Ano89w, Ano89x, Ano89y, Ano89z, Ano89-27, Ano89-28, Ano89-29, Ano82j, Ano83o, Ano84r, Ano85l, Ano86r, Ano87o, Ano88-49, Ano89l]. Bond [Hig88-98, Sha89]. Book [Ano82e, Ano82d, Ano83e, Ano83f, Ano83h, Ano83g, Ano83i, Ano83j, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano88k, Ano88l, EHH85a, HJH85a, HJH85b, HJH85c, Hig82c, Hig82b, Hig83b, Hig83c, Hig83d, Hig83e, Hig83f, RFB82a, RFB82b, Woo85a, Hig88-88, Hig88i]. Books [Hig83e]. boost [Hig88-83]. Borland [Ano88-101]. bosse [Bol89a]. Boston [Ano84j, HJH85c, Hig82c]. Bosworth [Hig83d]. both [Hig88-126, Sop88]. bottom [Ano88-59]. Bowden [Hig88-105]. Box [Hig83b, Hig88-90]. Bozman [Hig88-42]. Brad [Hig88-58]. brain [Hig88-145]. Brazil [Hig88s]. breach [Hig88-96, Hig88-91]. breached [Ano88-105]. break [Ano86m]. Breaking [Hig83e]. Breakpoints [Men88a]. breakup [Men85b]. Brian [Hig88-79]. brief [Ano86v]. briefly [Kur82]. briefs [Sch88]. British [Hun88a]. broad [Hig88-66]. Brookfield [Ano83j]. Brown [Ano88d, Ano88-95]. Bruce [Hig83d]. bug [Hig88-73]. bugs [Per88]. build [Hig88-113]. Bureau [Ano84g, Hig87d, LN88, Ano83n]. burglar [vH84]. Business [Hig83c, Hig88-134, KG86, KG87, dC85, HJH85a]. businesses [Ano88-110]. Butterworth [Hig82c]. Butterworths [HJH85c]. Buyers [Ano88m]. byte [Hig88-50]. bytes [Ano88-125, Hig88-168, Hig89n, Hig89o, Hig89p, Hig89q, Hig89r, Hig89s, Hig89t, Hig89u].

C [Ano88-45, Ano88-55, Ano88-72, Ano88-107]. cable [Ano88-132]. cable/video [Ano88-132]. Calamities [Hig88t]. Calendar [Ano82f, Ano82g, Ano82h, Ano83k, Ano83l, Ano83m, Ano84k, Ano84l, Ano84m, Ano84n, Ano85g, Ano85h, Ano85i, Ano85j, Ano86i, Ano86j, Ano86k, Ano86l, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, Ano88n, Ano88o, Ano88p, Ano88q, Ano88r, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k]. callback [Bis86]. called [Hig88-62]. calls [Mil88b, PH86]. campaign [Hig88l]. campus [Whi89]. Can [Hig88u, Hig88-139, Zaj88a, Hig88-30, Hig88-141, Ric88]. Canada [Ano83i, Bar88]. cancer [Hig88w]. can’t [Hig88-55]. capabilities [Ano86-32]. capability [SB84]. card [Ano88-73]. cards [Ano88-141, Hig88-118]. Carnahan [Ano84v]. Carole [Ano88m]. Carroll [HJH85c]. cartridge [Hig87g]. Case [Ano88-101, Hig88-146, Coh87e, Gla88, Hig88-39]. Cases [HP84, Bum84]. catch [Ano88-148]. catching [Hig88-107]. categories [San88]. causes [HP84]. CBEMA [Hig83c]. Center [Hig88-47, Ano88-47, vdBL84]. Central [Kim88]. certifiably [Bon83]. certification [PP88]. Cerullo [Hig88-101]. challenge [Hov86, Paa87c, Ano83e]. challenges [Ala86]. charade [Hig87a]. charges [Hig88-96]. checking [Thu87]. checklist [Ano88-133]. checklists [Hov83]. checksum [Coh87c, HC88]. Chernobyl [Hig88-147]. Chicago [Hig88-42]. chief [Raa88]. chill [Hig88-134]. China [Ano86-30]. chip [She88]. chosen [vdAvE86]. chosen-plaintext [vdAvE86]. Christine [Hig88-100]. Christmas [Hig88-149]. Christopher [Ano88-140]. Cipher [Hig87b, Hen82, Hig88-89]. ciphers [CR88, Hig83d]. Citations [Ano88x, Ano88y, Ano88v, Ano88-38, Ano88w, Ano88-27, Ano88-28, Ano88-44, Ano88-43, Ano88-50, Ano88-94, Ano88-100, Ano88-98, Ano88-99, Ano88-142, Ano88-143, Ano88-144]. city [Hig82a]. Clarke [Hig88-52]. classification [FG89, LDS^{+}88]. classified [Ano88-105]. Clifford [Hig88-122]. clouds [Ano88-104]. Co [Ano83j, Ano84i, Hig83e, RFB82b]. Coalition [Ano88s]. code [Ser85]. codes [Ano88-124, HJH85b, Hig83e, Hig83d]. Coffee [Ano86m]. collar [Bol88d]. colleges [Bri88]. colour [DR88]. combat [Ano88s, Hig88-159]. combined [Chr88]. come [Hig88-33]. comfort [Her87]. coming [Ano88-114]. command [Kug86]. comments [Ano87q]. commercial [Bra88]. Commission [Ano84g]. Committee [Ano88-78, Ano89n, Ano85m, Ano88-74, Ano88-75, Ano88-76, Ano88-77, Ano88-79, Ano89o, Ano89p, Ano89q, Ano89r, Ano89s, Ano89t, Ano89u]. Commonwealth [Ano89-39]. communication [Ano88-50, Rih87, Top84, Zaj88a, HJH85a]. Communications [Ano88-44, Hig88-52, Hig88-122, Hig88-126, Ano86-27, Ano87-37, Ano88t, Ano88u, Ano88-64, Bur88, Cer83, Hig88-147, Hig88-158, Hig88-108, Woo89]. Communities [Ano88v, Ano88-38, Ano88w, Ano88x, Ano88-27, Ano88-43, Ano88-50, Ano88-100, Ano88-98, Ano88-99, Ano88-142, Ano88-143, Ano88-144]. community [Sei84]. companies [Bri88]. Company [Ano83f, Ano83g, Ano84j, HJH85b, Hig83d, Ano88-130]. compels [Hig88-75]. complete [Ayo83]. complexity [Gag88b]. components [WZ86]. compression [Ano86q, Ano88-39, Ano88-40, Ano88-41, WC88]. compromise [Men88c]. compromised [Zaj86]. COMPSEC [Hig88d]. computation [FVTS87]. Computational [Coh89a]. Computer [Ano83n, Ano84o, Ano87e, Ano87f, Ano87-34, Ano88v, Ano88w, Ano88t, Ano88u, Ano88x, Ano88y, Ano88z, Ano88-27, Ano88-28, Ano88-30, Ano88-29, Ano88-31, Ano88-32, Ano88-33, Bri88, Coh87b, Com84, Elo88, For82, Gol83, Hig83c, Hig83f, Hig83a, Hig87c, Hig87e, Hig87d, Hig87f, Hig88k, Hig88l, Hig88v, Hig88w, Hig88x, Hig88z, Hig88y, Hig88-27, Hig88-28, Hig88-29, Hig88-30, Hig88-148, Hig88-31, Hig88-32, Hig88u, Hig88-35, Hig88-53, Hig88-58, Hig88-66, Hig88-72, Hig88-68, Hig88-69, Hig88-62, Hig88-76, Hig88-78, Hig88-79, Hig88-82, Hig88-103, Hig88-100, Hig88-104, Hig88-113, Hig88-130, Hoe86, Hun88a, Men89a, Mur83, Owe86, Paa84a, Paa84b, PH87, Paa87b, Saa87, Sol87, Sta86a, Won85, Ano82k, Ano82q, Ano83x, Ano83y, Ano84e, Ano84f, Ano84-29, Ano85a, Ano85u, Ano85v, Ano88-36, Ano88-60, Ano88-63, Ano88-91]. computer [Ano88-103, Ano88-114, Ano88-126, Ano88-128, Ano88-130, Ano88-145, Ano88-154, Ano88-156, Ano88-155, Ano89-39, BE89, BB84, BBFÅ82, Blo82, BS87, Car82, CI85, Cer85, Coh85, Coh88d, Coh88c, Coh89a, Coh89b, EHH85b, Elo85, FG89, Fif89, Fis84, Gle89, Guy83, HJH85c, HJH85d, Hig82b, HH82, Hig83b, Hig83i, Hig85e, Hig87a, Hig88a, Hig88-147, Hig88-33, Hig88-38, Hig88-41, Hig88-50, Hig88-159, Hig88-160, Hig88-92, Hig88d, Hig88-120, Hig88-112, Hig88-119, Hig88-171, HM86, Hof89a, HR88, Kar86, Kim88, LS87, Lob83, Men85a, Men89b, Mil88d, Mou82, Mur88, Neu82, Paa88, Per82, PR85b, PG87, RFB82a, RFB82b, Raa88, RB82, Ric86, RH86, Sch85, Spe87, Sta86b, Str88, SK88, War83, Web89b, Wis89, Woo82b, Woo85a, Zaj86, Zaj89b, bAG88, vdBL84, Ano82d, Ano83e, Ano88k, Ano88l]. computer-aided [HJH85c]. computer-based [Car82]. computer-mediated [BB84]. Computer-related [Men89a, Sol87]. computerized [Ano88-48]. computers [Ano84h, Ano87w, Ano88-54, Ano88-80, Ano88-109, Ano88-127, Hig83d, Hig87-30, Hig88u, Hun88c, KG86, Sto88, Hig88-91]. Computerworld [Hig88-42, Hig88-46, Hig88-55, Hig88-95, Hig88-98, Hig88-121, Hig88-137, Hig88-73]. computing [Har89, Har83, Hig88-107]. Concepts [Hig88-128, Kug86]. conceptual [FM83, FM84]. concern [Ano83e, Ano88-108, Ano88-136]. concerns [Hig88-102]. concise [PvL86]. Conference [Ano83q, Ano84v, Ano84-32, Ano87-34, Ano88-34, Paa87b, Ano83i, Ano84w, Ano85a, EHH85b, Fåk87b, Hig85e, Hig88x, Hig88d, Paa88, Paa84b]. conferences [Hig88h]. confidence [Ano88-46, Top84]. confidential [Hig88-81]. confidentiality [Ano85k, Cha85]. configuration [TC88]. configuring [CI85]. Confronting [Hig88-92]. confronts [Hig88w]. Congestion [Chi88]. Congress [Ano83e, Ano85v, Bec83, Hig88-171]. connection [Req88]. Connolly [Ano88g]. considerations [Ano88-134, Bra88]. CONST [Lun88]. constructive [Lun88]. consultants [Ano88-113]. consulting [Hig88-83]. containing [PG87]. context [BvS89, Woo88]. Contingency [Fri82, Sei84, Ano88-151, SK82]. Continuous [Koc84]. contract [Hig88-96]. Contractor [Ano88-81, Ano88-82]. Control [Ano87-34, Hig88u, Hig88-53, Hig88-68, Hig88-103, Hig88-104, Hov83, Paa87b, Ano82d, Ano86y, Ano87w, Ano87-40, Car82, Chi88, Dos86, Har86, Hig88-70, Hig88-78, Hig88-109, Hig88-131, Hig88-138, JL89, Kar86, Lun89, Men89e, MS89, Mou87, PH86, Paa88, Per82, San88, Ten86b, Yeo85, Ano83f]. controlled [OM84]. Controlling [Wei83, FG89]. Controls [Web89a, Ams88, Cer83, Cer85, GM85, Hig83a, Hig87-31, Mil85, Ten86a, Woo83a]. cool [Ano88-110]. copy [Hig88-39]. Copyright [Hig88-110, Lew88a, Ano88h, Hig88m, Hig88p, Hig88s, Hig88-43, Hig88-87, Par88a]. cornerstone [HR88]. corporate [Ano88-35, Hig88-137]. correct [Coh88a]. COSSAC [CI85]. Cost [BS87, Hig83a, Hig88-46, MP87]. Cost-effectiveness [BS87]. could [Ano88-59, Ano88-106, Gof88b, Hig88s, Hig88-136]. countermeasures [Com84, Hig88-130]. countersues [Gof88a]. countersuit [Hig88n]. Countries [Ano84g]. course [Ano88-36, Har83, Neu82]. court [Hig88-39]. courts [Bum84]. cracker [BBR87]. cracking [Hig87b]. cracks [Hig88-120]. Crash [Ano88-36]. crime [Ano83n, Ano88k, Ano88v, Ano88w, Blo82, Bol88d, Com84, Fif89, Gol83, HJH85d, Hig87-32, Hig88w, Saa87, Sta86a, Sta86b, Won85, bAG88]. criminal [Bol88d]. crisis [Bol88b]. critical [Ano88-55, Hig88-56]. Crockett [Ano88-106]. cross [Hig88x]. cryptananysis [CR88]. Crypto [Ano84p, Fåk87a, Ano86n, Fåk87b]. Cryptographic [Ser85, VA88, Ano86-32, Bur88, Coh87c, Erd86, HC88, SB84, Wei83]. cryptographically [Ayo83]. Cryptography [Ano88-38, Ano88-37, Fis84, GL82, Wil82a, BS82, Buc82, Gag88b, Wil82c, Woo82a, Ano82e]. Cryptologia [Hig88-89]. Cryptology [Por84, She88, Kah82, HJH85b]. cryptosystem [Jun87, Jun88, Sal88]. cryptosystems [Ano88-38, Ano88-37, Gag88a, Mul84, PR85a, She88]. Currah [Hig88-79]. Current [Ano82i, For82]. custom [She88]. cycle [BE89, Dos86, TR86a, TR86b].

D [Ano88-69, Ano88-137, Hig88-105, Woo85a]. D1 [Hig88-129]. D4 [Hig88-129]. damaging [Hig88-97]. Danca [Hig88-62]. Dance [Hig88-41]. Darryl [Ano88-34]. Data [All88, Ano85k, Ano88-39, Ano88-40, Ano88-41, Ano88-42, Ano88-43, Ano88-45, Ano88-100, Bur88, Cer83, Cha85, Eke86a, Gri88b, Hig85f, Hig87g, Hig87h, Hig87i, Hig88-94, Kra84, LK84, Mou84a, Sny88, Sop88, dC85, vdL86, Ala86, Ano84g, Ano84h, Ano84-30, Ano86q, Ano88-48, Ano88-61, Ano88-71, Ano88-113, Ano88-105, Ano88-123, Ano88-122, Ano88-130, Ano88-147, BB88, Blo82, Buu84, CW83, CJ86, Cou82, Cou88, DR88, EHH85a, FG89, FM83, Gof88b, Gri88a, Hig82a, Hig83f, Hig87-31, Hig87-32, Hig88k, Hig88l, Hig88-47, Hig88-85, Hig88-103, Hig88-97, Hig88-117, Hig88-123, Hig88-116, Hig88-130, Hig88-126, Hig88-141, Kur82, Men83, Men88b, Men89d, MS89, NCCW88, RFB82b, Ram89a, Ram89b, Top84, Woo89, Zaj88b, Zaj88c, Ano88-44, Ano84g]. data-base [EHH85a]. data-processing [CW83]. data-security [Ano88-113]. Database [Ano88x, Ano88-44, Ano88-43, Ano88v, Ano88-38, Ano88w, Ano88y, Ano88-27, Ano88-28, Ano88-50, Ano88-94, Ano88-100, Ano88-98, Ano88-99, Ano88-142, Ano88-143, Ano88-144, BM83, DJT87, GM85, Hig88-54, KTT89, MT88, Plu88, Thu87, vdL86]. databases [ALN88, Thu89]. Datacrime [Hig89g]. Datamation [Hig88m, Hig88-106]. dataveillance [Hig88-52]. David [Ano88-80, HJH85b, Ano88e, Ano88-151, Hig88-53, Hig88-113]. Davies [Hig83f, Hig88-53, Woo85a, Ano88e]. Davis [Ano88-80, Ano88-137]. day [Ano86y]. DB2 [Boz88a]. DBMS [TTT89]. DC [Ano83e, Hig83c]. DEA [Ano88-139]. dead [Ano88-89]. deadly [Hig88-119]. death [Hig87f]. debate [Hig88-61, Hig88-81]. DEC [Ano88-95, Hig88-33]. decade [MB88]. December [Ano88-38, Ano88-94, Ano88-142, Ano88-143, Ano85t, Ano86-35, Ano87-29, Ano88-120, Ano89-37]. Decentralized [Hig88-34]. Decision [Car83, Ten86a, BS87, Byl88, Fis84, Par88a, Sch85, Ten86b]. decisions [Ano84o]. decryption [Hen82]. Dedham [Hig83b]. default [Hig88v]. defence [Bra88]. defence-related [Bra88]. defense [Ano88-150, Coh88c]. defenses [Coh89b]. defensive [Bri88]. definition [Ano84i]. Delaware [Web89b]. delete [McD88]. Demands [Fåk87a]. demise [Ano86o].

Department [Ano83n, Gri88b]. Design [CJ86, Coh87d, Coh87e, PR85a, Voi88, Ayo83, Bon83, BM83, Coh85, Fåk87a, FM84, JL89, PR85b, TTT89]. designed [Bur88]. Designing [Coh88a]. Despite [Hig88-35]. destroys [Ano88-130]. detected [Hig88-58]. Detecting [She88]. Detection [Ten86b, Mal88]. deterioration [Ten86b]. deterring [Woo82b]. Developing [Per82]. Development [Ano88-46, Bea86a, Elo85, Kah82, TR86a, TR86b, Hig88-99]. developments [Ano88-154, Ano88-156, Ano88-155, vdG84]. devices [Ano87-41, Ano88-65, Mou84b, Spe87]. Dfl [Ano83i, EHH85a, HJH85a]. diagnosis [Mal88]. Dial [Zaj88a]. Dial-up [Zaj88a]. Dickerson [Ano88-47, Ano88-101]. Dictionary [Ano84h]. did [Hig88g]. didn’t [McD88]. Dierstein [Ano88l]. difference [Fåk86]. different [Ano88-107]. differential [Koc83]. Digital [Ano83f, BP89, Hig88-124, Coh88a]. dilemma [Hig88e]. dinosaur [Hig85d]. diodes [Coh88a]. dip [Ano88-106]. director [Hig88e]. Directory [Ano88-48, Kim88]. Disaster [BDE^{+}82, Hig88-36, Hun88b, Zaj89a, Ano86t, Ano88e, Ano88-59, Fri82, Hig88v, Hig88-35, Hig88-65, Hig88-98, Hig88-108, Hig88i, Men85b, Men88b]. disasters [Ano86-48, Ano88x]. discovered [Hig88-62]. discretionary [Kar86, Sch85]. Discussing [Bou88]. Disk [Hig88-142, Ano86t, Hig87p, Hig89w]. diskette [Woo85b]. display [Ano88-132, vE85]. displays [Ano88h, DR88]. Dispute [Hig88-93]. Distributed [Zaj88b, Ano84-30, BBFÅ82, Cal83, Coh87d, Hig88-114, MT88, Plu88]. distribution [Coh88d, Sch82]. Disturbance [Hig88-37]. DMBS [Ano88f]. do [Ano88k, Zaj86]. document [PP88]. documentation [HR88, Mil88c]. documents [Ano88-81, Ano88-82]. DOD [Hig88-38, Ano85a, Mun88a]. DoD/NBS [Ano85a]. DOE [Bur88]. Donald [Hig83f]. Don’t [Ano86q, Bol89a]. door [Ano86h]. Douglas [Hig88-95, Ano88a]. Douglass [Ano83f]. down [Ano87-43, Hig88-120, Hun88c]. downtime [Ano88-36]. DP [Hig88-34, Men85b]. Dudek [Ano88-96]. during [MB88]. dynamic [WCWG86].

easily [Ko88]. Eason [Ano83f]. easy [Ano86z, Ano88-113]. eavesdropping [Ano87p, Hig88-149, vE85]. ECOMA [Paa84b, Paa87a]. ECOMA-12 [Paa84b]. ECOMA-14 [Paa87a]. economy [Hig88-51]. Ed [Ano88-110]. edition [Ano83j]. Editor [Ano83i, Ano84e, Ano84f, Ano88-58, Ano88-92, Ano88-93, Ano89m, Ano89-30, Hig83b, Hig83c, Hig87o, Hig87l, Hig87m, Hig87n, Hig88-153, Hig88-154, Hig88-155, Hig88-156, Hig89h, Hig89i, Hig89j, Hig89k, Hig89l, Hig89m, Hig88-176, Hig89y, Hig82f]. Editorial [Ano84q, Hig82d, Hig82e, Hig83g, Hig83h, Hig84a, Hig84b, Hig84c, Hig85d, Hig85a, Hig85b, Hig85c, Hig88-73, Wil82b, Ano82j, Ano83o, Ano84r, Ano85l, Ano86r, Ano87o, Ano88-49, Ano89l]. editors [Ano82l, Ano82m, Ano85p, Ano86-29, HJH85a, Ano83r, Ano83-27, Ano84x, Ano85o, Ano87v, Ess86, Hig87v]. EDP [Ano83f, Bea86b, Men85a, Paa84a, Web88, vBH88]. EDP-audit [vBH88]. EDP-auditing [Paa84a]. Edpacs [Hig88-99]. EDS [Hig88-64]. Educating [Sta86b]. Edward [Hig88-105]. Effective [Woo83a, Ano83f, Hig88-69, Hig88-126, Ten86a]. effectiveness [BS87]. Effects [vdG84]. efforts [Hig88-46]. EFT [Ano88-137, Joh87]. EFT-POS [Joh87]. EFTS [Guy83]. ELAND [TC88]. Electromagnetic [Ano87p, Hig86a, HF86, Hig88-149, Hig88-150, vE85]. Electronic [Ano86s, Ano83g, Ano86-38, Ano88-111, BR83, DR88, Eke86b, Hig82c, Hig82a, KD86, LB82, MW88, RY86, Sha89, Web89a, vH87]. electronics [Ano88-81, Ano88-82]. Element [LDS^{+}88]. Element-level [LDS^{+}88]. elements [Hig88-124, PR85a]. embezzlement [Sol87]. Emerging [Kee89, War86a, War86b]. emphasis [Hig88-74]. employee [Hig88-121, Men89c]. enabling [BP89]. encipherment [PR85b]. encouraging [Men89b]. encrypted [Ano87r, Ano88-123]. Encryption [Ano88-44, Bet88, Bur88, Hig87k, Hig88-151, Ano88-42, Ano88-43, Ayo83, Ben88, Chr88, Fåk86, Hig87h, Hig87r, Hig87q, Hig88-82, Hig88-169, Kra84, OM84]. end [FG89]. end-user [FG89]. Engineering [Ano88v, Ano88-38, Ano88w, Ano88x, Ano88-27, Ano88-44, Ano88-43, Ano88-50, Ano88-100, Ano88-98, Ano88-99, Ano88-142, Ano88-143, Ano88-144, Ano88-94]. engines [Thu87]. Enhancing [Hig88-138, Woo83b]. Enigma [Hig83e, Ano86-40]. enough [Hig88-46]. ensure [Hig88-85]. entire [Hig88-30]. environment [Ano88-135, FG89, GM85, PW86, PW87b, RMS89, Zar88, vdL86, Hig88-102]. environments [BvS89]. EPA [Hig88-62]. epidemiology [Mur88]. Equipment [Ano83f, Hig83c, Ano88-139, Fåk87a, Mil85, Ser85]. equivalent [Ayo83]. Eric [Ano83j]. error [Hol88, War83]. errors [Gri88b, Hig88-148]. escrow [Ano88m]. espionage [Kap84]. essential [Sch87]. Establishing [Woo86a, Woo86b]. estimating [Ano88-46]. etc [Ano86-48]. Ethernet [Ano88-50]. Ethical [Coh88d]. ethics [Hig88-151]. Eugene [Ano88-64]. Europe [Cha85]. European [Ano84g, Paa84b, Ano87-34, Paa87b, Paa88]. evaluate [Hig87r]. Evaluating [Ano88-51, War83, FM83]. Evaluation [Gar89a, PHEW88, Ano88-28, Bac83, Hig88-172, Koc84]. Evelyn [Ano88-107]. even [Hig88-69]. events [Mal88]. ever [Hig88-115]. Exec [Hig88-39]. Executive [Ano82k, SK82]. executives [Hig88-103]. expectations [Bol89a]. experiment [Hol88]. experiments [Coh87b]. Expert [Ano88-52, Ano88-53, BPR88, Lon87, Mun88a, Hig88-99, Hig88-112, Mal88, TC88, Hig88-46]. exploiting [She88]. explore [TTT89]. exponentiation [BR89]. extending [Car87]. extension [Por82].

F [Ano88-64, HJH85a]. Fåk [Ano83i]. facilities [CI85]. facility [Kim88, Koc83]. Fact [Ano87p, Hig88-145]. factor [Hig88-106]. factors [Por82, Woo87b]. fail [Hig88k]. failures [Ano88-108, HM86]. fantasy [Ano87p, Hig88-145]. Fast [Hen82, Ano88-104, HC88]. Fault [Ano88-54, Hig88-40, Ano88-145, Mal88, Voi88]. fault-tolerant [Ano88-145]. Faulty [Ano88-55]. fax [Hig87-28]. February [Ano88-27, Ano88-100, Ano88-98, Ano88-99, Ano84y, Ano87y, Ano88-115, Ano88-144, Ano89-32]. federal [Hig88-27, Hig88-75, Hig88-41]. Feds [Ano88f]. feel [Byl88, Hig88-61]. fees [Ano88-90]. fever [Dav88]. ff [Ano88-128, Hig88-28, Hig88-46, Hig88-102, Hig88-106, Hig88-130]. fiber [Hig88-158]. fibres [Hig88-101]. fiction [Pol84]. field [Hig88t]. Fighting [Ano88-56]. fights [Hig88-38]. figures [VA88]. file [Ano88-42, CB88, OM84]. files [Ano88-48, Ano88-60, Gof88b, Hig88n, Koc83]. filter [Hig88b, Hig88-172, Hig88-173, Hig88-174]. filters [Ano88-129]. final [Hig88-100]. finger [Hig88-139]. Fire [Hig88-42, Sop88, Men89a, Men89d]. fires [Ano86-48]. firm [Bar88, Hig88-98]. firms [Par88b]. First [Ano83q, Paa87b, Ano83i, Bec83, Hig85e, HE85]. fits [Sch87]. Fitzgerald [Ano88-32, Ano88-97]. five [Gar89a]. flags [Joh87]. flaws [Boz88b]. flexible [Hig88-56]. Floods [Ano86-48]. floor [Gri88a]. Floppy [Woo85b]. flow [Blo82, PP88, RFB82b]. flows [EHH85a]. Flynn [Hig88n]. Focus [Hig88-46, Hig88-55, Ano88-69, Hig87d, Hig88-33]. focuses [Ano88-34]. foil [Ano88-59]. Footnote [Ano88-57]. forces [Hig88l]. forms [Hig88-37, Yeo85]. Fourth [Ano83j, Hig88d]. Framework [BE89, CI85]. Francis [Hig88m]. Frank [Hig88-124]. fraud [Bol89b, Car82, Hig88-132, War83]. free [Hig88-28]. frequency [Hug88]. Friedberg [Hig88-101]. friendly [BB84, Hig88-104, Mil88b]. Full [Ano88-151]. Full-time [Ano88-151]. fully [Hig88-111]. function [McD88, Str88, vBH88]. functional [Thu89]. functions [Ano88-46, Bea86a, Bea86b, MRW89]. funding [Hig88-45]. funds [Ano88-139, KD86, RY86, Sha89, Web89a, vH87]. Fuss [Ano88-59]. Future [Woo82a].

G [Ano82e, Ano83e, HJH85a]. GAO [Hig88-77]. gap [Ano88-80]. gaps [Hig88z]. Gardner [Hig88-72]. GE [Sha89]. General [Cer85, BR89]. Gerbil [Gla88]. German [Ano88-60]. get [Bet88]. Getting [Ano88-138, Men88b]. give [Hig88-87]. given [Hig88-74]. gives [Ano88-95, Hig88-112, Hig88-134]. global [Hig88-51, Sha89]. go [Ano88a, CM88, Hig88g]. Goode [Hig88-103, Hig88-94]. Goodwin [Ano88-102]. Gordon [Hig83e]. Government [EHH85b, Hig88k, Hig88l, Hig88-30, Hig88-35, Hig88-41, Hig88-58, Hig88-66, Hig88-72, Hig88-69, Hig88-62, Hig88-76, Hig88-78, Hig88-79, Hig88-82, Hig88-94, Hig88-100, Hig88-112, Hig88-113, Ano88-130, Men88e]. Gower [Ano83j]. graduate [Har83]. grand [Kah82]. greater [Hig88-115]. Green [Ano88-36]. Greenhouse [Ano88-29]. grew [Ano88-104]. Grim [Ano88-139]. ground [Hig88-28]. group [Buc82, EHH85b, Hig88-97, Paa84b]. groups [Ano87-32]. Grover [Ano88-55]. growing [Ano88-108, Hig88t, Hig88-92]. Gruzs [Hig88-28]. guard [Ano88-121]. Guarding [Ano88-61]. guards [Paa84a]. Guest [Ano87q, Wil82b]. Guidance [Fra88]. guide [Ano82d, Ano82k, Ano83g, Hig82b, HH82, Hig88-103, RFB82a, SK82]. Guidelines [Ano88-62, Mil88c, Hig88-66].

H [Ano88-135, Hig88-101, Gof88a, Hig88-43, Jon88]. H-P [Gof88a, Hig88-43, Jon88]. habits [Hig88-113]. Hacker [Hig88-122, Her87, Hig88-53]. Hackers [Hig88-157, Hig88c, Ano88-95, Hig88-130]. hacking [LS87]. hailed [Ano88-112]. Hal [Ano84i]. half [Zaj88c]. Hamilton [Ano88-68]. handbook [PES88, Fåk87b]. handle [Mal88]. handling [LAN86, PP88]. Handwriting [Ano88-63]. happen [Hig88-55]. happens [Ano88k]. Hard [Ano86t, Hig87p, Ano88-121]. hardbound [Ano83g]. hardcover [Ano84h]. Hardware [Ano88-64, Ano88-65, Fåk86, Hig87r, Hig88-119]. harmful [Hig88-92]. Harold [Ano83h, Hig88-82]. Harper [Hig88-101]. Harrison [Ano88-145]. Harry [RFB82b]. hash [MRW89]. Hashing [Hig88-44]. Hayden [Hig83d]. Head [Hig88l]. health [Kie88]. healthy [Hig88-113]. Heaton [Ano88-62]. held [Ano88-71, Ano88-104, Fåk87b]. Henniker [Ano88-62]. Henniker-Heaton [Ano88-62]. here [Hig88-55]. heterogeneous [PW86, PW87b]. Heyden [Ano84e]. hierarchical [Coh87d]. High [Ano87r, Ban82]. High-speed [Ano87r]. Highland [Ano83h]. highlighted [HJH85d]. Highlights [Ano88-66]. Hill [Ano84i, Hig83e]. hinder [Hig88-76]. hiring [Bol88c]. Historical [PHEW88]. history [Hig88-146]. HK [Ano88-67]. hobbyists [Ano88-60]. Hold [Ano88-68]. holistic [dB85]. Holland [Ano83i, EHH85a, HJH85a]. Hong [Ano88-69]. Horatius [Ano88-70]. Horse [LAN86]. hospital [LK84, Ano88-71]. Host [Erd86]. Houghton [Ano84j]. House [Hig83b, Hig88-45]. Howard [Ano88-45]. Howkins [Hig88-100]. HP [Hig88p]. HR [Hig88-49]. Hughes [Ano88-42]. human [Ala86, Por82, Woo87a]. humidity [Sop88]. hunt [Her88]. hurt [Gri88b]. hut [Hig83e]. hype [Hig88-50]. hyperbole [Hig88-148].

I/S [Mou87]. IBM [Boz88a, Buu84, Gof88b, WCWG86]. IC [Ano88-73]. idempotent [PR85a]. identification [Ano88-126, Hig88-118]. Identifying [Spe87]. IEEE [Ano88-128, Hig83f, Hig83f]. IFIP [Ano83i, Fåk87b, Ano83p, Ano83q, Ano83s, Ano84s, Ano84t, Ano85m, Ano86u, Ano87s, Ano88-66, Ano88-74, Ano88-75, Ano88-76, Ano88-77, Ano88-78, Ano88-79, Ano89n, Ano89o, Ano89p, Ano89q, Ano89r, Ano89s, Ano89t, Ano89u, Bec83, Hig88-162, Kug86, Owe86, Paa87c]. IFIP/sec [Ano83i, Ano83p, Ano83q, Ano86u, Ano87s, Hig88-162]. IFIP/Sec’84 [Ano84s, Ano84t]. IFIP/Sec’86 [Paa87c]. ignorance [Hig88-148]. II [Ano88-156, HF86, Hig87h, Hig87j, Hig88-169, Hig89w, She88]. iii [Hig83c]. im [MB88]. immune [Woo87a]. immunity [Hig88-87]. Impact [Hig83i, Guy83, Men85b, Mur86b]. implementation [Ban82, Buu84, Dos86, Erd86, Hig88-128]. implementations [BR89]. Implementing [Car88, Jun87, Jun88, Koc83]. implications [Coh88c, KD86]. Important [Elo88, Hig88-40]. improved [Por82]. improvement [HC88]. improvements.Part [JO89a, JO89b]. Improving [Pos87, Deh89]. incidents [Mou82]. includes [Hig88-88]. Incorporating [Yeo85]. increase [Hig88-45, Hig88-101]. increasingly [Hig88-40]. index [Ano82c, Ano82r, Ano83d, Ano83z, Ano84d, Ano84-31, Ano85f, Ano85x, Ano86g, Ano86-44, Ano87g, Ano87-38, Ano88i, Ano88-94, Ano88-146, Ano89c, Ano89-40]. indicted [Raa88]. individual [Ano86s, TTA^{+}88]. indoor [Hig88-124]. Industrial [Kap84, Ano88-53]. Industry [Par88a]. inexpensive [Hig88-76]. infect [Hig88-30, Hig88-137]. infecting [Ano89-39]. inference [Thu87]. Infinite [Top84]. info [CM88]. Informatics [Ano84-32]. Information [Ano86w, Ano88v, Ano88-38, Ano88w, Ano88x, Ano88-27, Ano88-44, Ano88-43, Ano88-50, Ano88-98, Ano88-99, Ano88-142, Ano88-143, Ano88-144, Gag88a, Hig83b, Hig87u, Hig88y, Hig88-47, Hig88-51, Hig88-52, Hig88-101, Hig88-105, Hig88-118, Hig88-114, Hig88-115, Paa87c, Pol84, Woo87b, ALN87, ALN88, Ano83g, Ano86-38, Ano87w, Ano88y, Ano88z, Ano88-65, Ano88-81, Ano88-82, Ano88-111, Aue85, BB84, BB88, Bol88b, Bol88c, Coh87a, Coh87d, Coh87e, Coh88a, Coh88b, FG89, Gua87, Har88, Hig82c, Hig83d, Hig88-152, HE85, LK84, Lun88, Men88e, Men89c, MB89, PR85b, Sch87, Sha89, Woo83a, Woo83b, WZ86, Woo87a, Woo88, Ano88-100, MB88, Ano84i]. informational [Bot83]. Informationstechnik [MB88]. InfoWorld [Hig88q, Hig88o, Hig88-71, Hig88-63, Hig88-64, Hig88-93, Hig88-96, Hig88-97, Hig88-120, Hig88n]. Innovative [Vil88]. insecurity [Hig88y]. INSPEC [Ano88-44, Ano88v, Ano88-38, Ano88w, Ano88x, Ano88-27, Ano88-43, Ano88-50, Ano88-100, Ano88-98, Ano88-99, Ano88-142, Ano88-143, Ano88-144]. install [Hov83]. Institute [Ano88-128, Ano86w]. Insurance [Ano88-80, Hig88-53, Hig88z]. insured [Hig88-111]. Insuring [Ric86]. integrated [Koc83, LK84]. Integrating [TR86b]. Integration [She88]. Integrity [Ano86w, BS82, Ano84i, Coh87c, Coh88b, Ram89a, Sha89]. Intellectual [Ano88-81, Ano88-82, Hig88-51]. interbank [FVTS87]. Interconnection [LN88, Ram89a]. interest [Paa84b]. interface [Hig88-93]. interfaces [Hug88]. interference [Hig88-150]. Intergrating [TR86a]. Intermittent [Koc84]. Internal [Har86, Ano88-124, GM85, Woo86a]. International [Ano82l, Ano83r, Ano83-27, Ano84x, Ano85o, Ano87v, Ano88-83, Ano88-84, Ano88-85, Ano88-86, Ano88-87, Ano88-88, Ano89v, Ano89w, Ano89x, Ano89y, Ano89z, Ano89-27, Ano89-28, Ano89-29, Blo82, Ess86, Hig87v, Kra84, Ano82m, Ano85p, Ano85v, Ano86-29, Com84, EHH85a, HJH85a, Ano84v, Ano86w]. Internet [Gar89b]. introduce [Cha88]. introduction [Har89, Hig83d, Hig88-54, RFB82b]. intruders [Ano88-45]. Investigating [Hig88-163]. investigation [Mou82, Sta86a]. investigators [Sta86a, Sta86b]. IP [Ram89b]. IPSA [Ano84w]. ISO [Kra84]. isolated [Hig88-28]. Israeli [Rad89]. issue [Ano88g, Ano88-110, Ano88-89, DeM89, Hig86c, Hig86d, Hig86e, Hig86f, Hig87s, Hig87t, Hig88m, Lew88a]. issues [Ano88-136, Cha88, Coh88d, Elo88, Hig88-84, JL89, MT88, Men88e, Owe86, RH86, War86a, War86b, Zar88, Hig88-105]. IV [Hig88d]. ix [Ano84h].

J [Ano88-32, Ano88-145, Ano88-140, Hig82b, Hig88-48, Hig88-93, Hig88-96, Hig88-99, Hig88-101, Hig88-105, Hig88-113, Hig88-137, RFB82a].

Jahrezehnts [MB88]. James [Ano83g, Ano84j, Hig82c, Hig88-98, Ano88g]. January [Ano88-44, Ano88-43, Ano82n, Ano83u, Ano88v, Ano88-38, Ano88-37, Ano88t, Ano88-30, Ano88-31, Ano88-33, Ano88-53, Ano88-54, Ano88-63, Ano88-94, Ano88-99, Ano88-141, Ano88-142, Ano88-143]. Japan [Ko88]. Jean [Hig88-42]. Jeff [Ano88-102]. Jerome [Ano88-133]. Jerry [Ano84h]. Jim [Hig88-92]. Jo [Ano88-139]. John [Ano82e, Ano82d, Ano83h, Ano83j, Ano84e, Ano84f, Ano84h, Ano88-128, HJH85c, Hig82b, Hig88-91, RFB82a, Woo85a, Hig88-129]. Johnston [Hig88-93, Hig88-96]. Jonathan [Hig88-92]. Jones [Ano88-47]. Joseph [Ano83h, Hig88-82, Hig88-47]. Jr [RFB82b]. Jr. [Ano88k]. Juan [HJH85a]. Jubilee [Ano84w]. Julie [Ano88-61]. July [Ano88-33, Ano88-28]. jump [Par88b]. June [Ano88-54, Ano88-128, Hig88-48, Hig88-92, Ano82p, Ano83w, Ano85r, Ano86-36, Ano87z, Ano88y, Ano88-103, Ano88-117, Ano89-33]. Justice [Ano83n, Bol88d, Ano83n, Ano88-90].

Kahn [HJH85b, HJH85b]. Karon [Hig88-50]. Kask [Ano88-35]. Katzan [RFB82b]. Keefe [Hig88-55]. Ken [Ano88-126, Hig88y]. Kerr [Ano88-112]. Kevin [Ano88-32, Ano88-97, Hig88-66, Hig88-78, Hig88-121]. Key [Joh87, Ano88-42, Ano88-45, Gag88a, Hig88-121, Jan86, Lon87, MP87, Mul84, PR85a, RY86, Sal88, Sch82, Wil82c]. key-storage [MP87]. keys [Ano88-111, Hig87q]. kill [Hig88-136]. Kingdom [Hig88-59]. Kirvan [Hig88-126]. knapsack [Hen82]. know [Hig88-140]. knowledge [SK88]. knowledge-based [SK88]. Knowles [Hig88-114]. Knox [Hig88-105]. Kong [Ano88-69]. Konheim [Ano82e]. KOs [Hig88-42]. Krayem [Hig88-118]. Krull [Hig88-104].

L [Ano88-65, Ano88-113, Hig88-30, Hig88-135, Woo85a]. L.A [Hig88-98]. lab [Mun88b]. lack [Hig88-35]. lacking [Hig88-34]. Lambert [Ano88-147, Ano88-134]. LAN [Ano88-112, Hig88-57, Hig88-58]. Lane [Hig83e]. language [Elo85, Sal88]. languages [Kug86]. LANS [Cha88, Hig88-40]. Lapid [Ano88-123]. laptops [Hig88-69]. Large [She88, Ano86x, Woo86b]. Larry [Ano88-138]. last [Hof89a, MB88, Web88]. Laurie [Hig88n]. law [Ano84o, Ano87e, Ano87f, Hig87e, Hig88s, Hig88x, Hig88-60, Hig88-75, Hig88-164]. Lawrence [Ano88-42]. laws [Hig88m]. Lawsuit [Ano88-68, Bar88]. layer [VK85]. leaking [Hig87-30]. leaky [Boz88a]. learning [Mal88]. leaving [Bri88]. Lee [Ano88-36]. Legal [Ano88-91, Hig88-59, Zaj89b, Gof88b, Hig88-44]. legislation [HJH85d]. Lehigh [vW89]. Lessons [Ano88-36, Ano88-148, Hoe86]. Letters [Ano88-92, Ano88-93, Ano89-30]. letzten [MB88]. level [DJT87, Hig88-67, LDS^{+}88]. liable [Hig88-39]. libraries [Mur82]. Library [Ano83e]. license [Ano88g, Ano88-110]. licensing [Ano88-90, Ano88-89, Ano88-140]. lie [Saa87]. life [BE89, Dos86, TR86a, TR86b]. lifetime [Car87]. light [Ano87-42]. limited [Ano86y]. limits [Ano88-46]. Linda [Ano88-29]. line [Ano88-59, Ano88-55, EHH85a, Koc84]. lines [Ano88-64, Kah82, Zaj88a]. link [Ano86-27, Hig88-79]. links [Ano88-99]. List [Hig83a, Ano88-129]. Lister [Hig88u]. Literature [Ano85e, Ano82a, Ano82b, Ano82i, Ano83a, Ano83b, Ano83c, Ano84a, Ano84b, Ano84c, Ano85c, Ano85d, Ano85b, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87e, Ano87c, Ano87f, Ano87d, Ano89a, Ano89b, HH82, Hig88j, Hig89a, Hig89b, Hig89c, Hig89d, Hig89e, Hig89f, Hug88]. Lobel [Ano88-133]. Local [Ano88-94, Ano88-72, Har86, Hig88-102, JL89, Sch82, TC88]. Lock [Ano87t, Ano88-45]. locks [Ano88-111]. login [Deh89]. London [Hig83e, Hun88c]. Long [Hig88-60, PH87]. longstanding [Hig88-93]. Look [Byl88, Hig88-88, Hig88-61]. Look-and-feel [Byl88]. looms [Byl88]. Loophole [Ano88-95]. Loose [Hig88-94]. loss [Ano83f]. losses [Hig88-125]. Lotus [Ano88-96, Hig88r, Hig88-95]. Louise [Ano83e]. LRAM [Gua87]. Ltd [Hig83e]. Ludlum [Ano88-151]. lure [Ano86x]. lures [Par88b]. Luxembourg [Ano84g].

M [Ano84f, Ano84h, Ano88-134, Ano88-147, HJH85c, Hig88-28, Hig88u, Hig88-47, Hig88-68, Hig88-101]. MA [Ano83f, Hig83b]. Mac [Hig88-93, Hig88-97, Hig88-109, Hig88-62, Hig88-63]. MAC-based [Hig88-109]. MacDonald. [Ano88-107]. Mace [Ano88h, Hig88-97]. machines [Ano88-95, Hig88-149]. Macintosh [Hig88-64]. Macmillan [HJH85b]. made [Ano86z, Fåk87a]. Madsen [Ano88-52]. Magazine [Hig88-48, Hig88-92]. mail [MW88]. mainframe [Ano86-27, Ano87-40, Ano88-99, Ano88-121, Req88]. Maintaining [Coh88b]. maintenance [Ano88-30, OR88]. Make [Her88, Hig88-65, Hig88-56]. makes [Hig88-126, Hun88a]. Making [BR83, Fis84]. Manage [Ano88-97]. Manage/38 [Ano88-97]. manageable [Fåk87a]. Management [Ano84-32, Won88b, Woo87b, Ano84e, Ano84f, Ano88-133, Ano88-138, Ano88-147, BR83, Bou88, DeM89, DJT87, Fåk87a, GPR^{+}88a, Gua87, Hig88-103, Lon87, MT88, Mou84a, Mun88b, Plu88, RY86, Sch85, Sch87, Thu87, Won85, Woo83b, Ano84i, Hig82b, RFB82a]. managerial [Hig87c]. managers [Ano82d, Ano83g]. Managing [Bea86b, Bol88b, NCCW88, HJH85c, Hig82c]. Mann [Ano88-134, Ano88-147]. manual [Ano83j]. Manufacturers [Hig83c]. March [Ano88-53, Ano88-144, Hig88-78, Hig88-99, Hig88-102, Ano85q, Ano86-33]. Mark [Ano88-108, Hig88-103, Hig88-94]. Market [Ano88-33, Hig88-84, EHH85a]. Markoff [Hig88-91, Hig88-129]. Marlene [Ano88-95]. Martin [Hig88-98]. Marvin [Ano84f]. matching [Gri88a]. mathematical [Gle89]. mathematics [Mul84]. Matzkin [Hig88-92]. May [Ano83i, Hig88q, Hig88m, Hig88o, Hig88l, Hig88-42, Hig88-52, Hig88-66, Hig88-71, Hig88-62, Hig88-95, Hig88-98, Hig88-113, Hig88-121, Hig88-122, Hig88-124, Hig88-126, Hig88-137, Hig88-135, Hig88-50, Hig88-87, Ano84-28, Ano89-31, Hig88n]. maze [Ano88-113]. MBS [Hig88-66]. McGraw [Ano84i, Hig83e, Hig88-112, Ano88-90]. McKinney. [Ano88-72]. McLeod [Hig88y]. MCSC [Hig88-67]. MD [Hig83f]. means [Woo89]. measure [Bol89a]. Measurement [Paa84b, SB84]. measures [BBFÅ82, Hig88-85, Hig88f, Woo85b, dB85]. measuring [Bol88a]. mechanisms [JO89a, JO89b, VK85]. Media [Hig88-148, Hig88-68]. mediated [BB84]. medical [Ano88-71]. Mediocrity [Ano88j]. meet [Blo82]. meeting [Ano88-66, HE85]. meetings [Ano87x]. meganetwork [Mad88]. member [Ano85o, Ano86-29, Ess86, Hig87v]. members [Ano82m, Ano83r, Ano83-27, Ano84x, Ano85m, Ano85p, Ano87v, Hig83f]. Memoriam [Ano84u]. memory [Kol88]. Menkus [Hig88k]. merit [VA88]. Message [Ano87u, Chr88, Sha89, MRW89]. messages [FVTS87]. metaphysical [Hig88-88]. method [Hig88-82]. Methodology [CW83, PES88, BE89, FM83, PPT89]. methods [BB84, Coh88c, SK88]. metrics [Ano88-142, Ano88-143]. Metro [Raa88]. Metro-North [Raa88]. Michael [Ano88-69, Hig88-101, Hig88-102, Hig88-105, Ano88-148]. Micro [Hig88-69, Hig88-84]. Microcomputer [Ano86y, Ano86z, Ano86-27, Ano88-98, Ano88-99, Hig85f, Hig88-70, Zar88, All88, Ano83h, Ano86e, BR83, Hig86b, Hig87r, Hig88-161, Hof89b, Voi88]. Microcomputer-mainframe [Ano86-27]. Microcomputers [Ano88-100, Ano86-37, Hig83i, KG87]. microprocessor [Ano88-73]. microprocessors [BR89]. micros [Hig88g, Hig88-127]. Microsoft [Ano88-101, Hig88n, Hig88q, Hig88m, Hig88p, Hig88-43, Hig88-71, Jon88, vdAvE86, Hig88-96]. Mifflin [Ano84j]. Millard [Ano88-140]. million [Ano88-127, Bec83]. Milnet [Sto88]. mind [Ano88-110]. minicomputer [Ano88-134]. minute [Bol89b]. Minutes [HE85]. MIPS [Hig88-94]. MIS [Ano88-102, Ano88-103, Bol88a, Bol89a, Hig88w, Hig88-137, Ko88]. misuse [LS87, Raa88]. Mitch [Ano88f]. MLS [TTT89]. MLS/DBMS [TTT89]. model [BB88, BvS89, FM84, Hig88-78, Kug86, Woo87a]. modeling [BS87]. modelling [Lun88]. Models [Coh89b]. modem [Ano87-35]. modems [Ano88-98]. Modern [Ano86-28]. modify [Hig86b]. Modular [PR85b, BR89]. modules [She88, WZ86]. money [LB82, Par88b, Sha89]. Monitoring [Hig88-72]. moral [BBR87]. Most [Ano88-104, Ano84j, Hig88-35]. mostly [CM88]. Moulton [Hig82a]. moves [Gri88a]. much [Hig88-46]. Multi [DJT87, Bec83, RFB82b]. Multi-level [DJT87]. multi-million [Bec83]. Multi-national [RFB82b]. multidestination [MW88]. multidisciplinary [Mou82]. Multilevel [MT88, Car88, Thu89]. multinational [Woo86b]. multiprocessing [Ano88-54]. Munro [Hig88-41]. Murray [Hig88-46, Ano88-135]. must [Hig88-132]. MVS [Hig88-79, PB83, PH84, PH86]. Myths [Hug88].

NASA [Ano88-60, Ano88-105]. NASD [Ano88-106]. nasty [Hig88-73]. National [Kie88, Ano83e, Ano88-52, Hig85e, Hig88d, Hig87d, LN88, RFB82b]. Navy [Hig88-58]. NBS [Ano85a, Ano88-114, Ano88-107, HH82, Hig88-45]. necessity [Ten86a]. need [Hig88-28]. needed [Hun88b]. needs [Com84, Mun88a, Mun88b]. Neil [Hig88-41, Hig88-48]. Nested [San88]. Net [Ano88-108, Ano88-52]. Network [Ams88, Ano88-27, Ano88-72, Ano88-137, Bur88, Cal83, Coh85, Coh87e, Hig88-76, Hig88-102, JL89, RH86, Sha89, Top84, Var89]. Networks [Ano85n, PW87a, Ano88-50, Ano88-94, Ayo83, Chi88, Coh87a, Coh87d, Coh88a, Har86, Hig88-42, Kar86, MP87, Sch82, TC88, Hig83f, Woo85a]. Neumann [Ano88-123]. Neutral [Ano88-109]. Never [Ano88-110]. News [Ano83s, Ano86-30, Hig88l, Hig88-30, Hig88-35, Hig88-41, Hig88-58, Hig88-66, Hig88-72, Hig88-69, Hig88-62, Hig88-76, Hig88-78, Hig88-79, Hig88-82, Hig88-100, Hig88-112, Hig88-113, Hig88-126, Sch88]. next [Hig88-49]. Nick [Hig88o, Hig88-63, Hig88-64, Hig88-120]. Nina [Hig88-124]. Nine [Ano83f]. Niv [Ano88-123]. NJ [Hig83d]. No [Ano88-113, Ano88-96, Hig88-82]. noise [Hig88-28]. Non [Ano88-53, BvS89, Hig83f]. Non-industrial [Ano88-53]. non-members [Hig83f]. non-secure [BvS89]. North [Ano83i, EHH85a, HJH85a, McD88, Raa88]. North-Holland [Ano83i, EHH85a, HJH85a]. Nostrand [RFB82b]. note [Ano86-31, Ano87x, Hig82f, Hig88h, Mul84]. Notes [Men83, Ano87-31, Hig87y, Hig87z, Plu88]. November [Ano82o, Ano83v, Ano84-27, Ano88w, Ano88u, Ano88z, Ano89-36]. NSA [Ano84j, Ano88-114, Ano88-107, Ano88-136, Hig88-76, Hig88-77]. NTIS [Ano87w, Ano88y, Ano88-28]. nuclear [Hoe86]. number [Por84]. numbers [Ano86x, Saa87].

object [KTT89]. object-oriented [KTT89]. observability [PW87a]. Observations [Mur83]. obtain [Hig88-160]. October [Ano88v, Ano88-37, Ano88w, Ano88t, Ano88z, Fåk87b, Hig83c, Ano87-28, Ano88-39, Ano88-40, Ano88-41, Ano88-119, Ano89-35]. off [Ano88-110, Hig88-38, Jon88]. office [Hig88-167, Kee89, Won88a, Zar88, Ano83t, EHH85b, Hig85d, Hig87w]. Officers [Mou87]. officials [Hig88-41, Hig88-66]. oil [Hig88-135]. old [Wil82a]. Oliver [Ano83j]. OMB [Hig88-78]. on-line [Ano88-55, EHH85a, Koc84]. one [Ano88-62, Bol89b, HC88]. Open [LN88, Kru89, Ram89a, Hig88c]. Operating [Hig88-79, Hig86b, PB83, Pos87, Voi88]. operation [Hoe86]. Operational [PHEW88]. operations [Erd86]. opinions [Ano84o]. OPM [Hig88-66]. optical [Hig88-101]. optics [Hig88-158]. Optimized [BR89]. options [Zaj89b]. ordered [Ano88-90]. ordering [Coh87e]. orderings [Coh87a, Coh87d]. Organization [Kra84, BE89, Bea86a]. Organizational [Str88]. organizations [Ano82q, Ano83x, Ano83y, Ano84-29, Ano85u, Hig88-30]. oriented [BB88, Har83, KTT89]. OSI [Hig88-114, Hor88, LN88]. OSIS [Rih87]. OT&E [PHEW88]. Other [Ano87x, Zaj88c]. outline [Hig88-67, Neu82]. outsmart [Fif89]. overlook [Ano86q]. overview [Ano83e, Hig88-165, Won88b]. ownership [MB89].

P [Ano88k, RFB82a, Ano88-128, Hig88k, Hig88n, Hig88q, Hig88o, Hig88-28, Hig88-35, Hig88-41, Hig88-46, Hig88-47, Hig88-48, Hig88-71, Hig88-69, Hig88-62, Hig88-64, Hig88-73, Hig88-78, Hig88-79, Hig88-82, Hig88-93, Hig88-95, Hig88-97, Hig88-98, Hig88-100, Hig88-102, Hig88-106, Hig88-120, Hig88-112, Hig88-113, Hig88-115, Hig88-121, Hig88-130, Hig88-135, Gof88a, Hig88-43, Jon88]. package [Ano86-41, Fra88, Hig87k]. packages [BR83, Elo83]. Pages [Ano82n, Ano82p, Ano82o, Ano83u, Ano83w, Ano83v, Ano84z, Ano84y, Ano84-28, Ano84-27, Ano85t, Ano85r, Ano85q, Ano85s, Ano86-35, Ano86-36, Ano86-33, Ano86-34, Ano87-30, Ano87-27, Ano87-29, Ano87y, Ano87z, Ano87-28, Ano88-116, Ano88-118, Ano88-120, Ano88-115, Ano88-117, Ano88-119, Ano89-38, Ano89-34, Ano89-37, Ano89-32, Ano89-33, Ano89-31, Ano89-36, Ano89-35, Ano84f, Ano84h, Ano84i, Ano84j, EHH85a, EHH85b, HJH85a, HJH85b, HJH85c, Woo85a, Ano84e, Ano84g]. Pakistani [Web89b]. palace [Ano84j]. panel [LS87]. paperback [Ano83g, Ano84h]. parallel [Koc84]. Park [Hig83d]. Parker [Hig88q, Hig88-71, Hig88-96, Ano88s, Ano88-101]. Part [Ano88-156, Hig87j, HF86, Ano88-62]. partial [Coh87a, Coh87d, Coh87e]. partners [Mil88a]. passkey [Lew88b]. password [Ano86y, BB84, CMRW88, JO89a, JO89b, Lew88b, Men88c, Por82, Woo83a]. Passwords [Hig87x, RMS89, Ano86f, Men88d]. path [BvS89]. Patricia [Hig88-55]. Patton [Ano88m]. Paul [Hig88-50, Hig88-106, Hig88-126]. Paula [Ano88-121]. pauper [Bis86]. pay [Ano88-90, Hig88-46]. payment [BP89]. PC [Hig88-50, Hig88-48, Hig88-92, Hig88-135, Ano88-121, Hig88-56, Hig88-65, Hig88-85, Hig88-140, Rad89]. PCs [Ano88-65, Hig88-56, Hig88-67]. Pemmaraju [Hig88-130]. Pendleton [Ano88-69]. penetration [PPT89]. Penguin [Hig83e]. Pentagon [Hig88-80]. Peoples [Ano86-30]. perfectly [Por84]. performance [Ano88-28, Bol88a, Paa84b]. permutations [Ayo83]. Perry [Ano82d]. person [Coh88b]. Personal [Ano87-41, Hig88-81, Sch84, Ano88-122, BB88, FG89]. Personnel [Zaj88c, Bol88c]. perspective [Cou88, Har86, Sny88]. Peter [Ano88d, Hig88-69, Hig88-89]. phases [Hig88-71]. Philip [Ano88-72]. Physical [Ano86-37, Men89e]. physician [Hig87i]. Physics [Ano88v, Ano88-38, Ano88w, Ano88x, Ano88-27, Ano88-44, Ano88-43, Ano88-50, Ano88-100, Ano88-98, Ano88-99, Ano88-142, Ano88-143, Ano88-144]. Picking [Hig88-82]. Pipe [HJH85a]. piracy [Ano88s, Ano88-56, Ano88-144, Bum84, Hig88-120]. pirates [Ano88-67, Hig88s]. pitch [Hig88-135]. Pitfalls [Bol88a, Gof88b]. pivotal [Ano88-46]. Placement [Ram89a]. placing [She88]. plagued [Hig88-135]. plaintext [vdAvE86]. plan [GPR^{+}88a, GPR^{+}88b, Hig88-35, Hig88-66, Hig88-65]. planners [Ano88-151]. Planning [Hig88-83, Woo89, Ano88-137, Ano88-147, Fri82, Hig88-34, Hig88-36, Men85b, Sei84, SK82, Woo88, Hig88-98]. plans [Ano88m]. plant [Hoe86]. point [MB88]. points [Hig88-72, HC88]. Police [Raa88]. Policies [Ano88-122, Woo82b, Lun89]. policy [Ano88-124, Elo88]. polyalphabetic [CR88]. pools [Hig88v]. poor [Coh88b]. portability [Ano88-31]. ports [Ano88-64]. POS [Joh87]. Positive [LS87]. postscript [Ano87n]. Potent [WZ86]. Potential [SK88, Ano89-39, JO89a, JO89b]. potentially [BvS89]. Power [Hig88-66, Hig88-78, Hig88-84, Hig88-121, Ano87-35, Ano88-102, Ano88-106, Ano88-149, Hig88-90, Hoe86]. power-on [Ano87-35]. pp [Ano82e, Ano82d, Ano83h, Ano83g, Ano83j, Hig82c, Hig82b, Hig83b, Hig83d, Hig83e, Hig83f, Hig88m, Hig88l, Hig88y, Hig88-30, Hig88u, Hig88-42, Hig88-50, Hig88-55, Hig88-53, Hig88-52, Hig88-58, Hig88-66, Hig88-72, Hig88-68, Hig88-63, Hig88-76, Hig88-91, Hig88-92, Hig88-103, Hig88-94, Hig88-96, Hig88-99, Hig88-89, Hig88-101, Hig88-105, Hig88-104, Hig88-118, Hig88-114, Hig88-122, Hig88-124, Hig88-126, Hig88-129, Hig88-137, RFB82a, RFB82b]. pp. [Ano83e, Ano83f, Hig83c]. pp.$39.00 [Ano83i]. practical [Coh89b, Hug88, Mou84b, Owe86]. practice [Ser85]. praises [Par88a]. Pre [Dos86]. Pre-implementation [Dos86]. predictor [CMRW88]. preparing [Mil88c]. Prerequisites [MS89]. Press [Ano83f]. Preventing [Wis89, Ano88-36]. Prevention [Ano88x, Hoe86, Gol83]. Preventive [Hig88-85]. Price [Woo85a]. primer [Bum84, Ano82e]. Principles [Gua87, Bon83, Par84]. Printing [EHH85b]. Privacy [Lea86, Ano86s, Ano88-124, Aue85, BB88, Hig83c, Hig88-81, Jay82, Kur82, War86a, War86b, WC88, Ano84g]. privacy-oriented [BB88]. private [ALN88, Mil88a]. privilege [Car88]. problem [Hig88-58, MW88]. problems [Hig88-74, Hig88-115, Hig88-131, Men89a, Zaj88b, Hig88-99]. procedures [Gua87, Hig88-100, Mur82]. Proceedings [Ano83i, EHH85b, LS87]. process [Elo83, NCCW88, TR86a, TR86b]. Processing [Ano88-123, Ano84h, Ben88, Cal83, CW83, Hig82a, HE85, Men83, Men88b, Men89d, Zaj88b]. Producing [Mur86a]. Product [Hig88-166]. products [Ano88-112, Hig88-165, Hig89g]. Professional [Ano82q, Ano83x, Ano83y, Ano84-29, Ano85u, Aue85]. profiles [PHEW88]. Program [Hig88-167, Ano86y, Ano86-39, Ben88, GPR^{+}88a, GPR^{+}88b, Hig82c, Hig87i, Hig88o, Hig88-123, Hig89x, Mur82]. Programmer [Hig88-86, HP84]. programs [Gar89a, Hig88-48, Kie88]. progress [Lea86]. progressed [Web88]. project [Dos86]. proliferation [FG89]. promises [Ano88-96]. promotion [Kie88]. Proper [Hig88-100]. property [Ano88-81, Ano88-82, Hig88-51, Ric88]. proprietary [MB89]. protect [Ano88-65, Hig88-161, Hig88-175, Sop88]. Protecting [ALN88, MB89, Ano88-102, Hig88-117, Men89c, Ano83h, Ano83g]. Protection [Ano86-38, Coh87a, Ano84g, Ano86-32, Ano88y, Ano88z, Ano88-64, Ano88-91, Ano88-96, Coh87c, Coh87e, Har83, Hig85f, Hig87i, Hig88-59, Hig88-165, Hig88-118, Hig88f, HE85, Kug86, LK84, Men89d, SB84, dC85, vdAvE86, Hig88-95]. protection-data [Ano84g]. protocol [VK85]. protocols [LN88, Mil88a, Ram89b, Var89]. Prototyping [TTT89]. provably [Coh88a]. provides [Ano88-45]. Pseudorandom [Por84]. Public [Ano88-124, Mil88a, Buc82, Gag88a, Mul84, PR85a, Sal88, Sch82, Wil82c]. public-key [Mul84, Sal88]. Public-private [Mil88a]. publication [Wei83]. publications [Ano88b, Ano88c]. Publishers [Hig82c]. Publishing [Ano83g, Ano83j, HJH85b]. Pujals [Hig88-47]. purpose [BR89]. puzzle [Ano84j]. puzzled [Hig88-41].

Q.E.D [Hig83b]. quality [Ano88-142, Ano88-143, BPR88, Hig88-84, Sch87, Sny88, TR86a, TR86b]. Quarterly [Hig88u, Hig88-53, Hig88-68, Hig88-103, Hig88-104]. questions [Lun89]. quit [Zaj88d].

R [Hig82b, Hig88-104]. Rachel [Ano88s, Ano88-101, Hig88-96, Hig88q, Hig88-71]. Rada [HJH85a]. radiation [Ano88-132, Hig86a, HF86, vE85]. Radio [Hug88]. raid [Ano88-67]. raids [Hig88-120]. raises [Hig88m]. raising [CMRW88]. Ralph [Hig88-68]. RAM [Ano86-39, Hig88-130]. RAM-resident [Ano86-39]. RAMSS [PES88]. Random [Ano87-31, Ano88-125, Hig87y, Hig87z, Hig88-168, Hig89n, Hig89o, Hig89p, Hig89q, Hig89r, Hig89s, Hig89t, Hig89u, Por84]. Raysman [Ano88d]. reading [Ano86-43, Ano87-42]. ready [Hig88l, Zaj89a]. Real [Mal88, Voi88]. Real-time [Mal88]. realities [Hug88]. Reality [Pol84]. really [Zaj89a]. reason [Zaj86]. reasons [Bol89a]. recognition [Ano88-33, Ano88-63]. record [Ano86s]. Recovery [Kol88, Ano86t, Ano88x, Ano88-59, BDE^{+}82, Hig87p, Hig88v, Hig88-34, Hig88-36, Hig88-65, Hig88-108, Hun88b, Men85b, Men88b, Zaj89a]. recur [Ano88-106]. Red [Hig88-88, Hig88i]. reducing [Hig88-89, MP87]. reduction [All88, Ano88-34]. redundancy [Hig88-89]. referees [Ano88-83, Ano88-84, Ano88-85, Ano88-86, Ano88-87, Ano88-88, Ano89v, Ano89w, Ano89x, Ano89y, Ano89z, Ano89-27, Ano89-28, Ano89-29]. reference [Kug86, Woo87a]. Reflections [RB82]. registration [Ano88h]. regs [Hig88-35, Hig88-67, Hig88-76]. regulate [Gri88a]. regulation [HJH85a, RFB82b]. Rein [Hig83b, Hig83c]. Reinhold [RFB82b]. related [Bra88, Men89a, Sol87]. relational [CJ86, Thu87, vdL86]. releases [Ano88-96, Hig88o]. reliability [Ano88-46, Mur86b]. Reliable [GPR^{+}88a, GPR^{+}88b, Hig88-90, Hig88-144, Mur86a]. Reliance [Hig88-101]. remark [MRW89]. remedy [Boz88a]. Remember [Ano86-40]. reminder [Ano87-36]. reminders [Ano86-42]. Remote [Ano88-126, Ano87-35]. removes [Hig88-95]. Report [Ano87-32, Ano88-127, Fåk87b, Hig87-27, Hig88d, Ano84j, Ano86-45, Ano86-49, Fåk88, HJH85d, Hig87-32, Hig88-77, Hig88-171, LS87]. reported [Hig88-163, Hig88-91]. Reports [Hig89v, Ano83t, Ano88-154, Ano88-156, Ano88-155, Hig87z]. reprieve [Bet88]. Republic [Ano86-30]. Requirements [Sor83, FM83, Jan86, MP87]. rescue [Hig88-116]. research [Har83]. research-oriented [Har83]. resident [Ano86-39]. resource [Ano88-126, Woo83b]. respond [Hig88-43]. response [Hig87c, Hig88n, Kug86]. responsibility [Aue85, Mou84a]. rest [Ano83i].

Reston [Ano83g]. retailer [Hig88-120]. rethink [Men89d]. retina [Hig88-139]. retrieval [ALN88]. reuse [Won88b]. revealed [Hor88]. Review [Ano82e, Ano82d, Ano83e, Ano83f, Ano83h, Ano83g, Ano83i, Ano83j, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano88k, Ano88l, Ano88-128, EHH85a, HJH85a, HJH85b, HJH85c, Hig82c, Hig82b, Hig83b, Hig83c, Hig83d, Hig83e, Hig83f, Hig88-28, Hig88-91, Hig88-92, Hig88-93, Hig88-103, Hig88-94, Hig88-95, Hig88-96, Hig88-97, Hig88-99, Hig88-98, Hig88-100, Hig88-101, Hig88-102, Hig88-105, Hig88-104, Hig88-130, Hig88-124, Pre89, RFB82a, RFB82b, Woo85a]. reviews [Hig88-166]. Revised [Ano88-129]. revisited [Ano86p, Hig86a, HF86, Hig87x, Hig87j, Hig88-157, Men89a]. Richard [Hig88-41, Ano88d, Hig88-62]. Rick [Ano88-105]. rifle [Ano88-60]. rights [Ano88-81, Ano88-82, Hig88-44, Ric88]. Riley [Ano88-124, Ano88-122]. Risk [Ano86-41, Ano87-33, Hig88-106, Hof89a, Hof89b, PES88, RY86, Won85, BR83, Car83, Gar89a, Gua87, HJH85c, Hig88v, Mun88b, PHEW88, TTA^{+}88, vE85]. risks [Ano88-34, Ano88-52, Bol88c, Ric86, War83]. Rivera [Hig88-30]. road [PH87]. Robert [Hig88-101, Ano88-65, Ano88-113, Hig88m, Hig88l, Hig88-135]. Rochelle [Hig83d]. rocked [JS88]. Rodney [Ano88-51]. Roger [Hig88-52]. Rola [Hig88-118]. role [FG89, Men85a, Men89c]. Rolf [Hig82a]. Rosemary [Ano88-68]. Rosenberg [Ano84h]. Rothenberg [Hig88-48]. routing [Chi88]. RSA [Ano87u, Jun87, Jun88]. Rudiger [Ano88l]. rule [Hig88-132]. rules [Hig88-39]. rulings [Hig88-87]. Rullo [Ano84e]. Russell [HJH85a]. Ryan [Hig88-137].

S [Ano83f, Ano88-121, Ano88-145, Hig88-42, RFB82b, Mou87]. Sabotage [Ano88-130, Hig88-126]. Safe [Hig88-107, Ano88-42, Hig88-28, Hig88-47]. safeguard [Hig88-100]. Safeguards [Par84]. safely [Ano87-43, Dav86]. safes [Sop88]. said [Gar89b]. Sandra [Ano88-134, Ano88-147]. Satellite [Hig88-108]. save [Hig88-141]. say [Ano88-113, Hig88-121]. saying [Ano88-152, Ano88-153, Ano89-41, Ano89-42]. says [Ano88-105, Ano88-127, Hig88-46, Hig88-97]. Scale [She88]. Scannell [Ano88-110]. scans [Hig88-139]. Scent [Par88b]. Scheier [Ano88-65, Ano88-113]. Scheirer [Hig88-135]. scheme [Bis86, OM84]. schemes [Lon87]. Schmemann [Ano88-60]. Schmitt [Ano88-145]. Schreider [Hig88-35]. Schweitzer [Ano83g, Hig82c]. Sciences [Hig83b]. Scott [Hig88-97, Ano88h]. scramblers [Jay82]. screen [Ano88h]. Sealing [LB82]. search [Hig88-132, Hig89x]. sec [Ano83i, Ano83p, Ano83q, Ano86u, Ano87s, Hig88-162]. Sec’84 [Ano84s, Ano84t]. Sec’86 [Paa87c]. Second [Ano85v, Ano87-34, DR88, Paa88]. Secret [Hig89w, Ano84j, Ano88l, Hig88-128, Hig88-129]. Secretdisk [Hig88-169]. Secrets [HJH85b, Hig88-82]. sections [Hig88x]. Secure [Ano87-35, Ano88-131, Har88, Hig87-28, Hig88-109, Mur86b, Ano87-37, Ano88-72, Ano88-122, Bon83, BvS89, BM83, CI85, CJ86, Coh85, CB88, Har89, Hig87g, Hig87q, Hig88-158, Hig88-139, Jan86, KTT89, Mil88a, MP87, MW88, Mur86a, PH84, Sor83, Sto88]. secured [Mil85, Zaj88a]. Securicom [Ano85w, Woo84]. Securing [Cal83]. Security [Ano83q, Ano84v, Ano84-30, Ano84-32, Ano85y, Ano86-42, Ano87-34, Ano87-36, Ano87-39, Ano88-132, Ano88-134, Ano88-136, Ano88-137, Ano88-138, Ano88-135, Bra88, Eke86b, Hig83a, Hig88-110, Hig88e, Hig88-112, Hig88-113, Hig88-114, Hig88-115, Hig88-116, Hig88-111, JL89, KG86, KG87, Kru89, Mou87, Mur82, Paa87b, Ram89b, Sch88, Sha89, Thu87, VK85, Won88a, WZ86, Ala86, Ams88, Ano82d, Ano82k, Ano82q, Ano83e, Ano83x, Ano83y, Ano84e, Ano84f, Ano84-29, Ano85a, Ano85k, Ano85u, Ano85v, Ano86e, Ano86-37, Ano86-39, Ano88f, Ano88y, Ano88z, Ano88-29, Ano88-148, Ano88-45, Ano88-52, Ano88-55, Ano88-71, Ano88-65, Ano88-103, Ano88-100, Ano88-109, Ano88-113, Ano88-114, Ano88-107, Ano88-112, Ano88-133, Ano88-144, Ano88-147, Ano88-154, Ano88-156, Ano88-155, Bac83, BE89, Bea86a, Bea86b, Bec83, BBFÅ82, BvS89, Bot83, Bou88]. security [Boz88a, BS82, BS87, BP89, Buu84, CW83, Car88, CMRW88, Cha85, Cha88, Cou82, Deh89, Dos86, DJT87, EHH85b, Eke86a, Elo83, Elo85, Elo88, Fis84, For82, FM83, FM84, Guy83, Hig82c, HH82, Hig82a, Hig83c, Hig83f, Hig83i, Hig85e, Hig85f, Hig86b, Hig87a, Hig87e, Hig87d, Hig87w, Hig88k, Hig88l, Hig88-56, Hig88-27, Hig88-33, Hig88-46, Hig88-54, Hig88-57, Hig88-58, Hig88-66, Hig88-67, Hig88-70, Hig88-68, Hig88-69, Hig88-74, Hig88-76, Hig88-75, Hig88-78, Hig88-79, Hig88-167, Hig88-85, Hig88-86, Hig88-88, Hig88d, Hig88-100, Hig88-105, Hig88-123, Hig88-121, Hig88g, Hig88-126, Hig88-138, Hig88-140, Hoe86, Hof89a, Hor88, HE85, Hov86, Joh87, Kee89, KD86, Kur82, Lob83, MT88, Men85a, Men88e, Men89e, Mil88b, Mil88d, Mou84a, Mou84b, Mun88b, Mur83, NWH84, Neu82, Owe86, PB83, Paa84a, Paa84b, PH87]. security [Paa87c, Paa88, Per82, PP88, Pol84, Pos87, Ram89a, RB82, RY86, Rih87, RH86, Sch84, Sch85, Sch87, Ser85, Str88, SK88, TR86a, TR86b, Var89, Won85, Woo83a, Woo83b, Woo85b, Woo86a, Woo86b, Woo87a, Woo87b, Woo88, Woo89, Zaj88c, Zar88, vH87, vdG84, vdL86, Ano84g, Hig82b, Hig83b, Hig83d, Hig88-103, Hig88-101, Hig88-104, RFB82a, Ano83i, Ano83j, Hig88-102, Woo85a]. security-privacy [Ano84g]. see [Boz88b]. seeks [Hig88-71, Hig88-64]. seen [Hig88-34]. sees [Ano88-52]. Seev [Ano88-123]. Selecting [Men89e]. Selection [Bol88c, Elo83, Fra88, Par84]. selective [Kur82]. Senate [Ano88-29, Ano88-139]. sensible [Hig88-117]. sensitive [Cou88]. separate [Ano88h]. September [Ano88u, Ano88x, Ano88-39, Ano88-40, Ano88-41, Ano85s, Ano86-34]. sequences [Por84]. Serge [Ano88-60]. series [Sto88]. serious [Ano86-31]. servers [CB88]. service [Sha89]. Services [Ano88v, Ano88-38, Ano88w, Ano88x, Ano88-27, Ano88-44, Ano88-43, Ano88-50, Ano88-100, Ano88-98, Ano88-99, Ano88-142, Ano88-143, Ano88-144, BDE^{+}82, Bur88, Ram89a]. Sesame [Hig88c]. Settle [Ano88-101]. settlements [Sha89]. seventh [EHH85b]. several [Hig88-37]. Seymour [Hig88-92]. sharing [Ano84-30, dC85]. shell [Ano88-138]. shift [Ano88-29]. ships [Hig88-58, Hig88-94]. shirt [Ano88-102]. shirt-sleeves [Ano88-102]. shop [Hig88-139]. Should [Zaj88d, Hig88-46, Hig88-140]. Shrink [Ano88-140]. Shrink-wrap [Ano88-140]. shrug [Jon88]. Sicht [MB88]. sides [Ano88-107]. siege [Hig88-29]. Signature [dB85, Ano88-51, WCWG86]. signatures [Hig88-64, Hig88-109]. signed [Hig88-110]. Silicon [Ano88d]. Silver [Ano84w, Hig83f]. simple [Deh89]. Simulation [Koc84]. sink [Hig88-94]. site [Ano88g, Ano88-110, Ano88-89]. sites [Hig88-34, JS88]. situation [Men89b]. six [Hig83e]. skies [Hun88a]. skirts [Ano88g]. slacking [Ano88-103]. slaps [Bar88, Hig88p]. sleeves [Ano88-102]. slow [Boz88a]. smack [Hig88-84]. small [KG86, KG87]. Smart [Ano88-141, Hig88-118]. Smartcards [Fif89]. SMITE [Har89]. Smoking [Hof89b]. Smole [Hig88-68]. snake [Hig88-135]. snake-oil [Hig88-135]. snoopers [Hig88-116]. Sobol [Hig88-102]. Socha [Hig88-99]. Societal [HM86]. Society [Hig83f]. SODA [KTT89]. Software [Ano88-142, Ano88-143, Ano88-144, Fåk86, GPR^{+}88a, GPR^{+}88b, Hig88-119, Hol88, PES88, PHEW88, Per88, Ano88m, Ano88s, Ano88t, Ano88u, Ano88-34, Ano88-30, Ano88-31, Ano88-46, Ano88-56, Ano88-65, Ano88-90, Ano88-91, Ano88-100, Ano88-131, BPR88, BR89, Bum84, Erd86, Fra88, Hig87r, Hig88-39, Hig88-59, Hig88-92, Hig88-113, Hig88-111, Kim88, Mil88c, Mur86a, Mur86b, OM84, OR88, PP88, TR86a, TR86b, Won88b]. Solutions [MW88, Hig88-131]. Some [Ano86-43, Buu84, Gag88b, Hig88-51, HC88, Lun89, Plu88]. Sons [Ano82e, Ano82d, Ano83h, Ano84e, Ano84f, Ano84h, Hig82b, RFB82a, Woo85a]. sources [Hig87u]. Soviet [Bol88d]. SPA [Hig88-120]. spark [Hig88t]. Speakers [Hig88-121]. Special [Ano89-39, Hig88f, Paa84b, HJH85d]. Specification [GM85, Elo85]. Spectrum [Ano88-128]. speech [Jay82]. speed [Ano87r, Ban82]. Speedier [Ano87-37]. split [Hig88q, Hig88-71]. spots [Hig88-72]. spread [Gle89, Hig88-31, Ano88-128]. Spring [Hig83f, Hig88-28, Hig88u, Hig88-53, Hig88-68, Hig88-103, Hig88-104]. spy [Ano88-127]. St [Ano88-104]. stage [Dos86]. Stalking [Hig88-122]. stand [Ano88-107]. Standard [Ano88-44, Bur88, Ano88-43, Bet88, Hig82a, Hig87h]. Standardization [Kra84]. standards [Ano88-27, Ano88-62, BS82, Woo86a, Woo86b, Hig87d, LN88]. Stang [Hig88-113]. Star [Hig88g]. STARS [GPR^{+}88a, GPR^{+}88b]. started [Men88b]. state [Lob83, Mur83, Gri88b]. state-of-the-art [Lob83]. States [Ano83i, Hig88-87]. statistical [ALN88]. Statistics [Ano83n, Hig88-144]. status [Ano86-45, Ano86-49]. stem [Hig88-119]. step [Bec83]. Stephanie [Ano88-89]. Stephen [Ano88-47]. Steps [Mil85, Ano83f]. Stevens [Ano88-108, Ano88-138]. Stevenson [Hig88-69]. still [Per88]. Stockholm [Ano83i, Fåk87b]. stocks [Hun88c]. Stoll [Hig88-122]. Stone [Ano88-121]. storage [Ano87r, Jan86, MP87]. storage/backup [Ano87r]. storm [Ano88-104, Ko88]. story [Hig83e]. strategic [Sny88]. Strategies [Car87]. strategy [Buu84, Per82, HJH85c]. street [CM88]. strict [Ano88-71]. strike [Bri88]. strikes [Hig88-141]. Structing [Hig88-123]. structure [Ano84i, Ano88-145]. structures [TTA^{+}88]. structuring [Str88]. Stuart [Hig88-93, Hig88-96]. Study [Ano89-39, Buc82, Coh87e, DR88]. Subject [Ano82r, Ano83z, Ano84-31, Ano85x, Ano86-44, Ano87-38, Ano88-146, Ano89-40]. subordinates [Bol89a]. success [Woo87b]. successful [Hig88-62]. sudden [Hig87f]. suggested [Ano86-43]. suit [Hig88q, Hig88m, Hig88p, Hig88-43, Hig88-71, Jon88, Ric88]. supervise [Mou87]. supervisor [PH86]. supplies [Ano88-149]. support [BS87, Car83, Ten86b, Ten86a]. Supportability [PES88, PHEW88]. suppresses [Hig88-77]. surfaces [Hig88-63]. Surreptitious [PB83]. surveillance [Ano88-32]. survey [JO89a, JO89b, Kie88, PvL86, RH86, Web89a]. Susan [Ano88-112]. suspected [Hig87-32]. Sweden [Ano83i, Fåk88, LB82]. switch [Sor83]. symphony [Hig88-133]. synthesis [Web89a]. System [Ano87-39, Hov83, Hov86, ALN87, Ano86e, Ano88-145, BPR88, BB88, Bol88b, BS87, Eke86a, Elo85, Gri88b, Hig83b, Hig86b, Hig88-161, Hig88-69, Hig88-79, Hig88-99, Hig88-105, Hig88-131, HM86, Hor88, KTT89, LK84, Men85b, Men88e, Men89e, Mou84b, PB83, Pos87, RY86, TC88, Voi88, Woo83a, WZ86, Woo87a, dC85, vH87, vdG84, Hig82a, Ano83h]. system/88 [Ano88-145]. systematic [Cou82]. Systems [Ano85y, GPR^{+}88a, GPR^{+}88b, Hig88-94, LN88, Ano82d, Ano84-30, Ano86s, Ano86-38, Ano88-148, Ano88-53, Ano88-51, Ano88-104, Ano88-134, Ano88-137, Ano88-138, Ano89-39, BB84, BBFÅ82, Bol88c, Bon83, BP89, CW83, Cer85, CM88, DJT87, FM84, Gua87, Hig88-28, Hig88-29, Hig88-78, Hig88-114, Hig88-116, HE85, Joh87, Koc84, Kru89, Lon87, Mal88, MT88, PH84, PP88, PR85b, Plu88, RFB82b, Ram89a, Ten86b, Ten86a, Thu87, Web89a, Wis89, Woo86a, Woo86b, Woo87a, Woo87b, Woo88, Yeo85].

T. [Hig88-114]. tab [Hig88-175]. tackle [Mun88b]. Tactical [Ano88-147]. tactics [Hig88-33]. Taft [Ano88-34]. take [Mil85]. Talbot [Hig82b, RFB82a]. talking [Zaj88d]. Taming [Hig88-124]. Tangible [Hig88-125]. Tari [Hig88-35]. Tate [Hig88-106]. TC [Ano88-66, Ano83s, Ano84-32, Ano86-45, Ano87-32, HE85]. TC-11 [Ano83s, Ano84-32]. TCP [Ram89b]. TCP/IP [Ram89b]. Teaching [Har83, Neu82]. Technical [Ano88-78, Ano89n, Ano85m, Ano88-74, Ano88-75, Ano88-76, Ano88-77, Ano88-79, Ano89o, Ano89p, Ano89q, Ano89r, Ano89s, Ano89t, Ano89u, GPR^{+}88b, Hov86, Woo86a, Woo86b, vBH88, vdG84]. Techniques [Ano88-39, Ano88-40, Ano88-41, Hig83a, Hig85f]. technological [Ala86]. Technology [Ano83t, Ano86-46, Ano86-47, Ano87-40, Ano87-41, GPR^{+}88a, GPR^{+}88b, Hig87-29, Hig88-28, Hig88-130, Hig88-126, Ano86-28, Ano88-80, Hig88-52, MB88, Mur83, Ano84v]. telecommunication [Hig88-101]. Telecommunications [Hig88-102, NWH84, Ano84h]. teleinformatics [GL82]. TeleTrusT [Rih87]. TeleTrusT-OSIS [Rih87]. telex [FVTS87]. tells [Mun88a]. Tempest [Hig87-30, Mun88a, Ano88-139, Hig88-127]. ten [RB82]. terminal [Hig88-164, Sor83]. terminalized [Eke86a]. terminals [Ano86-37, Jan86]. Terminology [Ano84g]. Terrell [Hig88-72]. terrorism [Blo82, Hig88-170, Mad88, Men83]. Test [PHEW88, Hig88b, Koc83]. Testing [LN88, PPT89]. text [WC88]. Their [Hig88-50, Ano89-39, Bol89a]. them [Bri88, Zaj88d]. themselves [Hig88-56]. Theory [Coh87b, Ano87w, Bot83, Gag88a, Gag88b, Gle89, Hof89a, Sal88]. there [Fåk86]. thief [Ano88-148]. third [Hig88-106]. Thomas [Ano84e, Ano88-122, Hig88-28]. those [Hig88g, Per88]. threat [Ano86-39, HP84, Hig88k, Hig88-92, Hig88-119, bAG88, Ano88l]. threats [LAN86]. Three [Ano83-27]. thwart [Hig88s, Hig88-130]. ties [Ano88-121]. tigers [Her88]. Tim [Ano88-90, Hig88-112]. time [Ano86y, Ano88-151, Hig88-82, Mal88, Men89d, Voi88]. Times [Hig88-91, Hig88-129]. timesharing [RMS89]. timing [Mal88]. Tips [Hig88-170]. today [Fåk86, bAG88]. together [Ano88-114]. tokens [Spe87]. told [Gof88b]. tolerance [Ano88-55, Hig88-40]. tolerant [Ano88-54, Ano88-145, Voi88]. Tom [Ano83f, Ano88-124]. tool [Ano86t, BR83, Hig88-118]. tools [Ano86-28, Ano88-131, Hig87p]. Top [Hig88-128, Hig88-129, Bou88]. topics [vdBL84]. topology [PH84]. total [Hig83i]. Tough [Hig88-130]. trading [Hig88-98]. trails [Ano86z, Boz88b]. training [CMRW88, Mun88a]. Transaction [Mil85, PP88, Plu88]. transactions [Ano88-55, Eke86b]. transborder [Blo82, EHH85a]. transfer [KD86, Ram89b, RY86, Sha89, Web89a]. transferability [Ano88-31]. transfers [vH87]. transmission [Ano88-42, Hig87-28]. transnational [RFB82b]. transparent [Hig88-169]. transport [VK85]. TRASEC [vH87]. Treating [Dav88]. trespassing [Ano86h]. tricky [Ano88-51]. Trojan [LAN86]. trouble [Hig88-141, Hun88a]. Troy [Ano88-64]. trust [Hig88u]. tug [Ano88-47]. tug-of-war [Ano88-47]. tunes [Hig88-133]. Turn [Hig83b, Hig83c]. turning [Ano88m]. Tutorial [NWH84, Wil82c, Hig83f]. Twenty [Bol89a]. Twenty-two [Bol89a]. twin [Paa84a]. Two [CB88, Bol89a, Hig88q, Hig88-71]. Typical [Hig88-131].

U.S [Ano83n, Hig88-91]. U.S. [HJH85d, Hig88-171, Men88e]. U.S.A. [Sto88]. UFS [Hig88-104]. unanswered [Lun89]. Unclassified [Mil88b, Cou88]. uncovers [Ano88-67]. understand [McD88]. Understanding [HR88, Men88c, Men88d]. unfair [Hig88r]. unforgettable [CM88]. unfriendly [Hun88a]. Uninterruptible [Ano88-149]. Unique [Bac83]. Unisys [Bar88]. unite [Hig88x]. United [Ano83i, Hig88-59]. units [vE85]. university [RMS89, Web89b]. UNIX [Hig88-74, Hig88-105]. unobservability [BP89]. Unprotected [Hig88-133, Ano88a]. Upcoming [Hig88h]. update [Hig88-173, Hig88-174]. updating [Joh87]. upper [Ano88-133]. UPS [Hig88-141]. urges [Ano88h]. USA [Hig83b]. Use [DR88, Bon83, Dav86, Mal88, Men88d, RMS89]. used [Ben88, KG87]. useful [BR83, Car87]. User [BB84, Fåk87b, Hig88-104, ALN87, Ano86-28, Ano88-32, FG89, Hig88-93, Hig88-116, Hig88-135, HR88, Mil88c, MB88, OM84, PH86, PW87a, PES88]. user-controlled [OM84]. User-friendly [BB84, Hig88-104]. Users [Boz88b, Ano88-112, Ano88-126, Gof88b, Hig88-140, Smi87, Spe87, Ano88-128]. using [All88, Ano87u, Ayo83, Koc83, Kol88, PR85a, Ten86b, dB85].

V [Hig88l]. VA [Ano83g]. vaccination [Hig88-62]. vaccine [Hig88-48]. vaccines [Hig88-119]. vague [Hig88m]. Valley [Ano88d]. value [Hig85d]. Vanessa [Ano88-139]. VCHECKER [Hig89x]. vendors [Ano88-136]. Verification [Var89, Ano88-51, Ano88-131, WCWG86, dB85]. Verifying [ALN87]. versus [Fåk86, Koc84]. Very [Ano87-42, She88]. vi [Hig82c, Hig83b, Hig83f]. victims [Hig89v]. video [Ano88-132, Hig88-164, vE85]. view [Ano86f, Ano88-35, Sch85, Sch87, Thu89]. viewpoint [vH84]. viii [HJH85b, Hig83d]. Viiveke [Ano83i]. violation [Car88, PB83]. viral [Dav88, Hig88o]. Virginia [Ano88-96]. virtual [Kol88]. Virus [Ano88-150, Hig88-136, Hig88i, Hig88-172, Hig88-173, Hig88-174, Hig88-135, Ano88-129, Coh88d, Fåk88, Gla88, Hig87i, Hig88-143, Hig88a, Hig88-145, Hig88-146, Hig88-38, Hig88-41, Hig88-159, Hig88-160, Hig88b, Hig88-163, Hig88-62, Hig88-63, Hig88-64, Hig88-165, Hig88-97, Hig88f, Hig88-171, Hig88-144, Hig88-134, Hig89g, Hig89x, Men89b, Rad89, Web89b, vW89]. Virus-plagued [Hig88-135]. VIRUSCAN [Pre89]. Viruses [DeM89, Hig88-137, Whi89, Zaj88d, Ano88l, Ano88-128, Ano89-39, Bri88, Coh87b, Coh88c, Coh89a, Coh89b, Gle89, Hig87f, Hig88-30, Hig88-148, Hig88-31, Hig88-50, Hig88-80, Hig88-112, Hig88-119, Mur88, PG87, Wis89, Zaj89b, Hig88-92]. vital [Hig88-130]. Vizachero [Ano88-105]. VLSI [She88]. VM [Ano88-135, Hig88-138]. VMSecure [Hig88-138]. Voelcher [Ano88-128]. Voice [Hig88-139, Ano88-33, Hig88-126]. Volume [Ano84e, Ano84f, Ano82c, Ano82r, Ano83d, Ano83z, Ano84d, Ano84-31, Ano85f, Ano85x, Ano86g, Ano86-44, Ano87g, Ano87-38, Ano88i, Ano88-146, Ano88-81, Ano88-82]. Volumes [PHEW88]. VT [Ano83j]. vulnerabilities [Kee89, Sch84]. vulnerability [HM86]. vulnerable [Fåk88, Hig88-129].

W [Hig83f, Hig88-94, Woo85a]. Wall [Ano88-104]. Wandel [MB88]. Wanted [Ano86-48, Ano88-151]. war [Ano88-47]. warned [Hig88-135]. warning [Hig88-175]. wars [Hig88g]. was [Gar89b]. Washington [Ano83e, Hig83c, EHH85b, Hig88-63]. wasn’t [Gla88]. watch [Ano86-46, Ano86-47, Ano87-40, Ano87-41, Hig87-29]. wave [Ano88-112]. way [Ano88-113, Deh89]. Wayne [Hig88-99, Ano88-52]. Wayner [Hig88-89]. weak [Hig88-72, Hig88-79, HC88]. Weaknesses [JO89a, JO89b, She88]. weathers [Ko88]. Webb [Ano83f]. WEEK [Hig88-115, Hig88-50, Hig88-135, CM88]. weekly [Hig88-32]. Weiss [Ano88-126]. Welchman [Hig83e]. welcome [Hig88c]. Wellesley [Hig83b]. WG [Fåk87b, Kug86, Ano86-49]. Where [Blo82, Hig88g]. while [Ano88-104]. white [Bol88d]. white-collar [Bol88d]. who [Hig88-130]. wide [Ano88-154, Ano88-156, Ano88-155, Bur88]. Wiley [Ano82e, Ano82d, Ano83h, Ano84e, Ano84f, Ano84h, Hig82b, RFB82a, Woo85a]. Wilkinson [Ano88-89]. will [Mun88b]. William [Ano82d, Ano88-55, Ano88-135, Hig88-46]. Wilson [Ano83j]. Wily [Hig88-122]. within [bAG88]. without [Hig88b, PW87a]. Wofsey [Ano84f]. word [Ben88, Cal83, Smi87]. WordPerfect [Ben88]. work [Hig88-48, Mil88a]. workers [Ano88-127]. Working [Ano84-32, Ano87-32, Fåk87b]. workplace [Ano83g, Ano88-108]. works [Kur82]. worksheet [BR83]. worksite [Kie88]. Workstation [Req88, Hig88-86, Hig88-103, Sch84, Vil88]. Workstation/mainframe [Req88]. World [Ano88-154, Ano88-156, Ano88-155, Ano83i, Mad88]. World-wide [Ano88-154, Ano88-156, Ano88-155]. worm [Gar89b]. worms [Whi89]. worries [Ano88-128]. worse [Hig88-50]. Would [Hig87-32, Hig88-45]. wrap [Ano88-140]. Write [Ano87-43, Hig88-175]. Write-protect [Hig88-175].

x [HJH85a]. xi [Ano83f, Ano84e, Hig82b, RFB82a]. xii [RFB82b]. xiii [Ano82d, Ano84i]. xiv [Ano82e, Ano83e, Ano83g, Ano83j, EHH85a]. xix [Ano84f, Woo85a]. xvi [HJH85c]. xxii [Ano83h]. xxxvi [Ano83i].

years [RB82, Web88]. Yeheskel [Ano88-123]. York [Ano82e, Ano82d, Ano83h, Ano84e, Ano84f, Ano84h, Ano84i, HJH85b, Hig82b, Hig82a, Hig83e, Hig88-91, Hig88-129, RFB82a, RFB82b, Woo85a]. Yvo [Ano88-62].

Zajac [Ano88k].

References

Alagar:1986:HAT

[Ala86] V. S. Alagar. A human approach to the technological challenges in data security. Computers & Security, 5(4):328–335, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

           (electronic). URL https://www.sciencedirect.com/science/
           article/pii/0167404886900544.

Allen:1988:DAR

[All88] Bruce M. Allen. Data acquisition, reduction, and analysis using a microcomputer. Computers & Security, 7(5):513, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

           (electronic). URL https://www.sciencedirect.com/science/
           article/pii/0167404888902301.

Ahituv:1987:VAI

[ALN87] Niv Ahituv, Yeheskel Lapid, and Seev Neumann. Verifying the authentication of an information system user. Computers & Security, 6(2):152–157, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900861.

Ahituv:1988:PSD

[ALN88] Niv Ahituv, Yeheskel Lapid, and Seev Neumann. Protecting statistical databases against retrieval of private information. Computers & Security, 7(1):59–63, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905044.

Amsel:1988:NSA

[Ams88] Ellen Amsel. Network security and access controls. Computers & Security, 7(1):53–57, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905032.

Anonymous:1982:ARAa

[Ano82a] Anonymous. Abstracts of recent articles and literature. Computers & Security, 1(2):193–199, June 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404882900153.

Anonymous:1982:ARAb

[Ano82b] Anonymous. Abstracts of recent articles and literature. Computers & Security, 1(3):293–297, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900505.

Anonymous:1982:AIV

[Ano82c] Anonymous. Author index to volume 1. Computers & Security, 1(3):303–304, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404882900530.

Anonymous:1982:BRCb

[Ano82d] Anonymous. Book review: Computer control and security: a guide for managers and systems analysts: William E. Perry: New York:

John Wiley and Sons, 1981. xiii + 207 pp. $22.95. Computers & Security, 1(1):85, January 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404882900311.

Anonymous:1982:BRCa

[Ano82e] Anonymous. Book review: Cryptography: a primer: Alan G. Konheim: New York: John Wiley and Sons, 1981. xiv + 432 pp. $34.95. Computers & Security, 1(1):84, January 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488290030X.

Anonymous:1982:Ca

[Ano82f] Anonymous. Calendar. Computers & Security, 1(1):92–94, January 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404882900335.

Anonymous:1982:Cb

[Ano82g] Anonymous. Calendar. Computers & Security, 1(2):200, June 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404882900165.

Anonymous:1982:Cc

[Ano82h] Anonymous. Calendar. Computers & Security, 1(3):298, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404882900517.

Anonymous:1982:CLA

[Ano82i] Anonymous. Current literature: Abstracts of recent articles and literature. Computers & Security, 1(1):86–91, January 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900323.

Anonymous:1982:EB

[Ano82j] Anonymous. Editorial Board. Computers & Security, 1(1):ifc, January 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404882900177.

Anonymous:1982:EGC

[Ano82k] Anonymous. Executive guide to computer security. Computers & Security, 1(3):233–241, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404882900396.

Anonymous:1982:IBE

[Ano82l] Anonymous. International board of editors. Computers & Security, 1(1):3–8, January 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404882900190.

Anonymous:1982:NMI

[Ano82m] Anonymous. New members of the international board of editors. Computers & Security, 1(2):97–98, June 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900025.

Anonymous:1982:PJa

[Ano82n] Anonymous. Pages 1–94 (January 1982). Computers & Security, 1 (1):??, January 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1982:PN

[Ano82o] Anonymous. Pages 201–304 (November 1982). Computers & Security, 1(3):??, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1982:PJb

[Ano82p] Anonymous. Pages 95–200 (June 1982). Computers & Security, 1(2):??, June 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1982:POC

[Ano82q] Anonymous. Professional organizations in computer security. Computers & Security, 1(2):187–190, June 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900116.

Anonymous:1982:SIV

[Ano82r] Anonymous. Subject index to volume 1. Computers & Security, 1(3):299–302, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404882900529.

Anonymous:1983:ARAa

[Ano83a] Anonymous. Abstracts of recent articles and literature. Computers & Security, 2(1):88–96, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900457.

Anonymous:1983:ARAb

[Ano83b] Anonymous. Abstracts of recent articles and literature. Computers & Security, 2(2):194–205, June 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900597.

Anonymous:1983:ARAc

[Ano83c] Anonymous. Abstracts of recent articles and literature. Computers & Security, 2(3):286–296, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900159.

Anonymous:1983:AIV

[Ano83d] Anonymous. Author index to volume 2. Computers & Security, 2(3):309, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488390024X.

Anonymous:1983:BRC

[Ano83e] Anonymous. Book review: Computer security: an overview of national concern and challenge: Louise G. Becker: Washington DC: Library of Congress, 1983. xiv + 241 pp. Computers & Security, 2(3):299–300, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900202.

Anonymous:1983:BRN

[Ano83f] Anonymous. Book review: Nine steps to effective EDP loss control: Tom S. Eason and Douglass A. Webb: Bedford, MA: Digital Press, Digital Equipment Company, 1982. xi + 177 pp. Computers & Security, 2(3):299, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900196.

Anonymous:1983:BRPb

[Ano83g] Anonymous. Book review: Protecting information in the electronic workplace: a guide for managers: James A. Schweitzer: Reston VA: Reston Publishing Company, Inc., 1983. xiv + 146 pp. $19.95 (hardbound) and $14.95 (paperback). Computers & Security, 2(3): 299, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404883900184.

Anonymous:1983:BRPa

[Ano83h] Anonymous. Book review: Protecting your microcomputer system: Harold Joseph Highland: New York: John Wiley and Sons, Inc., 1984. xxii + 244 pp. $14.95. Computers & Security, 2(3): 298, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404883900172.

Anonymous:1983:BRSa

[Ano83i] Anonymous. Book review: Security, IFIP/sec ’83: Viiveke Fåk, Editor: Proceedings of the first conference, Stockholm, Sweden, 16–19 May 1983. Amsterdam: North-Holland, 1983. xxxvi + 328 pp.$39.00 (in United States and Canada). Dfl. 95.00 (rest of the world). Computers & Security, 2(3):297–298, November 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900160.

Anonymous:1983:BRSb

[Ano83j] Anonymous. Book review: Security manual: Eric Oliver and John Wilson: [Fourth edition]. Brookfield VT: Gower Publishing Co., 1983. xiv + 228 pp. $12.50. Computers & Security, 2(3): 300, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404883900214.

Anonymous:1983:Ca

[Ano83k] Anonymous. Calendar. Computers & Security, 2(1):97–98, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404883900469.

Anonymous:1983:Cb

[Ano83l] Anonymous. Calendar. Computers & Security, 2(2):206–211, June 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404883900603.

Anonymous:1983:Cc

[Ano83m] Anonymous. Calendar. Computers & Security, 2(3):301–304, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404883900226.

Anonymous:1983:CCB

[Ano83n] Anonymous. Computer crime: Bureau of Justice Statistics, U.S. Department of Justice. Computers & Security, 2(2):186–188, June 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404883900573.

Anonymous:1983:EB

[Ano83o] Anonymous. Editorial Board. Computers & Security, 2(1):ifc, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404883900251.

Anonymous:1983:IS

[Ano83p] Anonymous. IFIP/Sec ’83. Computers & Security, 2(2):99–101, June 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404883900470.

Anonymous:1983:ISI

[Ano83q] Anonymous. IFIP/Sec ’83 — IFIP’s first security conference. Computers & Security, 2(2):120–131, June 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900500.

Anonymous:1983:NMI

[Ano83r] Anonymous. New members of the international board of Editors. Computers & Security, 2(3):214–217, November 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900020.

Anonymous:1983:NIT

[Ano83s] Anonymous. News from IFIP TC-11. Computers & Security, 2(3):281–282, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900135.

Anonymous:1983:OTA

[Ano83t] Anonymous. The Office of technology assessment reports. Computers & Security, 2(1):85–87, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900445.

Anonymous:1983:PJa

[Ano83u] Anonymous. Pages 1–98 (January 1983). Computers & Security, 2 (1):??, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1983:PN

[Ano83v] Anonymous. Pages 213–309 (November 1983). Computers & Security, 2(3):??, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1983:PJb

[Ano83w] Anonymous. Pages 99–211 (June 1983). Computers & Security, 2(2):??, June 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1983:POCa

[Ano83x] Anonymous. Professional organizations in computer security. Computers & Security, 2(2):184–185, June 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900561.

Anonymous:1983:POCb

[Ano83y] Anonymous. Professional organizations in computer security. Computers & Security, 2(3):283–285, November 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900147.

Anonymous:1983:SIV

[Ano83z] Anonymous. Subject index to volume 2. Computers & Security, 2(3):305–308, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900238.

Anonymous:1983:TNM

[Ano83-27] Anonymous. Three new members of the international board of Editors. Computers & Security, 2(1):3–4, January 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900275.

Anonymous:1984:ARAa

[Ano84a] Anonymous. Abstracts of recent articles and literature. Computers & Security, 3(1):61–72, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900361.

Anonymous:1984:ARAb

[Ano84b] Anonymous. Abstracts of recent articles and literature. Computers & Security, 3(2):151–163, May 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900609.

Anonymous:1984:ARAc

[Ano84c] Anonymous. Abstracts of recent articles and literature. Computers & Security, 3(4):328–341, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900142.

Anonymous:1984:AIV

[Ano84d] Anonymous. Author index to volume 3. Computers & Security, 3(4):350–351, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900178.

Anonymous:1984:BRAa

[Ano84e] Anonymous. Book review: Advances in computer security management. Volume 1: Thomas A. Rullo, editor: New York, (Heyden) John Wiley and Sons, 1980. 245 + xi pages. Computers & Security, 3(1):58, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900324.

Anonymous:1984:BRAb

[Ano84f] Anonymous. Book review: Advances in computer security management. Volume 2: Marvin M. Wofsey, editor: New York: John Wiley and Sons, 1983. 268 + xix pages. $39.95. Computers & Security, 3(1):58–59, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900336.

Anonymous:1984:BRDa

[Ano84g] Anonymous. Book review: Data protection-data security-privacy: Luxembourg: Terminology Bureau of the Commission of the European Countries, 1981. 776 pages. Computers & Security, 3(1): 57–58, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900312.

Anonymous:1984:BRDb

[Ano84h] Anonymous. Book review: Dictionary of computers, data processing and telecommunications: Jerry M. Rosenberg: New York: John Wiley and Sons, 1984. 614 + ix pages. $29.95 hardcover, $14.95 paperback. Computers & Security, 3(1):60, February 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488490035X.

Anonymous:1984:BRI

[Ano84i] Anonymous. Book review: Information integrity: a structure for its definition and management: Hal B. Becker: New York: Mc-Graw Hill Book Co., 1983. 237 + xiii pages. $26.95. Computers & Security, 3(1):59–60, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900348.

Anonymous:1984:BRP

[Ano84j] Anonymous. Book review: The puzzle palace: a report on NSA, America’s most secret agency: James Bamford: Boston: Houghton Mifflin Company, 1982, 465 pages. $16.95. Computers & Security, 3 (1):57, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900300.

Anonymous:1984:Ca

[Ano84k] Anonymous. Calendar. Computers & Security, 3(1):73–75, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404884900373.

Anonymous:1984:Cb

[Ano84l] Anonymous. Calendar. Computers & Security, 3(2):165–167, May 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404884900610.

Anonymous:1984:Cc

[Ano84m] Anonymous. Calendar. Computers & Security, 3(3):249–251, August 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404884900476.

Anonymous:1984:Ce

[Ano84n] Anonymous. Calendar. Computers & Security, 3(4):342–345, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900154.

Anonymous:1984:CLD

[Ano84o] Anonymous. Computer law decisions and opinions. Computers & Security, 3(3):237–248, August 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900464.

Anonymous:1984:Cd

[Ano84p] Anonymous. Crypto ’84. Computers & Security, 3(4):324–325, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900129.

Anonymous:1984:E

[Ano84q] Anonymous. Editorial. Computers & Security, 3(1):1–2, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404884900191.

Anonymous:1984:EB

[Ano84r] Anonymous. Editorial Board. Computers & Security, 3(1):ifc, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488490018X.

Anonymous:1984:ISa

[Ano84s] Anonymous. IFIP/Sec’84. Computers & Security, 3(1):49–50, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900270.

Anonymous:1984:ISb

[Ano84t] Anonymous. IFIP/Sec’84. Computers & Security, 3(2):143–144, May 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404884900580.

Anonymous:1984:M

[Ano84u] Anonymous. In memoriam. Computers & Security, 3(4):253– 255, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900014.

Anonymous:1984:ICC

[Ano84v] Anonymous. International Carnahan Conference on Security Technology. Computers & Security, 3(2):135–139, May 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900555.

Anonymous:1984:ISJ

[Ano84w] Anonymous. The IPSA Silver Jubilee conference. Computers & Security, 3(2):139–140, May 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900567.

Anonymous:1984:NMI

[Ano84x] Anonymous. New members of the international board of Editors. Computers & Security, 3(2):80, May 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488490049X.

Anonymous:1984:PF

[Ano84y] Anonymous. Pages 1–75 (February 1984). Computers & Security, 3 (1):??, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1984:PA

[Ano84z] Anonymous. Pages 169–251 (August 1984). Computers & Security, 3(3):??, August 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1984:PN

[Ano84-27] Anonymous. Pages 253–351 (November 1984). Computers & Security, 3(4):??, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1984:PM

[Ano84-28] Anonymous. Pages 77–167 (May 1984). Computers & Security, 3(2):??, May 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1984:POC

[Ano84-29] Anonymous. Professional organizations in computer security. Computers & Security, 3(2):145–150, May 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900592.

Anonymous:1984:SDD

[Ano84-30] Anonymous. Security in distributed data sharing systems. Computers & Security, 3(4):325–327, November 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900130.

Anonymous:1984:SIV

[Ano84-31] Anonymous. Subject index to volume 3. Computers & Security, 3(4):346–349, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900166.

Anonymous:1984:TWC

[Ano84-32] Anonymous. TC-11 working conference on informatics security management. Computers & Security, 3(1):50–52, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404884900282.

Anonymous:1985:DNC

[Ano85a] Anonymous. 7th DoD/NBS computer security conference. Computers & Security, 4(3):229–237, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900331.

Anonymous:1985:ARAc

[Ano85b] Anonymous. Abstract of recent articles and literature. Computers & Security, 4(3):251–263, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900434.

Anonymous:1985:ARAa

[Ano85c] Anonymous. Abstracts of recent articles and literature. Computers & Security, 4(1):83–95, March 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900112.

Anonymous:1985:ARAb

[Ano85d] Anonymous. Abstracts of recent articles and literature. Computers & Security, 4(2):175–181, June 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900240.

Anonymous:1985:ARAd

[Ano85e] Anonymous. Abstracts of recent articles and literature. Computers & Security, 4(4):349–362, December 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900537.

Anonymous:1985:AIV

[Ano85f] Anonymous. Author index to volume 4. Computers & Security, 4(4):367, December 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900562.

Anonymous:1985:Ca

[Ano85g] Anonymous. Calendar. Computers & Security, 4(1):97–99, March 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404885900124.

Anonymous:1985:Cb

[Ano85h] Anonymous. Calendar. Computers & Security, 4(2):183–184, June 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404885900252.

Anonymous:1985:Cc

[Ano85i] Anonymous. Calendar. Computers & Security, 4(3):265–266, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404885900446.

Anonymous:1985:Cd

[Ano85j] Anonymous. Calendar. Computers & Security, 4(4):363–364, December 1985. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404885900549.

Anonymous:1985:DSC

[Ano85k] Anonymous. Data security and confidentiality. Computers & Security, 4(1):75–81, March 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900100.

Anonymous:1985:EB

[Ano85l] Anonymous. Editorial Board. Computers & Security, 4(1): ifc, March 1985. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488590001X.

Anonymous:1985:ITC

[Ano85m] Anonymous. IFIP technical committee 11 members. Computers & Security, 4(2):153–159, June 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900215.

Anonymous:1985:N

[Ano85n] Anonymous. Networks ’84. Computers & Security, 4(3):242–243, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404885900355.

Anonymous:1985:NMIa

[Ano85o] Anonymous. New member of the international board of Editors. Computers & Security, 4(1):3, March 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900033.

Anonymous:1985:NMIb

[Ano85p] Anonymous. New members of the international board of editors. Computers & Security, 4(3):187, September 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900276.

Anonymous:1985:PM

[Ano85q] Anonymous. Pages 1–100 (March 1985). Computers & Security, 4 (1):??, March 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1985:PJ

[Ano85r] Anonymous. Pages 101–184 (June 1985). Computers & Security, 4(2):??, June 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1985:PS

[Ano85s] Anonymous. Pages 185–266 (September 1985). Computers & Security, 4(3):??, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1985:PD

[Ano85t] Anonymous. Pages 267–367 (December 1985). Computers & Security, 4(4):??, December 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1985:POC

[Ano85u] Anonymous. Professional organizations in computer security. Computers & Security, 4(2):173–174, June 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900239.

Anonymous:1985:SIC

[Ano85v] Anonymous. Second international congress on computer security. Computers & Security, 4(2):161–171, June 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900227.

Anonymous:1985:S

[Ano85w] Anonymous. Securicom ’85. Computers & Security, 4(3):238– 242, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404885900343.

Anonymous:1985:SIV

[Ano85x] Anonymous. Subject index to volume 4. Computers & Security, 4(4):365–366, December 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900550.

Anonymous:1985:SS

[Ano85y] Anonymous. Systems security. Computers & Security, 4(4):343– 347, December 1985. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404885900525.

Anonymous:1986:ARAa

[Ano86a] Anonymous. Abstracts of recent articles and literature. Computers & Security, 5(1):55–67, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901185.

Anonymous:1986:ARAb

[Ano86b] Anonymous. Abstracts of recent articles and literature. Computers & Security, 5(2):150–164, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901392.

Anonymous:1986:ARAc

[Ano86c] Anonymous. Abstracts of recent articles and literature. Computers & Security, 5(3):257–272, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886900179.

Anonymous:1986:ARAd

[Ano86d] Anonymous. Abstracts of recent articles and literature. Computers & Security, 5(4):350–359, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886900593.

Anonymous:1986:AMS

[Ano86e] Anonymous. Advanced microcomputer security system. Computers & Security, 5(4):285–287, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886900453.

Anonymous:1986:AVA

[Ano86f] Anonymous. Another view about passwords. Computers & Security, 5(4):279–281, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886900428.

Anonymous:1986:AIV

[Ano86g] Anonymous. Author index to volume 5. Computers & Security, 5(4):367–368, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886900647.

Anonymous:1986:BDT

[Ano86h] Anonymous. “Back door” trespassing. Computers & Security, 5 (2):94, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901264.

Anonymous:1986:Ca

[Ano86i] Anonymous. Calendar. Computers & Security, 5(1):80–82, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404886901227.

Anonymous:1986:Cc

[Ano86j] Anonymous. Calendar. Computers & Security, 5(2):177–178, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404886901586.

Anonymous:1986:Cd

[Ano86k] Anonymous. Calendar. Computers & Security, 5(3):273–275, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886900180.

Anonymous:1986:Ce

[Ano86l] Anonymous. Calendar. Computers & Security, 5(4):363–364, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886900623.

Anonymous:1986:CBa

[Ano86m] Anonymous. Coffee break. Computers & Security, 5(3):190, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886900088.

Anonymous:1986:Cb

[Ano86n] Anonymous. CRYPTO ’85. Computers & Security, 5(2):170–176, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404886901574.

Anonymous:1986:D

[Ano86o] Anonymous. The demise of the DES. Computers & Security, 5 (1):3–4, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901057.

Anonymous:1986:R

[Ano86p] Anonymous. The DES revisited. Computers & Security, 5(4):281– 282, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488690043X.

Anonymous:1986:DOD

[Ano86q] Anonymous. Don’t overlook data compression. Computers & Security, 5(1):8–9, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901100.

Anonymous:1986:EB

[Ano86r] Anonymous. Editorial Board. Computers & Security, 5(1): ifc, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488690101X.

Anonymous:1986:ERS

[Ano86s] Anonymous. Electronic record systems and individual privacy. Computers & Security, 5(4):361–362, December 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886900611.

Anonymous:1986:HDD

[Ano86t] Anonymous. Hard disk disaster recovery tool. Computers & Security, 5(3):190–192, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488690009X.

Anonymous:1986:IS

[Ano86u] Anonymous. IFIP/Sec ’86. Computers & Security, 5(1):6–7, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404886901070.

Anonymous:1986:BL

[Ano86v] Anonymous. In brief . . .. Computers & Security, 5(3):187–188, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886900052.

Anonymous:1986:III

[Ano86w] Anonymous. International Information Integrity Institute. Computers & Security, 5(4):287, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886900465.

Anonymous:1986:LLN

[Ano86x] Anonymous. The lure of large numbers. Computers & Security, 5(4):287, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886900477.

Anonymous:1986:MAC

[Ano86y] Anonymous. Microcomputer access control limited by password, time, day, and/or program. Computers & Security, 5(3):185–186, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886900040.

Anonymous:1986:MAT

[Ano86z] Anonymous. Microcomputer audit trails made easy. Computers & Security, 5(3):188–189, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886900064.

Anonymous:1986:MMC

[Ano86-27] Anonymous. Microcomputer-mainframe communications link. Computers & Security, 5(1):8, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://

www.sciencedirect.com/science/article/pii/0167404886901094.

Anonymous:1986:MTT

[Ano86-28] Anonymous. Modern technology tools for user authentication. Computers & Security, 5(3):184–185, September 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886900039.

Anonymous:1986:NMI

[Ano86-29] Anonymous. New member of the international board of editors. Computers & Security, 5(1):2, March 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886901033.

Anonymous:1986:NPR

[Ano86-30] Anonymous. News from the peoples Republic of China. Computers & Security, 5(1):3, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901045.

Anonymous:1986:SN

[Ano86-31] Anonymous. On a serious (?) note. Computers & Security, 5 (2):99, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901318.

Anonymous:1986:CPC

[Ano86-32] Anonymous. On cryptographic protection of capabilities. Computers & Security, 5(2):98–99, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901306.

Anonymous:1986:PM

[Ano86-33] Anonymous. Pages 1–82 (March 1986). Computers & Security, 5 (1):??, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1986:PS

[Ano86-34] Anonymous. Pages 179–275 (September 1986). Computers & Security, 5(3):??, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1986:PD

[Ano86-35] Anonymous. Pages 277–368 (December 1986). Computers & Security, 5(4):??, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1986:PJ

[Ano86-36] Anonymous. Pages 83–178 (June 1986). Computers & Security, 5(2):??, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1986:PSM

[Ano86-37] Anonymous. Physical security of microcomputers and terminals. Computers & Security, 5(1):7–8, March 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886901082.

Anonymous:1986:PIE

[Ano86-38] Anonymous. Protection of information in electronic systems. Computers & Security, 5(4):360, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488690060X.

Anonymous:1986:RRP

[Ano86-39] Anonymous. RAM-resident program security threat. Computers & Security, 5(4):282–285, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886900441.

Anonymous:1986:RE

[Ano86-40] Anonymous. Remember the Enigma! Computers & Security, 5(4):288–289, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886900489.

Anonymous:1986:RAP

[Ano86-41] Anonymous. Risk analysis package. Computers & Security, 5(3):189–190, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886900076.

Anonymous:1986:SR

[Ano86-42] Anonymous. Security reminders. Computers & Security, 5(2): 97–98, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488690129X.

Anonymous:1986:SSR

[Ano86-43] Anonymous. Some suggested reading. Computers & Security, 5 (2):97, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901288.

Anonymous:1986:SIV

[Ano86-44] Anonymous. Subject index to volume 5. Computers & Security, 5(4):365–366, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886900635.

Anonymous:1986:TSR

[Ano86-45] Anonymous. TC 11 status report. Computers & Security, 5(1): 72–73, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901203.

Anonymous:1986:TWa

[Ano86-46] Anonymous. Technology watch. Computers & Security, 5(1): 4–6, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901069.

Anonymous:1986:TWb

[Ano86-47] Anonymous. Technology watch. Computers & Security, 5(2): 94–97, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901276.

Anonymous:1986:WFF

[Ano86-48] Anonymous. Wanted: Floods, fires, disasters, etc. Computers & Security, 5(1):10, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901112.

Anonymous:1986:WSR

[Ano86-49] Anonymous. WG 11:4 status report. Computers & Security, 5(2): 165–166, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901409.

Anonymous:1987:ARAa

[Ano87a] Anonymous. Abstracts of recent articles and literature. Computers & Security, 6(1):80–90, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887901295.

Anonymous:1987:ARAb

[Ano87b] Anonymous. Abstracts of recent articles and literature. Computers & Security, 6(2):180–190, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900927.

Anonymous:1987:ARAd

[Ano87c] Anonymous. Abstracts of recent articles and literature. Computers & Security, 6(4):352–362, August 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900721.

Anonymous:1987:ARAf

[Ano87d] Anonymous. Abstracts of recent articles and literature. Computers & Security, 6(6):517–526, December 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900332.

Anonymous:1987:ARAc

[Ano87e] Anonymous. Abstracts of recent articles and literature: Computer law. Computers & Security, 6(3):269–281, June 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887901088.

Anonymous:1987:ARAe

[Ano87f] Anonymous. Abstracts of recent articles and literature: Computer law. Computers & Security, 6(5):432–441, October 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900162.

Anonymous:1987:AIV

[Ano87g] Anonymous. Author index to volume 6. Computers & Security, 6(6):530–531, December 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900368.

Anonymous:1987:Ca

[Ano87h] Anonymous. Calendar. Computers & Security, 6(1):95–96, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887901325.

Anonymous:1987:Cb

[Ano87i] Anonymous. Calendar. Computers & Security, 6(2):191–192, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900939.

Anonymous:1987:Cc

[Ano87j] Anonymous. Calendar. Computers & Security, 6(3):282–283, June 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488790109X.

Anonymous:1987:Cd

[Ano87k] Anonymous. Calendar. Computers & Security, 6(4):363–364, August 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900733.

Anonymous:1987:Ce

[Ano87l] Anonymous. Calendar. Computers & Security, 6(5):442–443, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900174.

Anonymous:1987:Cf

[Ano87m] Anonymous. Calendar. Computers & Security, 6(6):527, December 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900344.

Anonymous:1987:P

[Ano87n] Anonymous. DOS postscript. Computers & Security, 6(1):15– 16, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887901209.

Anonymous:1987:EB

[Ano87o] Anonymous. Editorial Board. Computers & Security, 6(1):ifc, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887901106.

Anonymous:1987:EEF

[Ano87p] Anonymous. Electromagnetic eavesdropping — fact of fantasy? Computers & Security, 6(1):12–14, February 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887901167.

Anonymous:1987:GC

[Ano87q] Anonymous. Guest comments. Computers & Security, 6(5):382– 384, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887900101.

Anonymous:1987:HSE

[Ano87r] Anonymous. High-speed encrypted storage/backup. Computers & Security, 6(5):370–373, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900034.

Anonymous:1987:IS

[Ano87s] Anonymous. IFIP/Sec. Computers & Security, 6(1):12, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887901155.

Anonymous:1987:LI

[Ano87t] Anonymous. Lock it up. Computers & Security, 6(5):380–381, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900083.

Anonymous:1987:MAU

[Ano87u] Anonymous. Message authentication using the RSA. Computers & Security, 6(5):373–376, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900046.

Anonymous:1987:NMI

[Ano87v] Anonymous. New members of the international board of Editors. Computers & Security, 6(1):4–7, February 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488790112X.

Anonymous:1987:NCC

[Ano87w] Anonymous. NTIS computers, control and information theory. Computers & Security, 6(5):426–431, October 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900150.

Anonymous:1987:OMN

[Ano87x] Anonymous. Other meetings of note. Computers & Security, 6(1): 14, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887901180.

Anonymous:1987:PF

[Ano87y] Anonymous. Pages 1–96 (February 1987). Computers & Security, 6 (1):??, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1987:PJ

[Ano87z] Anonymous. Pages 193–283 (June 1987). Computers & Security, 6(3):??, June 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1987:PAb

[Ano87-27] Anonymous. Pages 285–364 (August 1987). Computers & Security, 6(4):??, August 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1987:PO

[Ano87-28] Anonymous. Pages 365–443 (October 1987). Computers & Security, 6(5):??, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1987:PD

[Ano87-29] Anonymous. Pages 445–531 (December 1987). Computers & Security, 6(6):??, December 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1987:PAa

[Ano87-30] Anonymous. Pages 97–192 (April 1987). Computers & Security, 6(2):??, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1987:RN

[Ano87-31] Anonymous. Random notes. Computers & Security, 6(5):381– 382, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887900095.

Anonymous:1987:RAT

[Ano87-32] Anonymous. Report about TC 11 working groups. Computers & Security, 6(1):14, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887901179.

Anonymous:1987:RA

[Ano87-33] Anonymous. Risk assessment. Computers & Security, 6(4):297– 298, August 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887900630.

Anonymous:1987:SEC

[Ano87-34] Anonymous. Second European conference on computer audit, control and security. Computers & Security, 6(1):94, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887901313.

Anonymous:1987:SRP

[Ano87-35] Anonymous. Secure remote power-on modem. Computers & Security, 6(5):376–378, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900058.

Anonymous:1987:SR

[Ano87-36] Anonymous. Security reminder. Computers & Security, 6(1): 15, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887901192.

Anonymous:1987:SSC

[Ano87-37] Anonymous. Speedier secure communications. Computers & Security, 6(5):379–380, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900071.

Anonymous:1987:SIV

[Ano87-38] Anonymous. Subject index to volume 6. Computers & Security, 6(6):528–529, December 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900356.

Anonymous:1987:SS

[Ano87-39] Anonymous. System security 1986. Computers & Security, 6(1): 91–93, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887901301.

Anonymous:1987:TWA

[Ano87-40] Anonymous. Technology watch — access control to mainframe. Computers & Security, 6(4):295–296, August 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900629.

Anonymous:1987:TWP

[Ano87-41] Anonymous. Technology watch — personal authentication devices. Computers & Security, 6(1):10–11, February 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887901143.

Anonymous:1987:VLR

[Ano87-42] Anonymous. Very light reading. Computers & Security, 6(4):298– 299, August 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887900642.

Anonymous:1987:WIS

[Ano87-43] Anonymous. Write it down safely. Computers & Security, 6(5): 379, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488790006X.

Anonymous:1988:GUD

[Ano88a] Anonymous. 1-2-3 to go unprotected: Douglas Barney. Computers & Security, 7(1):102, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905172.

Anonymous:1988:APa

[Ano88b] Anonymous. Addresses of publications. Computers & Security, 7(1):108–110, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905597.

Anonymous:1988:APb

[Ano88c] Anonymous. Addresses of publications. Computers & Security, 7 (4):433, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906505.

Anonymous:1988:ASV

[Ano88d] Anonymous. Aftershocks for Silicon Valley: Richard Raysman and Peter Brown. Computers & Security, 7(2):213, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903409.

Anonymous:1988:ADD

[Ano88e] Anonymous. Anatomy of a disaster: David Davies. Computers & Security, 7(1):101, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905159.

Anonymous:1988:AFA

[Ano88f] Anonymous. ANSI, Feds address DMBS security: Mitch Betts. Computers & Security, 7(2):220, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903860.

Anonymous:1988:ASS

[Ano88g] Anonymous. Apollo skirts site license issue: James Connolly. Computers & Security, 7(2):217, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903690.

Anonymous:1988:AUS

[Ano88h] Anonymous. Apple urges separate copyright registration of screen displays: Scott mace. Computers & Security, 7(1):101, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888905111.

Anonymous:1988:AIV

[Ano88i] Anonymous. Author index to volume 7. Computers & Security, 7(6):617–618, December 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900144.

Anonymous:1988:BM

[Ano88j] Anonymous. Battling mediocrity. Computers & Security, 7(2): 217, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888903677.

Anonymous:1988:BRCa

[Ano88k] Anonymous. Book review: Computer crime — What to do before it happens: Bernard P. Zajac, Jr. Computers & Security, 7(1): 102, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888905160.

Anonymous:1988:BRCb

[Ano88l] Anonymous. Book review: Computer viruses — a secret threat: Rudiger Dierstein. Computers & Security, 7(2):215, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903537.

Anonymous:1988:BTT

[Ano88m] Anonymous. Buyers turning toward software escrow plans: Carole Patton. Computers & Security, 7(2):215–216, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903574.

Anonymous:1988:Ca

[Ano88n] Anonymous. Calendar. Computers & Security, 7(1):111, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888905603.

Anonymous:1988:Cb

[Ano88o] Anonymous. Calendar. Computers & Security, 7(2):224, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903951.

Anonymous:1988:Cc

[Ano88p] Anonymous. Calendar. Computers & Security, 7(3):332, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888901095.

Anonymous:1988:Cd

[Ano88q] Anonymous. Calendar. Computers & Security, 7(4):435, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888906517.

Anonymous:1988:Ce

[Ano88r] Anonymous. Calendar. Computers & Security, 7(6):616, December 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888900132.

Anonymous:1988:CCSa

[Ano88s] Anonymous. Coalition to combat software piracy: Rachel Parker. Computers & Security, 7(2):219, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903781.

Anonymous:1988:CCSb

[Ano88t] Anonymous. Computer communications software. January 1975– October 1986. Computers & Security, 7(5):519–520, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488890274X.

Anonymous:1988:CCSc

[Ano88u] Anonymous. Computer communications software. November 1986-September 1987. Computers & Security, 7(5):520, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888902751.

Anonymous:1988:CCJa

[Ano88v] Anonymous. Computer crime. January 1975–October 1986 (citations from the INSPEC: Information Services for the Physics and Engineering Communities database). Computers & Security, 7(5): 513, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902283.

Anonymous:1988:CCN

[Ano88w] Anonymous. Computer crime. November 1986–October 1987 (citations from the INSPEC: Information Services for the Physics and Engineering Communities database). Computers & Security, 7(5): 513, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902295.

Anonymous:1988:CDP

[Ano88x] Anonymous. Computer disasters: Prevention and recovery. 1975– September 1986 (citations from the INSPEC: Information Services for the Physics and Engineering Communities database). Computers & Security, 7(5):510, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488890209X.

Anonymous:1988:CISa

[Ano88y] Anonymous. Computer information security and protection. June 1985–April 1987 (citations from the NTIS database). Computers & Security, 7(5):509, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902064.

Anonymous:1988:CISb

[Ano88z] Anonymous. Computer information security and protection. November 1986–October 1987. Computers & Security, 7(5):511, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902179.

Anonymous:1988:CNS

[Ano88-27] Anonymous. Computer network standards. 1975–February 1987 (citations from the INSPEC: Information Services for the Physics and Engineering Communities database). Computers & Security, 7(5):514–515, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902398.

Anonymous:1988:CPE

[Ano88-28] Anonymous. Computer performance evaluation. July 1983–April 1988 (citations from the NTIS database). Computers & Security, 7 (5):511, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902180.

Anonymous:1988:CSS

[Ano88-29] Anonymous. Computer security shift is approved by Senate: Linda Greenhouse. Computers & Security, 7(2):213–214, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903458.

Anonymous:1988:CSM

[Ano88-30] Anonymous. Computer software maintenance. January 1970– August 1987. Computers & Security, 7(5):518, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902635.

Anonymous:1988:CST

[Ano88-31] Anonymous. Computer software transferability and portability. January 1970–August 1987. Computers & Security, 7(5):518, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888902647.

Anonymous:1988:CUS

[Ano88-32] Anonymous. Computer user surveillance: Kevin J. Fitzgerald. Computers & Security, 7(2):215, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903525.

Anonymous:1988:CVR

[Ano88-33] Anonymous. Computer voice recognition: Market aspects. January 1983–July 1987. Computers & Security, 7(5):518, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888902623.

Anonymous:1988:CFR

[Ano88-34] Anonymous. Conference focuses on reduction of software risks: Darryl K. Taft. Computers & Security, 7(2):216, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903604.

Anonymous:1988:CVA

[Ano88-35] Anonymous. The corporate view: Alex Kask. Computers & Security, 7(2):220, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903835.

Anonymous:1988:CCL

[Ano88-36] Anonymous. Crash course: Lessons in preventing computer downtime: Lee Green. Computers & Security, 7(2):217, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903653.

Anonymous:1988:CCJc

[Ano88-37] Anonymous. Cryptography and cryptosystems. January 1970– October 1987. Computers & Security, 7(5):519, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902738.

Anonymous:1988:CCJb

[Ano88-38] Anonymous. Cryptography and cryptosystems. January 1987– December 1987 (citations from the INSPEC: Information Services for the Physics and Engineering Communities database). Computers & Security, 7(5):518, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902593.

Anonymous:1988:DCTb

[Ano88-39] Anonymous. Data compression: Techniques and applications. October 1983–September 1986. Computers & Security, 7(5):518, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888902611.

Anonymous:1988:DCTc

[Ano88-40] Anonymous. Data compression: Techniques and applications. October 1983–September 1986. Computers & Security, 7(5):519, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888902684.

Anonymous:1988:DCTd

[Ano88-41] Anonymous. Data compression: Techniques and applications. October 1986–September 1987. Computers & Security, 7(5):519, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888902696.

Anonymous:1988:DEK

[Ano88-42] Anonymous. Data encryption is key to safe file transmission: Lawrence E. Hughes. Computers & Security, 7(2):221, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903926.

Anonymous:1988:DESb

[Ano88-43] Anonymous. Data encryption standard. 1975–January 1987 (citations from the INSPEC: Information Services for the Physics and Engineering Communities database). Computers & Security, 7(5): 511, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902155.

Anonymous:1988:DESa

[Ano88-44] Anonymous. Data Encryption Standard. January 1975–January 1988 (citations from the INSPEC: Information Services for the Physics and Engineering Communications Database). Computers & Security, 7(5):511, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902143.

Anonymous:1988:DSP

[Ano88-45] Anonymous. Data security provides the key to lock out intruders: Howard C. Berkowitz. Computers & Security, 7(2):216, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903628.

Anonymous:1988:DCL

[Ano88-46] Anonymous. Development of confidence limits by pivotal functions for estimating software reliability. Computers & Security, 7(1): 107, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888905500.

Anonymous:1988:DCTa

[Ano88-47] Anonymous. Dickerson center of tug-of-war: Stephen Jones. Computers & Security, 7(2):213, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903422.

Anonymous:1988:DCD

[Ano88-48] Anonymous. Directory of computerized data files, 1987. Computers & Security, 7(5):519, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902672.

Anonymous:1988:EB

[Ano88-49] Anonymous. Editorial Board. Computers & Security, 7(1):2, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888904865.

Anonymous:1988:ECN

[Ano88-50] Anonymous. Ethernet communication networks. 1978-1986 (citations from the INSPEC: Information Services for the Physics and Engineering Communities database). Computers & Security, 7(5): 515, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902428.

Anonymous:1988:ESV

[Ano88-51] Anonymous. Evaluating signature verification systems is tricky: Rodney Beatson. Computers & Security, 7(2):221–222, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488890394X.

Anonymous:1988:ESNa

[Ano88-52] Anonymous. Expert sees national security risks in BASIS net: Wayne Madsen. Computers & Security, 7(2):221, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903902.

Anonymous:1988:ESNb

[Ano88-53] Anonymous. Expert systems: Non-industrial applications. January 1984–March 1987. Computers & Security, 7(5):519, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888902702.

Anonymous:1988:FTC

[Ano88-54] Anonymous. Fault tolerant computers and multiprocessing. January 1985–June 1987. Computers & Security, 7(1):107, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888905494.

Anonymous:1988:FTA

[Ano88-55] Anonymous. Faulty tolerance adds security to critical on-line transactions: William C. Grover. Computers & Security, 7(2):216, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903586.

Anonymous:1988:FSP

[Ano88-56] Anonymous. Fighting software piracy. Computers & Security, 7(1): 101, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888905135.

Anonymous:1988:F

[Ano88-57] Anonymous. Footnote. Computers & Security, 7(4):346, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888905718.

Anonymous:1988:E

[Ano88-58] Anonymous. From the editor. Computers & Security, 7(1):2, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888904877.

Anonymous:1988:FBL

[Ano88-59] Anonymous. Fuss over bottom line could foil disaster recovery. Computers & Security, 7(2):220, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903859.

Anonymous:1988:GCH

[Ano88-60] Anonymous. German computer hobbyists rifle NASA’s files: Serge Schmemann. Computers & Security, 7(1):102, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905214.

Anonymous:1988:GYD

[Ano88-61] Anonymous. Guarding your data: Julie Anderson. Computers & Security, 7(2):219, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488890377X.

Anonymous:1988:GSP

[Ano88-62] Anonymous. Guidelines or standards! Part one: Yvo Henniker-Heaton. Computers & Security, 7(1):101, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905147.

Anonymous:1988:HRC

[Ano88-63] Anonymous. Handwriting recognition by computer. January 1970– August 1987. Computers & Security, 7(5):518, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902659.

Anonymous:1988:HPC

[Ano88-64] Anonymous. Hardware protection of communications ports and lines: Eugene F. Troy. Computers & Security, 7(2):219, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903811.

Anonymous:1988:HSS

[Ano88-65] Anonymous. Hardware, software security devices protect PCs and inside information: Robert L. Scheier. Computers & Security, 7 (2):218, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888903720.

Anonymous:1988:HIT

[Ano88-66] Anonymous. Highlights of IFIP TC 11 meeting. Computers & Security, 7(3):307–308, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900387.

Anonymous:1988:HRU

[Ano88-67] Anonymous. HK raid uncovers Autodesk pirates. Computers & Security, 7(1):102, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905202.

Anonymous:1988:HLR

[Ano88-68] Anonymous. Hold that lawsuit: Rosemary Hamilton. Computers & Security, 7(2):213, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903446.

Anonymous:1988:HKF

[Ano88-69] Anonymous. Hong Kong focus: Michael D. Pendleton. Computers & Security, 7(2):214, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903483.

Anonymous:1988:H

[Ano88-70] Anonymous. Horatius. Computers & Security, 7(2):215, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903562.

Anonymous:1988:HMD

[Ano88-71] Anonymous. Hospital’s medical data held in strict security. Computers & Security, 7(1):106, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905433.

Anonymous:1988:HSL

[Ano88-72] Anonymous. How to secure a local area network: Philip C. McKinney. Computers & Security, 7(2):216, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903598.

Anonymous:1988:ICM

[Ano88-73] Anonymous. IC card with a microprocessor. Computers & Security, 7(5):519, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902726.

Anonymous:1988:ITCa

[Ano88-74] Anonymous. IFIP technical committee 11. Computers & Security, 7(1):95–96, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905093.

Anonymous:1988:ITCb

[Ano88-75] Anonymous. IFIP technical committee 11. Computers & Security, 7(2):205–206, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903379.

Anonymous:1988:ITCc

[Ano88-76] Anonymous. IFIP technical committee 11. Computers & Security, 7(3):305–306, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900375.

Anonymous:1988:ITCd

[Ano88-77] Anonymous. IFIP technical committee 11. Computers & Security, 7(4):415–416, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905822.

Anonymous:1988:ITCe

[Ano88-78] Anonymous. IFIP technical committee 11. Computers & Security, 7(5):507–508, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902052.

Anonymous:1988:ITCf

[Ano88-79] Anonymous. IFIP technical committee 11. Computers & Security, 7(6):607–608, December 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900119.

Anonymous:1988:ICT

[Ano88-80] Anonymous. Insurance and computers: the technology gap: David Davis. Computers & Security, 7(2):215, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903549.

Anonymous:1988:IPRa

[Ano88-81] Anonymous. Intellectual property rights in an age of electronics and information. Contractor documents. Volume 1. Computers & Security, 7(5):512, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488890226X.

Anonymous:1988:IPRb

[Ano88-82] Anonymous. Intellectual property rights in an age of electronics and information. Contractor documents. Volume 2. Computers & Security, 7(5):513, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902313.

Anonymous:1988:IBRa

[Ano88-83] Anonymous. International board of referees. Computers & Security, 7(1):45–46, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905019.

Anonymous:1988:IBRb

[Ano88-84] Anonymous. International board of referees. Computers & Security, 7(2):165–166, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903331.

Anonymous:1988:IBRc

[Ano88-85] Anonymous. International board of referees. Computers & Security, 7(3):269–270, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900326.

Anonymous:1988:IBRd

[Ano88-86] Anonymous. International board of referees. Computers & Security, 7(4):371–372, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905779.

Anonymous:1988:IBRe

[Ano88-87] Anonymous. International board of referees. Computers & Security, 7(5):473–474, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488890199X.

Anonymous:1988:IBRf

[Ano88-88] Anonymous. International board of referees. Computers & Security, 7(6):557–558, December 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900065.

Anonymous:1988:SLD

[Ano88-89] Anonymous. Is site licensing a dead issue?: Stephanie Wilkinson. Computers & Security, 7(2):218, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903744.

Anonymous:1988:JOP

[Ano88-90] Anonymous. Justice ordered to pay software licensing fees: Tim McGraw. Computers & Security, 7(1):101, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905123.

Anonymous:1988:LPC

[Ano88-91] Anonymous. Legal protection for computer software. Computers & Security, 7(2):213, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903410.

Anonymous:1988:LEa

[Ano88-92] Anonymous. Letters to the editor. Computers & Security, 7(2): 115–116, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888903215.

Anonymous:1988:LEb

[Ano88-93] Anonymous. Letters to the editor. Computers & Security, 7(3): 229–231, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488890017X.

Anonymous:1988:LAN

[Ano88-94] Anonymous. Local area networks. January 1986–December 1986 (citations from the engineering index database). Computers & Security, 7(5):515, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488890243X.

Anonymous:1988:LGH

[Ano88-95] Anonymous. Loophole gives hackers access to DEC machines: Marlene Brown. Computers & Security, 7(2):221, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903938.

Anonymous:1988:LPN

[Ano88-96] Anonymous. Lotus promises new releases, no protection: Virginia dudek. Computers & Security, 7(1):102, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905184.

Anonymous:1988:MKF

[Ano88-97] Anonymous. Manage/38: Kevin Fitzgerald. Computers & Security, 7(2):215, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888903550.

Anonymous:1988:MMF

[Ano88-98] Anonymous. Microcomputer modems. 1975–February 1987 (citations from the INSPEC: Information Services for the Physics and Engineering Communities database). Computers & Security, 7(5): 514, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902386.

Anonymous:1988:MML

[Ano88-99] Anonymous. Microcomputer to mainframe links. January 1975– February 1988 (citations from the INSPEC: Information Services for the Physics and Engineering Communities database). Computers & Security, 7(5):512, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902210.

Anonymous:1988:MDS

[Ano88-100] Anonymous. Microcomputers: Data and software security. 1974– February 1987 (citations from the INSPEC: Information Services for the Physics and Engineering Communities database). Computers & Security, 7(5):512, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902246.

Anonymous:1988:MBS

[Ano88-101] Anonymous. Microsoft, Borland settle Dickerson case: Rachel Parker. Computers & Security, 7(2):213, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903434.

Anonymous:1988:MAS

[Ano88-102] Anonymous. MIS assumes a shirt-sleeves approach to protecting power: Jeff Goodwin. Computers & Security, 7(2):218, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903756.

Anonymous:1988:MCS

[Ano88-103] Anonymous. MIS computer security slacking: June Altman. Computers & Security, 7(1):102, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905196.

Anonymous:1988:MSH

[Ano88-104] Anonymous. Most systems held fast while storm clouds grew on wall st: Alan Alper. Computers & Security, 7(2):218, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903732.

Anonymous:1988:NSC

[Ano88-105] Anonymous. NASA says classified data not breached: Rick Vizachero. Computers & Security, 7(1):102–103, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888905226.

Anonymous:1988:NPD

[Ano88-106] Anonymous. NASD power dip could recur: Barton Crockett. Computers & Security, 7(2):220, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903872.

Anonymous:1988:NNS

[Ano88-107] Anonymous. NBS and NSA stand on different sides of security: Evelyn C. MacDonald. Computers & Security, 7(2):221, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903914.

Anonymous:1988:NFG

[Ano88-108] Anonymous. Net failures growing concern in workplace: Mark Stevens. Computers & Security, 7(2):217, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903641.

Anonymous:1988:NCA

[Ano88-109] Anonymous. Neutral computers, auditing and security. Computers & Security, 7(5):520, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902763.

Anonymous:1988:NMB

[Ano88-110] Anonymous. “Never mind!” businesses cool off on site license issue: Ed Scannell. Computers & Security, 7(1):103, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905238.

Anonymous:1988:NLK

[Ano88-111] Anonymous. New locks and keys for electronic information. Computers & Security, 7(1):89–93, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905081.

Anonymous:1988:NWL

[Ano88-112] Anonymous. New wave of LAN security products hailed by users: Susan Kerr. Computers & Security, 7(2):217, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903665.

Anonymous:1988:NEW

[Ano88-113] Anonymous. No easy way through data-security maze, consultants say: Robert L. Scheier. Computers & Security, 7(2):218, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903719.

Anonymous:1988:NNC

[Ano88-114] Anonymous. NSA, NBS coming together on computer security. Computers & Security, 7(2):217, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903689.

Anonymous:1988:PF

[Ano88-115] Anonymous. Pages 1–111 (February 1988). Computers & Security, 7(1):??, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1988:PAa

[Ano88-116] Anonymous. Pages 114–224 (April 1988). Computers & Security, 7(2):??, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1988:PJ

[Ano88-117] Anonymous. Pages 228–332 (June 1988). Computers & Security, 7(3):??, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1988:PAb

[Ano88-118] Anonymous. Pages 334–435 (August 1988). Computers & Security, 7(4):??, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1988:PO

[Ano88-119] Anonymous. Pages 438–530 (October 1988). Computers & Security, 7(5):??, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1988:PD

[Ano88-120] Anonymous. Pages 532–619 (December 1988). Computers & Security, 7(6):??, December 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1988:PMT

[Ano88-121] Anonymous. PC–mainframe ties hard to guard: Paula S. Stone. Computers & Security, 7(2):216–217, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890363X.

Anonymous:1988:PSP

[Ano88-122] Anonymous. Policies for secure personal data: Thomas B. Riley. Computers & Security, 7(2):219, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890380X.

Anonymous:1988:PED

[Ano88-123] Anonymous. Processing encrypted data: Niv Ahituv, Yeheskel Lapid, and Seev Neumann. Computers & Security, 7(1):103, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488890524X.

Anonymous:1988:PPI

[Ano88-124] Anonymous. Public policy and internal privacy codes: Tom Riley. Computers & Security, 7(1):103, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905263.

Anonymous:1988:RBB

[Ano88-125] Anonymous. Random bits and bytes. Computers & Security, 7(2): 117–127, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888903227.

Anonymous:1988:RIA

[Ano88-126] Anonymous. Remote identification and authentication of computer resource users: Ken Weiss. Computers & Security, 7(2): 214, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888903495.

Anonymous:1988:RSCa

[Ano88-127] Anonymous. Report says computers spy on 7 million workers in US. Computers & Security, 7(1):103–104, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905299.

Anonymous:1988:RSCb

[Ano88-128] Anonymous. Review: Spread of computer viruses worries users: John Voelcher The Institute, IEEE Spectrum, June 1988, p. 1 ff. Computers & Security, 7(4):427, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906177.

Anonymous:1988:RLV

[Ano88-129] Anonymous. Revised list of virus filters. Computers & Security, 7(3):259–260, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900302.

Anonymous:1988:SAC

[Ano88-130] Anonymous. Sabotage aimed at computer company destroys government data. Computers & Security, 7(5):522–523, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888902878.

Anonymous:1988:SSV

[Ano88-131] Anonymous. Secure software verification tools. Computers & Security, 7(5):511, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902192.

Anonymous:1988:SBC

[Ano88-132] Anonymous. Security basics for cable/video display radiation. Computers & Security, 7(3):330–331, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888901083.

Anonymous:1988:SCU

[Ano88-133] Anonymous. A security checklist for upper management: Jerome Lobel. Computers & Security, 7(2):219, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903793.

Anonymous:1988:SCM

[Ano88-134] Anonymous. Security considerations for minicomputer systems acquisitions: Sandra M. Mann Lambert. Computers & Security, 7 (2):219, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888903823.

Anonymous:1988:SVE

[Ano88-135] Anonymous. Security in the VM environment: William H. Murray. Computers & Security, 7(2):214, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903501.

Anonymous:1988:SIC

[Ano88-136] Anonymous. Security issues concern NSA, vendors. Computers & Security, 7(1):106, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905457.

Anonymous:1988:SPE

[Ano88-137] Anonymous. Security planning for EFT network systems: A. D. Davis. Computers & Security, 7(2):214–215, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903513.

Anonymous:1988:SSG

[Ano88-138] Anonymous. Security systems: Getting management to shell out: Larry Stevens. Computers & Security, 7(2):220–221, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903896.

Anonymous:1988:SAF

[Ano88-139] Anonymous. Senate approves funds for DEA Tempest equipment: Vanessa Jo Grim. Computers & Security, 7(2):216, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903616.

Anonymous:1988:SWL

[Ano88-140] Anonymous. Shrink-wrap licensing: Christopher J. Millard. Computers & Security, 7(2):214, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488890346X.

Anonymous:1988:SCJ

[Ano88-141] Anonymous. Smart cards. January 1983–August 1987. Computers & Security, 7(5):518–519, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902660.

Anonymous:1988:SQMa

[Ano88-142] Anonymous. Software quality and metrics. January 1975– December 1986 (citations from the INSPEC: Information Services for the Physics and Engineering Communities database). Computers & Security, 7(5):517, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488890257X.

Anonymous:1988:SQMb

[Ano88-143] Anonymous. Software quality and metrics. January 1987– December 1987 (citations from the INSPEC: Information Services for the Physics and Engineering Communities database). Computers & Security, 7(5):517–518, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902581.

Anonymous:1988:SSP

[Ano88-144] Anonymous. Software security and piracy. February 1984–March 1986 (citations from the INSPEC: Information Services for the Physics and Engineering Communities database). Computers & Security, 7(5):513, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902325.

Anonymous:1988:SSF

[Ano88-145] Anonymous. The structure of system/88, a fault-tolerant computer: E. S. Harrison and E. J. Schmitt. Computers & Security, 7(2):218, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888903707.

Anonymous:1988:SIV

[Ano88-146] Anonymous. Subject index to volume 7. Computers & Security, 7(6):619, December 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900156.

Anonymous:1988:TPD

[Ano88-147] Anonymous. Tactical planning for data security management: Sandra M. Mann Lambert. Computers & Security, 7(1):103, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888905251.

Anonymous:1988:CTL

[Ano88-148] Anonymous. To catch a thief: Lessons in systems security: Michael Ball. Computers & Security, 7(2):220, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903884.

Anonymous:1988:UPS

[Ano88-149] Anonymous. Uninterruptible power supplies. Computers & Security, 7(2):218–219, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903768.

Anonymous:1988:VDA

[Ano88-150] Anonymous. Virus defense alert. Computers & Security, 7(2): 156, 158, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888903306.

Anonymous:1988:WFT

[Ano88-151] Anonymous. Wanted: Full-time contingency planners: David A. Ludlum. Computers & Security, 7(2):220, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903847.

Anonymous:1988:WSa

[Ano88-152] Anonymous. What others are saying. Computers & Security, 7(2): 210–212, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888903392.

Anonymous:1988:WSb

[Ano88-153] Anonymous. What others are saying. Computers & Security, 7(3): 316–319, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888900417.

Anonymous:1988:WWRa

[Ano88-154] Anonymous. World-wide reports on developments in computer security. Computers & Security, 7(1):97–100, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890510X.

Anonymous:1988:WWRc

[Ano88-155] Anonymous. World-wide reports on developments in computer security. Computers & Security, 7(4):417–422, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905834.

Anonymous:1988:WWRb

[Ano88-156] Anonymous. World-wide reports on developments in computer security: Part II. Computers & Security, 7(2):207–209, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903380.

Anonymous:1989:ARAa

[Ano89a] Anonymous. Abstracts of recent articles and literature. Computers & Security, 8(7):633–640, November 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900576.

Anonymous:1989:ARAb

[Ano89b] Anonymous. Abstracts of recent articles and literature. Computers & Security, 8(8):733–739, December 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900126.

Anonymous:1989:AI

[Ano89c] Anonymous. Author index. Computers & Security, 8(8):743, December 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488990014X.

Anonymous:1989:Ca

[Ano89d] Anonymous. Calendar. Computers & Security, 8(1):83, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488990045X.

Anonymous:1989:Cb

[Ano89e] Anonymous. Calendar. Computers & Security, 8(2):173, April 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889900734.

Anonymous:1989:Cc

[Ano89f] Anonymous. Calendar. Computers & Security, 8(3):270, May 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889901107.

Anonymous:1989:Cd

[Ano89g] Anonymous. Calendar. Computers & Security, 8(4):363, June 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889900989.

Anonymous:1989:Ce

[Ano89h] Anonymous. Calendar. Computers & Security, 8(5):456, August 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488990028X.

Anonymous:1989:Cf

[Ano89i] Anonymous. Calendar. Computers & Security, 8(6):547–548, October 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889900849.

Anonymous:1989:Cg

[Ano89j] Anonymous. Calendar. Computers & Security, 8(7):641–642, November 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889900588.

Anonymous:1989:Ch

[Ano89k] Anonymous. Calendar. Computers & Security, 8(8):741, December 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889900138.

Anonymous:1989:EB

[Ano89l] Anonymous. Editorial Board. Computers & Security, 8(1):2, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889900291.

Anonymous:1989:E

[Ano89m] Anonymous. From the editor. Computers & Security, 8(7):550– 552, November 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889900461.

Anonymous:1989:ITCa

[Ano89n] Anonymous. IFIP technical committee 11. Computers & Security, 8(1):67–68, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900412.

Anonymous:1989:ITCb

[Ano89o] Anonymous. IFIP technical committee 11. Computers & Security, 8(2):161–162, April 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900710.

Anonymous:1989:ITCc

[Ano89p] Anonymous. IFIP technical committee 11. Computers & Security, 8(3):257–258, May 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889901089.

Anonymous:1989:ITCd

[Ano89q] Anonymous. IFIP technical committee 11. Computers & Security, 8(4):299–302, June 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900904.

Anonymous:1989:ITCe

[Ano89r] Anonymous. IFIP technical committee 11. Computers & Security, 8(5):443–444, August 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900266.

Anonymous:1989:ITCf

[Ano89s] Anonymous. IFIP technical committee 11. Computers & Security, 8(6):535–536, October 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900825.

Anonymous:1989:ITCg

[Ano89t] Anonymous. IFIP technical committee 11. Computers & Security, 8(7):631–632, November 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900564.

Anonymous:1989:ITCh

[Ano89u] Anonymous. IFIP technical committee 11. Computers & Security, 8(8):731–732, December 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900114.

Anonymous:1989:IBRa

[Ano89v] Anonymous. International board of referees. Computers & Security, 8(1):33–34, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900369.

Anonymous:1989:IBRb

[Ano89w] Anonymous. International board of referees. Computers & Security, 8(2):121–122, April 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900679.

Anonymous:1989:IBRc

[Ano89x] Anonymous. International board of referees. Computers & Security, 8(3):207, May 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488990103X.

Anonymous:1989:IBRd

[Ano89y] Anonymous. International board of referees. Computers & Security, 8(4):345–346, June 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900953.

Anonymous:1989:IBRe

[Ano89z] Anonymous. International board of referees. Computers & Security, 8(5):397, August 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900217.

Anonymous:1989:IBRf

[Ano89-27] Anonymous. International board of referees. Computers & Security, 8(6):493, October 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900783.

Anonymous:1989:IBRg

[Ano89-28] Anonymous. International board of referees. Computers & Security, 8(7):605, November 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900527.

Anonymous:1989:IBRh

[Ano89-29] Anonymous. International board of referees. Computers & Security, 8(8):691, December 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900072.

Anonymous:1989:LE

[Ano89-30] Anonymous. Letters to the editor. Computers & Security, 8(2): 87–90, April 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889900606.

Anonymous:1989:PM

[Ano89-31] Anonymous. Pages 176–270 (May 1989). Computers & Security, 8(3):??, May 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1989:PF

[Ano89-32] Anonymous. Pages 2–83 (February 1989). Computers & Security, 8 (1):??, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1989:PJ

[Ano89-33] Anonymous. Pages 272–363 (June 1989). Computers & Security, 8(4):??, June 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1989:PAb

[Ano89-34] Anonymous. Pages 366–456 (August 1989). Computers & Security, 8(5):??, August 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1989:PO

[Ano89-35] Anonymous. Pages 458–548 (October 1989). Computers & Security, 8(6):??, October 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1989:PN

[Ano89-36] Anonymous. Pages 550–642 (November 1989). Computers & Security, 8(7):??, November 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1989:PD

[Ano89-37] Anonymous. Pages 644–749 (December 1989). Computers & Security, 8(8):??, December 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1989:PAa

[Ano89-38] Anonymous. Pages 86–173 (April 1989). Computers & Security, 8(2):??, April 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

Anonymous:1989:SAS

[Ano89-39] Anonymous. Special abstract: Study of computer “viruses” and their potential for infecting Commonwealth computer systems. Computers & Security, 8(1):69–71, February 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900424.

Anonymous:1989:SI

[Ano89-40] Anonymous. Subject index. Computers & Security, 8(8):745– 749, December 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889900151.

Anonymous:1989:WSa

[Ano89-41] Anonymous. What others are saying. Computers & Security, 8(1): 72–73, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889900436.

Anonymous:1989:WSb

[Ano89-42] Anonymous. What others are saying. Computers & Security, 8(4): 347–352, June 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889900965.

Auerbach:1985:PRI

[Aue85] Isaac L. Auerbach. Professional responsibility for information privacy. Computers & Security, 4(2):103–107, June 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900148.

Ayoub:1983:DCE

[Ayo83] F. Ayoub. The design of complete encryption networks using cryptographically equivalent permutations. Computers & Security, 2(3):261–267, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488390010X.

Backwith:1983:UAS

[Bac83] Nigel Backwith. Unique approach to security evaluation. Computers & Security, 2(1):35–40, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900329.

ben-Aaron:1988:TCC

[bAG88] Diana ben Aaron and Lee Green. Today’s computer crime: the threat from within. Computers & Security, 7(1):107, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888905548.

Banerjee:1982:HSI

[Ban82] Subodh K. Banerjee. High speed implementation of DES. Computers & Security, 1(3):261–267, November 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900451.

Barker:1988:UCS

[Bar88] Paul Barker. Unisys Canada slaps $500 k lawsuit on firm. Computers & Security, 7(5):522, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902830.

Barton:1984:UFP

[BB84] Ben F. Barton and Marthalee S. Barton. User-friendly password methods for computer-mediated information systems. Computers & Security, 3(3):186–195, August 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900403.

Biskup:1988:PMD

[BB88] Joachim Biskup and Hans Hermann Brüggeman. The personal model of data:: Towards a privacy-oriented information system. Computers & Security, 7(6):575–597, December 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900090.

Blom:1982:SMD

[BBFÅ82] Rolf Blom, Jan-Olof Brüer, Viiveke Fåk, and S. Ingvar Åkersten. On security measures in distributed computer systems. Computers & Security, 1(2):113–122, June 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404882900049.

Baird:1987:MC

[BBR87] Bruce J. Baird, Lindsay L. Baird, and Ronald P. Ranauro. The moral cracker? Computers & Security, 6(6):471–478, December 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900289.

Ball:1982:DRS

[BDE^{+}82] Leslie D. Ball, Gail Dentch, Mary Emerson, Martha Lewis, Susan McWhorter, and Frank Turgeon. Disaster recovery services. Computers & Security, 1(3):216–225, November 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900360.

Badenhorst:1989:FML

[BE89] K. P. Badenhorst and Jan H. P. Eloff. Framework of a methodology for the life cycle of computer security in an organization. Computers & Security, 8(5):433–442, August 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900254.

Beatson:1986:DOA

[Bea86a] John G. Beatson. Development and organization of the audit and security functions. Computers & Security, 5(2):168, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404886901495.

Beatson:1986:MEA

[Bea86b] John G. Beatson. Managing the EDP audit and security functions. Computers & Security, 5(3):201–206, September 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886900118.

Beckman:1983:IFS

[Bec83] Kristian Beckman. IFIP’s first security congress: the multi-million step. Computers & Security, 2(2):117–119, June 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900494.

Bennett:1988:AEA

[Ben88] John Bennett. Analysis of the encryption algorithm used in the WordPerfect word processing program. Computers & Security, 7(1):105, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905366.

Betts:1988:ESG

[Bet88] Mitch Betts. Encryption standard to get reprieve. Computers & Security, 7(1):106, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905470.

Bishop:1986:PCS

[Bis86] Matt Bishop. A pauper’s callback scheme. Computers & Security, 5(2):141–144, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901379.

BloomBecker:1982:ICC

[Blo82] Jay BloomBecker. International computer crime: Where terrorism and transborder data flow meet. Computers & Security, 1(1):41– 53, January 1982. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404882900244.

Bussolati:1983:TNA

[BM83] U. Bussolati and G. Martella. Towards a new approach to secure database design. Computers & Security, 2(1):49–62, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404883900342.

Bologna:1988:PMM

[Bol88a] G. J. Bologna. Pitfalls in measuring MIS performance. Computers & Security, 7(2):137–138, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903264.

Bologna:1988:MIS

[Bol88b] Jack Bologna. Managing an information system crisis. Computers & Security, 7(3):245–247, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900260.

Bologna:1988:SRH

[Bol88c] Jack Bologna. Selection risks in hiring information systems personnel. Computers & Security, 7(4):353–355, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905731.

Bologna:1988:SWC

[Bol88d] Jack Bologna. Soviet white-collar crime and criminal justice. Computers & Security, 7(6):553–556, December 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900053.

Bologna:1989:TTR

[Bol89a] G. J. Bologna. Twenty-two reasons why MIS subordinates don’t measure up to their bosses’ expectations. Computers & Security, 8(5):395–396, August 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900205.

Bologna:1989:OMF

[Bol89b] Jack Bologna. The one minute fraud auditor. Computers & Security, 8(1):29–31, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900357.

Bonyun:1983:UAP

[Bon83] David A. Bonyun. The use of architectural principles in the design of certifiably secure systems. Computers & Security, 2(2): 153–162, June 1983. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404883900536.

Bottom:1983:ITS

[Bot83] Norman R. Bottom. An informational theory of security. Computers & Security, 2(3):275–280, November 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900123.

Bound:1988:DST

[Bou88] William A. J. Bound. Discussing security with top management. Computers & Security, 7(2):129–130, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903239.

Bozman:1988:ISR

[Boz88a] Jean S. Bozman. IBM slow to remedy leaky DB2 security. Computers & Security, 7(1):104, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905330.

Bozman:1988:USF

[Boz88b] Jean S. Bozman. Users see flaws in audit trails, access authority. Computers & Security, 7(1):104, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905342.

Burk:1989:DPS

[BP89] Holger Bürk and Andreas Pfitzmann. Digital payment systems enabling security and unobservability. Computers & Security, 8(5): 399–416, August 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889900229.

Baum:1988:ESS

[BPR88] William E. Baum, Judith H. Podell, and G. N. Romstedt. Expert system for software quality assurance. Computers & Security, 7(5): 512, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902258.

Bound:1983:MRA

[BR83] William A. J. Bound and Dennis R. Ruth. Making risk analysis a useful management tool with microcomputer electronic worksheet packages. Computers & Security, 2(2):102–115, June 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900482.

Bong:1989:OSI

[BR89] Dieter Bong and Christoph Ruland. Optimized software implementations of the modular exponentiation on general purpose microprocessors. Computers & Security, 8(7):621–630, November 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889900552.

Braum:1988:SCD

[Bra88] H. Braum. Security considerations — defence-related and commercial. Computers & Security, 7(5):513, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902271.

Bridges:1988:CVS

[Bri88] Linda Bridges. Computer viruses strike colleges and companies, leaving them defensive. Computers & Security, 7(5):522, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888902866.

Branstad:1982:ISS

[BS82] Dennis K. Branstad and Miles E. Smid. Integrity and security standards based on cryptography. Computers & Security, 1(3):255– 260, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488290044X.

Bui:1987:CEM

[BS87] Tung Bui and Taracad R. Sivasankaran. Cost-effectiveness modeling for a decision support system in computer security. Computers & Security, 6(2):139–151, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488790085X.

Buck:1982:PCS

[Buc82] R. Creighton Buck. The public cryptography study group. Computers & Security, 1(3):249–254, November 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900438.

Bump:1984:PSP

[Bum84] Morrison Bump. A primer on software piracy cases in the courts. Computers & Security, 3(2):123–134, May 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900543.

Burnham:1988:DES

[Bur88] B. Burnham. DES (data encryption standard) cryptographic services designed for the DOE wide band communications network. Computers & Security, 7(5):510, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890212X.

Buurmeijer:1984:IDS

[Buu84] F. Buurmeijer. IBM’s data security strategy: Some implementation aspects. Computers & Security, 3(4):273–277, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404884900051.

Boshoff:1989:PCM

[BvS89] W. H. Boshoff and S. H. von Solms. A path context model for addressing security in potentially non-secure environments. Computers & Security, 8(5):417–425, August 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900230.

Byles:1988:LFD

[Byl88] Torrey Byles. “look-and-feel” decision looms. Computers & Security, 7(5):521, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902775.

Callaghan:1983:SDW

[Cal83] David R. Callaghan. Securing the distributed word processing network. Computers & Security, 2(1):78–81, January 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488390038X.

Carroll:1982:CCB

[Car82] John M. Carroll. The control of computer-based fraud. Computers & Security, 1(2):123–138, June 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404882900050.

Carroll:1983:DSR

[Car83] John M. Carroll. Decision support for risk analysis. Computers & Security, 2(3):230–236, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900056.

Carroll:1987:SEU

[Car87] John M. Carroll. Strategies for extending the useful lifetime of DES. Computers & Security, 6(4):300–313, August 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900654.

Carroll:1988:IMS

[Car88] John M. Carroll. Implementing multilevel security by violation privilege. Computers & Security, 7(6):563–573, December 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888900089.

Cohen:1988:TSF

[CB88] Fred Cohen and Stephen Bruhns. Two secure file servers. Computers & Security, 7(4):409–414, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905810.

Cerullo:1983:DCC

[Cer83] Michael J. Cerullo. Data communications controls. Computers & Security, 2(1):67–72, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900366.

Cerullo:1985:GCC

[Cer85] Michael J. Cerullo. General controls in computer systems. Computers & Security, 4(1):33–45, March 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900070.

Chamoux:1985:DSC

[Cha85] J. P. Chamoux. Data security and confidentiality in Europe. Computers & Security, 4(3):207–210, September 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488590029X.

Chandler:1988:LIN

[Cha88] Harriette L. Chandler. LANS introduce new security issues. Computers & Security, 7(1):106, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905469.

Chien:1988:CCR

[Chi88] A. A. Chien. Congestion control in routing networks. Computers & Security, 7(5):515, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902416.

Christoffersson:1988:MAE

[Chr88] Per Christoffersson. Message authentication and encryption combined. Computers & Security, 7(1):65–71, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905056.

Carroll:1985:CFA

[CI85] John M. Carroll and W. R. Mac Iver. COSSAC: a framework for analyzing and configuring secure computer facilities. Computers & Security, 4(1):5–12, March 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900045.

Carroll:1986:DSR

[CJ86] John M. Carroll and Helmut Jürgensen. Design of a secure relational data base. Computers & Security, 5(2):167, June 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886901422.

Crockett:1988:ASU

[CM88] Barton Crockett and Cassimir J. Medford. After the street’s unforgettable week all info systems are (mostly) go. Computers & Security, 7(1):107, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905512.

Carroll:1988:PPT

[CMRW88] John M. Carroll, Robert B. Mowat, Lynda E. Robbins, and David Wiseman. The password predictor — a training aid for raising security awareness. Computers & Security, 7(5):475–481, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888902003.

Cohen:1985:SCN

[Coh85] Fred Cohen. A secure computer network design. Computers & Security, 4(3):189–205, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900288.

Cohe:1987:PAI

[Coh87a] Fred Cohe. Protection and administration of information networks with partial orderings. Computers & Security, 6(2):118–128, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900824.

Cohen:1987:CVT

[Coh87b] Fred Cohen. Computer viruses: Theory and experiments. Computers & Security, 6(1):22–35, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887901222.

Cohen:1987:CCI

[Coh87c] Fred Cohen. A cryptographic checksum for integrity protection. Computers & Security, 6(6):505–510, December 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900319.

Cohen:1987:DAD

[Coh87d] Fred Cohen. Design and administration of distributed and hierarchical information networks under partial orderings. Computers & Security, 6(3):219–228, June 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887901039.

Cohen:1987:DPI

[Coh87e] Fred Cohen. Design and protection of an information network under a partial ordering: a case study. Computers & Security, 6(4): 332–338, August 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488790068X.

Cohen:1988:DPC

[Coh88a] Fred Cohen. Designing provably correct information networks with digital diodes. Computers & Security, 7(3):279–286, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488890034X.

Cohen:1988:MPP

[Coh88b] Fred Cohen. Maintaining a poor person’s information integrity. Computers & Security, 7(5):489–494, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902027.

Cohen:1988:ICV

[Coh88c] Fred Cohen. On the implications of computer viruses and methods of defense. Computers & Security, 7(2):167–184, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903343.

Cohen:1988:EIC

[Coh88d] Frederick Cohen. Ethical issues in computer virus distribution. Computers & Security, 7(4):335–336, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905627.

Cohen:1989:CAC

[Coh89a] Fred Cohen. Computational aspects of computer viruses. Computers & Security, 8(4):297–298, June 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900898.

Cohen:1989:MPD

[Coh89b] Fred Cohen. Models of practical defenses against computer viruses. Computers & Security, 8(2):149–160, April 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900709.

Comer:1984:CCN

[Com84] Michael Comer. Computer crime needs international countermeasures. Computers & Security, 3(1):53–55, February 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900294.

Courtney:1982:SAD

[Cou82] Robert H. Courtney. A systematic approach to data security. Computers & Security, 1(2):99–112, June 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900037.

Courtney:1988:APS

[Cou88] Robert H. Courtney. Another perspective on sensitive but unclassified data. Computers & Security, 7(1):19–23, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888904968.

Carroll:1988:ACP

[CR88] John M. Carroll and Lynda Robbins. The automated cryptananysis of polyalphabetic ciphers. Computers & Security, 7(1):104, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888905354.

Caroll:1983:MSA

[CW83] John M. Caroll and Oi-Lun Wu. Methodology for security analysis of data-processing systems. Computers & Security, 2(1):24– 34, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404883900317.

Davies:1986:HUS

[Dav86] Donald W. Davies. How to use the DES safely. Computers & Security, 5(2):169, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901562.

David:1988:TVF

[Dav88] Jon David. Treating viral fever. Computers & Security, 7(3): 255–258, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888900296.

deBruyne:1985:SVU

[dB85] P. de Bruyne. Signature verification using holistic measures. Computers & Security, 4(4):309–315, December 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900495.

duCroix:1985:DSA

[dC85] A. J. du Croix. Data sharing and access protection in business system 12. Computers & Security, 4(4):317–323, December 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900501.

Dehnad:1989:SWI

[Deh89] Khosrow Dehnad. A simple way of improving the login security. Computers & Security, 8(7):607–611, November 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900539.

DeMaio:1989:VMI

[DeM89] Harry B. DeMaio. Viruses — a management issue. Computers & Security, 8(5):381–388, August 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900187.

Dwyer:1987:MLS

[DJT87] Patricia A. Dwyer, George D. Jelatis, and Bhavani M. Thuraisingham. Multi-level security in database management systems. Computers & Security, 6(3):245–251, June 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887901052.

Doswell:1986:ACS

[Dos86] Roger T. Doswell. The audit, control and security of the project life cycle — pre-implementation stage. Computers & Security, 5 (2):168, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901513.

David:1988:UCE

[DR88] H. David and J. Roelofsen. Use of colour on electronic data displays (second study). Computers & Security, 7(5):512, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888902222.

EHH:1985:BRT

[EHH85a] EHH. Book review: Transborder data flows: access to the international on-line data-base market: Amsterdam: North-Holland, 1984. 130 + xiv pages, Dfl 75.00/US $28.75. Computers & Security, 4(3):249, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900410.

EHH:1985:PSC

[EHH85b] EHH. Proceedings of the seventh computer security group conference: Washington: US Government Printing Office, 1984 — 776-101/1027. 115 pages. Computers & Security, 4(3):249–250, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404885900422.

Ekebrink:1986:DST

[Eke86a] Ivan Ekebrink. Data security in terminalized system. Computers & Security, 5(4):325–327, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886900532.

Ekerbrink:1986:SET

[Eke86b] Ivan Ekerbrink. Security of electronic transactions. Computers & Security, 5(2):168, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901483.

Eloff:1983:SPS

[Elo83] Jan H. P. Eloff. Selection process for security packages. Computers & Security, 2(3):256–260, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900093.

Eloff:1985:DSL

[Elo85] Jan H. P. Eloff. The development of a specification language for a computer security system. Computers & Security, 4(2):143–147, June 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404885900197.

Eloff:1988:CSP

[Elo88] Jan H. P. Eloff. Computer security policy: Important issues. Computers & Security, 7(6):559–562, December 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900077.

Erdem:1986:HCO

[Erd86] Hilmi Erdem. Host cryptographic operations: a software implementation. Computers & Security, 5(4):344–346, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488690057X.

Essen:1986:NMI

[Ess86] Johan Essen. New member of the international board of Editors. Computers & Security, 5(2):84, June 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886901240.

Faak:1986:SVH

[Fåk86] Viiveke Fåk. Software versus hardware encryption — is there any difference today? Computers & Security, 5(2):167, June 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886901434.

Faak:1987:CMM

[Fåk87a] Viiveke Fåk. Crypto management made manageable — demands on crypto equipment design. Computers & Security, 6(1):36– 40, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887901234.

Faak:1987:RWC

[Fåk87b] Viiveke Fåk. Report from the working conference on “the Crypto User’s Handbook” held in Stockholm, October 22–23, 1986 by IFIP WG 11:4. Computers & Security, 6(4):351, August 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488790071X.

Faak:1988:WVV

[Fåk88] Viiveke Fåk. Are we vulnerable to a virus attack?. A report from Sweden. Computers & Security, 7(2):151–155, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890329X.

Feuerlicht:1989:RCI

[FG89] J. Feuerlicht and P. Grattan. The role of classification of information in controlling data proliferation in end-user personal computer environment. Computers & Security, 8(1):59–66, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889900400.

Fifield:1989:SOC

[Fif89] Ken J. Fifield. Smartcards outsmart computer crime. Computers & Security, 8(3):247–255, May 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889901077.

Fisher:1984:CCS

[Fis84] Warren W. Fisher. Cryptography for computer security: Making the decision. Computers & Security, 3(3):229–233, August 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404884900440.

Fry:1983:CME

[FM83] Brian G. P. Fry and William F. Main. A conceptual methodology for evaluating security requirements for data assets. Computers & Security, 2(3):237–241, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900068.

Fugini:1984:ACM

[FM84] M. Fugini and G. Martella. ACTEN: a conceptual model for security systems design. Computers & Security, 3(3):196–214, August 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404884900415.

Fordyce:1982:CSC

[For82] Samantha Fordyce. Computer security: a current assessment. Computers & Security, 1(1):9–16, January 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900207.

Frankel:1988:GSP

[Fra88] S. Frankel. Guidance on software package selection. Computers & Security, 7(5):515, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902404.

Friedman:1982:CDP

[Fri82] Stanley Friedman. Contingency and disaster planning. Computers & Security, 1(1):34–40, January 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404882900232.

Fernandez:1987:ACA

[FVTS87] C. Fernández, A. Vaquero, J. M. Troya, and J. M. Sánchez. Automating the computation of authenticators for interbank telex messages. Computers & Security, 6(5):396–402, October 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900125.

Gaglione:1988:ITP

[Gag88a] A. M. Gaglione. Information theory and public key cryptosystems. Computers & Security, 7(5):511, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902167.

Gaglione:1988:SCT

[Gag88b] A. M. Gaglione. Some complexity theory for cryptography. Computers & Security, 7(5):519, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902714.

Gardner:1989:EFR

[Gar89a] Phillip E. Gardner. Evaluation of five risk assessment programs. Computers & Security, 8(6):479–485, October 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488990076X.

Gardner:1989:IWW

[Gar89b] Phillip E. Gardner. The Internet worm: What was said and when. Computers & Security, 8(4):305–316, June 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900928.

Guillou:1982:CT

[GL82] L. C. Guillou and B. Lorig. Cryptography and teleinformatics. Computers & Security, 1(1):27–33, January 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900220.

Glath:1988:CGV

[Gla88] Raymond M. Glath. The case of the “Gerbil virus” that wasn’t. Computers & Security, 7(5):451–453, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888901952.

Gleissner:1989:MTS

[Gle89] Winfried Gleissner. A mathematical theory for the spread of computer viruses. Computers & Security, 8(1):35–41, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889900370.

Gal:1985:SIA

[GM85] Graham Gal and William E. McCarthy. Specification of internal accounting controls in a database environment. Computers & Security, 4(1):23–32, March 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900069.

Goeltz:1986:W

[Goe86] Ted Goeltz. Why not DES? Computers & Security, 5(1):24– 27, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901148.

Goff:1988:HPC

[Gof88a] Leslie Goff. H-P countersues Apple. Computers & Security, 7(5): 521, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902817.

Goff:1988:IUT

[Gof88b] Leslie Goff. IBM users told data files could have legal pitfalls. Computers & Security, 7(1):105, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905408.

Goldstein:1983:CCP

[Gol83] Bruce Goldstein. Computer crime and its prevention. Computers & Security, 2(1):63–66, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900354.

Greene:1988:STAa

[GPR^{+}88a] Joseph S. Greene, T. Probert, W. Riddle, C. Giese, and T. L. Quindry. Software technology for adaptable, reliable systems (STARS) program management plan. Computers & Security, 7(5): 515, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902453.

Greene:1988:STAb

[GPR^{+}88b] Joseph S. Greene, T. Probert, W. Riddle, C. Giese, and T. L. Quindry. Software technology for adaptable, reliable systems (STARS) technical program plan. Computers & Security, 7(5):515– 516, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902465.

Grimm:1988:BRD

[Gri88a] Vanessa Jo Grimm. Bill to regulate data matching moves to floor. Computers & Security, 7(5):521, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902799.

Grimm:1988:DEH

[Gri88b] Vanessa Jo Grimm. Data errors hurt State Department system. Computers & Security, 7(1):105–106, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905421.

Guarro:1987:PPL

[Gua87] Sergio B. Guarro. Principles and procedures of the LRAM approach to information systems risk analysis and management. Computers & Security, 6(6):493–504, December 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900307.

Guynes:1983:EIC

[Guy83] Steve Guynes. EFTS impact on computer security. Computers & Security, 2(1):73–77, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900378.

Hartson:1983:TPC

[Har83] H. Rex Hartson. Teaching protection in computing: a researchoriented graduate course. Computers & Security, 2(3):248–255, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404883900081.

Harper:1986:ICL

[Har86] Robert M. Harper. Internal control in local area networks: an accountant’s perspective. Computers & Security, 5(1):28–35, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488690115X.

Hargrove:1988:SIA

[Har88] Robert Hargrove. Secure information is accurate information. Computers & Security, 7(1):104, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905329.

Harrold:1989:ISA

[Har89] C. L. Harrold. An introduction to the SMITE approach to secure computing. Computers & Security, 8(6):495–505, October 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889900795.

Huang:1988:SWP

[HC88] Yue Jiang Huang and Fred Cohen. Some weak points of one fast cryptographic checksum algorithm and its improvement. Computers & Security, 7(5):503–505, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902040.

Hoving:1985:MFM

[HE85] Per L. Hoving and Johan Essén. Minutes from the first meeting of TC 11, security and protection in information processing systems. Computers & Security, 4(2):149–152, June 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900203.

Henry:1982:FDA

[Hen82] Paul S. Henry. Fast decryption algorithm for the knapsack cipher. Computers & Security, 1(1):80–83, January 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900293.

Herschberg:1987:HC

[Her87] I. S. Herschberg. The hackers’ comfort. Computers & Security, 6(2):133–138, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900848.

Herschberg:1988:MTH

[Her88] I. S. Herschberg. Make the tigers hunt for you. Computers & Security, 7(2):197–203, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903367.

Highland:1986:ERRb

[HF86] Harold Joseph Highland and Viiveke Fåk. Electromagnetic radiation revisited, part II. Computers & Security, 5(3):181–184, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886900027.

Highland:1982:GNC

[HH82] Esther Harris Highland and Harold Joseph Highland. A guide to NBS computer security literature. Computers & Security, 1(2): 164–176, June 1982. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404882900098.

Highland:1982:SSS

[Hig82a] Esther H. Highland. System security standard for electronic data processing: Rolf Moulton New York: the city of New York, 1981. $7.50. Computers & Security, 1(2):192, June 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900141.

Highland:1982:BRMb

[Hig82b] Harold Joseph Highland. Book review: Management guide to computer security: J. R. Talbot, New York: John Wiley and Sons, 1981. xi + 180 pp. $29.95. Computers & Security, 1(2):191–192, June 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488290013X.

Highland:1982:BRMa

[Hig82c] Harold Joseph Highland. Book review: Managing information security: a program for the electronic information age: James A. Schweitzer Boston: Butterworth (Publishers), Inc., 1982. vi + 133 pp. $15.95. Computers & Security, 1(2):191, June 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900128.

Highland:1982:Ea

[Hig82d] Harold Joseph Highland. Editorial. Computers & Security, 1(1): 1–2, January 1982. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404882900189.

Highland:1982:Eb

[Hig82e] Harold Joseph Highland. Editorial. Computers & Security, 1 (2):95, June 1982. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404882900013.

Highland:1982:EN

[Hig82f] Harold Joseph Highland. Editor’s note. Computers & Security, 1(3):242, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404882900402.

Highland:1983:CST

[Hig83a] Esther H. Highland. Computer security techniques: List of controls by cost. Computers & Security, 2(2):189–193, June 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900585.

Highland:1983:BRA

[Hig83b] Harold Joseph Highland. Book review: Advances in computer system security: Rein Turn, editor: Dedham MA: Artech House, Inc., 1981. vi + 403 pp. (Q.E.D. Information Sciences, PO Box 181, Wellesley, MA 02181, USA) $29.50. Computers & Security, 2(1): 84, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404883900421.

Highland:1983:BRCa

[Hig83c] Harold Joseph Highland. Book review: CBEMA privacy and security bibliography (October 1982): Rein Turn, editor: Washington DC: Computer and Business Equipment Manufacturers Association, 1982, iii + 51 pp. Computers & Security, 2(1):83, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404883900408.

Highland:1983:BRCb

[Hig83d] Harold Joseph Highland. Book review: Codes, ciphers and computers: an introduction to information security: Bruce Bosworth: Rochelle Park NJ: Hayden Book Company, Inc., 1982. viii + 259 pp., $13.95. Computers & Security, 2(1):83–84, January 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488390041X.

Highland:1983:BRH

[Hig83e] Harold Joseph Highland. Book review: The hut six story: Breaking the Enigma codes: Gordon Welchman: London: Allen Lane (Penguin Books, Ltd) and New York: McGraw Hill Book Co., Inc., 1982. x + 326 pp., $12.95. Computers & Security, 2(1):82– 83, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404883900391.

Highland:1983:BRT

[Hig83f] Harold Joseph Highland. Book review: Tutorial: the security of data in networks: Donald W. Davies: Silver Spring MD: IEEE Computer Society, 1981. vi + 241 pp., $20.00 to non-members and $15.00 to members of IEEE. Computers & Security, 2(1): 84, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404883900433.

Highland:1983:Ea

[Hig83g] Harold Joseph Highland. Editorial. Computers & Security, 2(1): 1–2, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404883900263.

Highland:1983:Eb

[Hig83h] Harold Joseph Highland. Editorial. Computers & Security, 2(3): 213, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404883900019.

Highland:1983:IMT

[Hig83i] Harold Joseph Highland. Impact of microcomputers on total computer security. Computers & Security, 2(2):171–183, June 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488390055X.

Highland:1984:Ea

[Hig84a] Harold Joseph Highland. Editorial. Computers & Security, 3(2): 77–79, May 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900488.

Highland:1984:Eb

[Hig84b] Harold Joseph Highland. Editorial. Computers & Security, 3(3): 169–170, August 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900385.

Highland:1984:Ec

[Hig84c] Harold Joseph Highland. Editorial. Computers & Security, 3(4): 257, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900026.

Highland:1985:Ea

[Hig85a] Harold Joseph Highland. Editorial. Computers & Security, 4(1): 1–2, March 1985. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404885900021.

Highland:1985:Eb

[Hig85b] Harold Joseph Highland. Editorial. Computers & Security, 4(2): 101–102, June 1985. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404885900136.

Highland:1985:Ec

[Hig85c] Harold Joseph Highland. Editorial. Computers & Security, 4(4):267–268, December 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900458.

Highland:1985:EOA

[Hig85d] Harold Joseph Highland. Editorial: Office automation or on the value of being a dinosaur. Computers & Security, 4(3):185–186, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404885900264.

Highland:1985:FNC

[Hig85e] Harold Joseph Highland. The first national computer security conference. Computers & Security, 4(1):65–73, March 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900094.

Highland:1985:MSD

[Hig85f] Harold Joseph Highland. Microcomputer security: Data protection techniques. Computers & Security, 4(2):123–134, June 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900161.

Highland:1986:ERRa

[Hig86a] Harold Joseph Highland. Electromagnetic radiation revisited. Computers & Security, 5(2):85–93, June 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886901252.

Highland:1986:HMM

[Hig86b] Harold Joseph Highland. How to modify a microcomputer operating system for added security. Computers & Security, 5(4):340– 343, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886900568.

Highland:1986:ILa

[Hig86c] Harold Joseph Highland. In this issue . . .. Computers & Security, 5(1):1, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901021.

Highland:1986:ILb

[Hig86d] Harold Joseph Highland. In this issue . . .. Computers & Security, 5(2):83, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901239.

Highland:1986:ILc

[Hig86e] Harold Joseph Highland. In this issue . . .. Computers & Security, 5(3):179–180, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886900015.

Highland:1986:ILd

[Hig86f] Harold Joseph Highland. In this issue . . .. Computers & Security, 5(4):277–278, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886900416.

Highland:1987:CCS

[Hig87a] Harold Joseph Highland. The charade of computer security. Computers & Security, 6(2):108–109, April 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900794.

Highland:1987:CC

[Hig87b] Harold Joseph Highland. Cipher cracking. Computers & Security, 6(3):205, June 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887901015.

Highland:1987:CAM

[Hig87c] Harold Joseph Highland. Computer abuse and managerial response. Computers & Security, 6(6):448, December 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900204.

Highland:1987:CSN

[Hig87d] Harold Joseph Highland. Computer security: a new focus at the National Bureau of Standards. Computers & Security, 6(2):102– 103, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887900770.

Highland:1987:CSL

[Hig87e] Harold Joseph Highland. Computer security and law. Computers & Security, 6(3):203–205, June 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887901003.

Highland:1987:CVS

[Hig87f] Harold Joseph Highland. Computer viruses and sudden death! Computers & Security, 6(1):8–10, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887901131.

Highland:1987:DCS

[Hig87g] Harold Joseph Highland. Data cartridge for secure data. Computers & Security, 6(6):449–451, December 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900216.

Highland:1987:DES

[Hig87h] Harold Joseph Highland. Data encryption standard II? Computers & Security, 6(3):195–196, June 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900952.

Highland:1987:DPV

[Hig87i] Harold Joseph Highland. Data physician — a virus protection program. Computers & Security, 6(1):73–79, February 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887901283.

Highland:1987:RPI

[Hig87j] Harold Joseph Highland. The DES revisited — part II. Computers & Security, 6(2):100–101, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900769.

Highland:1987:EP

[Hig87k] Harold Joseph Highland. Encryption package. Computers & Security, 6(3):199–202, June 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900988.

Highland:1987:Ea

[Hig87l] Harold Joseph Highland. From the editor. Computers & Security, 6(3):193–194, June 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900940.

Highland:1987:Eb

[Hig87m] Harold Joseph Highland. From the editor. Computers & Security, 6(5):365–367, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900010.

Highland:1987:Ec

[Hig87n] Harold Joseph Highland. From the editor. Computers & Security, 6(6):445–446, December 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900186.

Highland:1987:EL

[Hig87o] Harold Joseph Highland. From the editor . . .. Computers & Security, 6(4):285–287, August 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900605.

Highland:1987:HDR

[Hig87p] Harold Joseph Highland. Hard disk recovery tools. Computers & Security, 6(6):458–462, December 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900253.

Highland:1987:HSY

[Hig87q] Harold Joseph Highland. How secure are your encryption keys? Computers & Security, 6(2):99–100, April 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900757.

Highland:1987:HEM

[Hig87r] Harold Joseph Highland. How to evaluate microcomputer encryption software and hardware. Computers & Security, 6(3):229–244, June 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887901040.

Highland:1987:ILa

[Hig87s] Harold Joseph Highland. In this issue . . .. Computers & Security, 6(1):1–3, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887901118.

Highland:1987:ILb

[Hig87t] Harold Joseph Highland. In this issue . . .. Computers & Security, 6 (2):97–98, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887900745.

Highland:1987:IS

[Hig87u] Harold Joseph Highland. Information sources. Computers & Security, 6(3):198–199, June 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900976.

Highland:1987:NMI

[Hig87v] Harold Joseph Highland. New member of the international board of Editors. Computers & Security, 6(6):447, December 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900198.

Highland:1987:OAS

[Hig87w] Harold Joseph Highland. Office automation security. Computers & Security, 6(6):452–457, December 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488790023X.

Highland:1987:PR

[Hig87x] Harold Joseph Highland. Passwords revisited. Computers & Security, 6(6):451–452, December 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900228.

Highland:1987:RN

[Hig87y] Harold Joseph Highland. Random notes. Computers & Security, 6(6):463, December 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900265.

Highland:1987:RNR

[Hig87z] Harold Joseph Highland. Random notes and reports. Computers & Security, 6(2):109–110, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900800.

Highland:1987:RB

[Hig87-27] Harold Joseph Highland. Report from Beijing. Computers & Security, 6(3):196–198, June 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900964.

Highland:1987:SFT

[Hig87-28] Harold Joseph Highland. Secure fax transmission. Computers & Security, 6(5):368–370, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900022.

Highland:1987:TW

[Hig87-29] Harold Joseph Highland. Technology watch. Computers & Security, 6(2):103–108, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900782.

Highland:1987:TLC

[Hig87-30] Harold Joseph Highland. Tempest over leaking computers. Computers & Security, 6(6):457–458, December 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900241.

Highland:1987:UDC

[Hig87-31] Harold Joseph Highland. US data controls. Computers & Security, 6(3):202–203, June 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488790099X.

Highland:1987:WYR

[Hig87-32] Harold Joseph Highland. Would you report a suspected data crime to the authorities? Computers & Security, 6(4):288–295, August 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900617.

Highland:1988:BCV

[Hig88a] Dr. Harold Joseph Highland. Blitzed by a computer virus. Computers & Security, 7(3):237–239, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900223.

Highland:1988:HTV

[Hig88b] Dr. Harold Joseph Highland. How to test a virus filter without a virus. Computers & Security, 7(3):235–237, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890020X.

Highland:1988:OSH

[Hig88c] Dr. Harold Joseph Highland. Open Sesame — hackers welcome. Computers & Security, 7(3):232–233, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900181.

Highland:1988:RFN

[Hig88d] Dr. Harold Joseph Highland. Report on the fourth national computer security conference: COMPSEC IV. Computers & Security, 7(3):309–311, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900399.

Highland:1988:SDD

[Hig88e] Dr. Harold Joseph Highland. Security director’s dilemma. Computers & Security, 7(3):233–235, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900193.

Highland:1988:SVP

[Hig88f] Dr. Harold Joseph Highland. Special virus protection measures. Computers & Security, 7(3):239–242, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900235.

Highland:1988:SWS

[Hig88g] Dr. Harold Joseph Highland. Star wars security or where did those micros go? Computers & Security, 7(3):243–244, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900259.

Highland:1988:UCN

[Hig88h] Dr. Harold Joseph Highland. Upcoming conferences of note. Computers & Security, 7(3):242–243, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900247.

Highland:1988:VDR

[Hig88i] Dr. Harold Joseph Highland. Virus disaster red book. Computers & Security, 7(3):237, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900211.

Highland:1988:ARAb

[Hig88j] Esther H. Highland. Abstracts of recent articles and literature. Computers & Security, 7(6):609–615, December 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900120.

Highland:1988:AFA

[Hig88k] Esther H. Highland. Agencies fail to appreciate threat to data security: Belden Menkus, Government Computer New, April 29, 1988, p. 36. Computers & Security, 7(4):430, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906347.

Highland:1988:ARF

[Hig88l] Esther H. Highland. Agencies ready forces for data security campaign: Robert V. Head, Government Computer News, May 27, 1988, pp. 27, 31. Computers & Security, 7(4):424, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888905913.

Highland:1988:AMSb

[Hig88m] Esther H. Highland. Apple–Microsoft suit raises issue of vague copyright laws: Robert Francis, Datamation, May 15, 1988, pp. 22, 26. Computers & Security, 7(4):423, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905895.

Highland:1988:AFR

[Hig88n] Esther H. Highland. Apple files response to Microsoft’s countersuit: Laurie flynn InfoWorld, May 2, 1988, p. 5. Computers & Security, 7 (4):423, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488890586X.

Highland:1988:ARAa

[Hig88o] Esther H. Highland. Apple releases “anti-viral” program: Nick Arnett InfoWorld, May 2, 1988, p. 8. Computers & Security, 7(4): 426, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906086.

Highland:1988:ASM

[Hig88p] Esther H. Highland. Apple slaps Microsoft, HP with copyright suit. Computers & Security, 7(3):321–322, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900521.

Highland:1988:AMSa

[Hig88q] Esther H. Highland. Apple/Microsoft suit split in two: Rachel Parker, InfoWorld, May 23, 1988, p. 1. Computers & Security, 7 (4):423, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888905883.

Highland:1988:ALU

[Hig88r] Esther H. Highland. Are Apple and Lotus unfair? Computers & Security, 7(3):322, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900557.

Highland:1988:BCL

[Hig88s] Esther H. Highland. Brazil copyright law could thwart pirates. Computers & Security, 7(3):321, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900478.

Highland:1988:CSG

[Hig88t] Esther H. Highland. Calamities spark growing field. Computers & Security, 7(3):326–327, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900855.

Highland:1988:CWT

[Hig88u] Esther H. Highland. Can we trust computers?: Andrew M. Lister, Computer Control Quarterly, Spring 1988, pp. 31–35. Computers & Security, 7(4):425, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888906001.

Highland:1988:CBP

[Hig88v] Esther H. Highland. Computer backup pools, disaster recovery, and default risk. Computers & Security, 7(3):324, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900727.

Highland:1988:CCM

[Hig88w] Esther H. Highland. Computer crime: MIS confronts a cancer. Computers & Security, 7(3):327, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900909.

Highland:1988:CCS

[Hig88x] Esther H. Highland. Computer cross sections unite at law conference. Computers & Security, 7(3):321, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900491.

Highland:1988:CIK

[Hig88y] Esther H. Highland. Computer insecurity: Ken McLeod, Information Age, April 1988, pp. 89–93. Computers & Security, 7(4): 424, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888905962.

Highland:1988:CIG

[Hig88z] Esther H. Highland. Computer insurance: the gaps. Computers & Security, 7(3):322–323, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900594.

Highland:1988:CSF

[Hig88-27] Esther H. Highland. Computer security: the federal approach. Computers & Security, 7(3):320–321, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900454.

Highland:1988:CSN

[Hig88-28] Esther H. Highland. Computer systems need isolated ground that is safe and noise free: Thomas M. Gruzs, Computer Technology Review, Spring 1988, p. 103 ff. Computers & Security, 7(4):432– 433, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906487.

Highland:1988:CSU

[Hig88-29] Esther H. Highland. Computer systems under siege. Computers & Security, 7(3):325, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900752.

Highland:1988:CVC

[Hig88-30] Esther H. Highland. Computer viruses can infect entire organizations: Angel L. Rivera, Government Computer News, April 29, 1988, pp. 37, 43. Computers & Security, 7(4):426, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888906098.

Highland:1988:CVS

[Hig88-31] Esther H. Highland. Computer viruses spread. Computers & Security, 7(3):325, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900764.

Highland:1988:CW

[Hig88-32] Esther H. Highland. Computer weekly. Computers & Security, 7 (3):323, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888900661.

Highland:1988:DCS

[Hig88-33] Esther H. Highland. DEC computer security tactics come into focus. Computers & Security, 7(3):328, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890096X.

Highland:1988:DDS

[Hig88-34] Esther H. Highland. Decentralized DP sites seen lacking in recovery planning. Computers & Security, 7(3):328, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900971.

Highland:1988:DRM

[Hig88-35] Esther H. Highland. Despite regs, most agencies lack disaster plan: Tari Schreider, Government Computer News, April 29, 1988, p. 54. Computers & Security, 7(4):431, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906402.

Highland:1988:DRP

[Hig88-36] Esther H. Highland. Disaster recovery planning. Computers & Security, 7(3):323, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900624.

Highland:1988:DSF

[Hig88-37] Esther H. Highland. Disturbance in several forms. Computers & Security, 7(3):328, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900958.

Highland:1988:DFC

[Hig88-38] Esther H. Highland. DOD fights off computer virus. Computers & Security, 7(3):325–326, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900788.

Highland:1988:ELS

[Hig88-39] Esther H. Highland. Exec liable in software copy case, court rules. Computers & Security, 7(3):322, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900569.

Highland:1988:FTI

[Hig88-40] Esther H. Highland. Fault tolerance increasingly important for LANS. Computers & Security, 7(3):327, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900879.

Highland:1988:FOP

[Hig88-41] Esther H. Highland. Federal officials puzzled by computer virus attacks: Neil Munro and Richard A. Dance, Government Computer News, April 29, 1988, p. 89. Computers & Security, 7(4): 426, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906074.

Highland:1988:FKC

[Hig88-42] Esther H. Highland. Fire KOs Chicago networks: Jean S. Bozman, Computerworld, May 16, 1988, pp. 1, 121. Computers & Security, 7 (4):432, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906451.

Highland:1988:HPMb

[Hig88-43] Esther H. Highland. H-P, Microsoft respond to Apple copyright suit. Computers & Security, 7(3):322, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900545.

Highland:1988:HLR

[Hig88-44] Esther H. Highland. Hashing out legal rights. Computers & Security, 7(3):321, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488890051X.

Highland:1988:HBW

[Hig88-45] Esther H. Highland. House Bill would increase 1989 funding for NBS. Computers & Security, 7(3):322, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900570.

Highland:1988:HME

[Hig88-46] Esther H. Highland. How much is enough? Expert says security efforts should pay, not cost: William Murray, Computerworld Focus, April 6, 1988, p. 30 ff. Computers & Security, 7(4):428, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888906232.

Highland:1988:HSY

[Hig88-47] Esther H. Highland. How safe is your data: Joseph m. pujals information center, April 1988, p. 46. Computers & Security, 7(4): 428, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906219.

Highland:1988:HVP

[Hig88-48] Esther H. Highland. How vaccine programs work: Neil J. Rothenberg, PC Magazine, June 28, 1988, p. 35. Computers & Security, 7 (4):427, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906165.

Highland:1988:HWN

[Hig88-49] Esther H. Highland. HR 145: What’s next. Computers & Security, 7(3):320, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888900442.

Highland:1988:HBC

[Hig88-50] Esther H. Highland. The hype behind computer viruses: Their bark may be worse than their “byte”: Paul Karon PC Week, May 31, 1988, pp. 49, 53. Computers & Security, 7(4):427, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488890613X.

Highland:1988:IPS

[Hig88-51] Esther H. Highland. Information property: Some intellectual property aspects of the global information economy. Computers & Security, 7(3):322, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900582.

Highland:1988:ITD

[Hig88-52] Esther H. Highland. Information technology and dataveillance: Roger A. Clarke, Communications of the ACM, May 1988, pp. 498–512. Computers & Security, 7(4):433, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906499.

Highland:1988:IHD

[Hig88-53] Esther H. Highland. Insurance and the hacker: David Davies, Computer Control Quarterly, Spring 1988, pp. 28–30. Computers & Security, 7(4):425, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905998.

Highland:1988:IDS

[Hig88-54] Esther H. Highland. An introduction to database security. Computers & Security, 7(3):329, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888901046.

Highland:1988:ICH

[Hig88-55] Esther H. Highland. It can’t happen here: Patricia Keefe, Computerworld Focus, April 6, 1988, pp. 13–16. Computers & Security, 7 (4):428, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906244.

Highland:1988:CMP

[Hig88-56] Esther H. Highland. It’s critical to make PC security as flexible as PCs themselves. Computers & Security, 7(3):329–330, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488890106X.

Highland:1988:LS

[Hig88-57] Esther H. Highland. LAN security. Computers & Security, 7 (3):324, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888900697.

Highland:1988:LSP

[Hig88-58] Esther H. Highland. LAN security problem detected on Navy ships: Brad Bass, Government Computer News, April 1, 1988, pp. 1, 84. Computers & Security, 7(4):430, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906323.

Highland:1988:LPS

[Hig88-59] Esther H. Highland. Legal protection of software in the United Kingdom. Computers & Security, 7(3):323, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900636.

Highland:1988:LAL

[Hig88-60] Esther H. Highland. Long arm of the law. Computers & Security, 7(3):321, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488890048X.

Highland:1988:LFD

[Hig88-61] Esther H. Highland. The “look and feel” debate. Computers & Security, 7(3):322, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900533.

Highland:1988:MVD

[Hig88-62] Esther H. Highland. Mac virus discovered at EPA, vaccination called successful: Richard A. Danca, Government Computer News, May 13, 1988, p. 4. Computers & Security, 7(4):426, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888906104.

Highland:1988:MVSa

[Hig88-63] Esther H. Highland. Mac virus surfaces in Washington: Nick Arnett InfoWorld, April 11, 1988, pp. 1, 3. Computers & Security, 7 (4):425, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906025.

Highland:1988:MVSb

[Hig88-64] Esther H. Highland. Macintosh virus seeks EDS signatures: Nick Arnett InfoWorld, April 25, 1988, p. 8. Computers & Security, 7 (4):426, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906062.

Highland:1988:MPP

[Hig88-65] Esther H. Highland. Make a plan for PC disaster recovery. Computers & Security, 7(3):329, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888901009.

Highland:1988:MOO

[Hig88-66] Esther H. Highland. MBS, OPM officials plan broad security guidelines: Kevin Power, Government Computer News, May 13, 1988, pp. 1, 92. Computers & Security, 7(4):423–424, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905901.

Highland:1988:MRO

[Hig88-67] Esther H. Highland. MCSC regs outline basic security level for PCs. Computers & Security, 7(3):330, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888901071.

Highland:1988:MSR

[Hig88-68] Esther H. Highland. Media security: Ralph M. Smole, Computer Control Quarterly, Spring 1988, pp. 36–37. Computers & Security, 7(4):425, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906013.

Highland:1988:MSSb

[Hig88-69] Esther H. Highland. Micro security system effective even on laptops: Peter Stevenson, Government Computer News, April 29, 1988, p. 48. Computers & Security, 7(4):431, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906384.

Highland:1988:MSC

[Hig88-70] Esther H. Highland. Microcomputer security and control. Computers & Security, 7(3):326, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900806.

Highland:1988:MSSa

[Hig88-71] Esther H. Highland. Microsoft seeks to split Apple suit into two phases: Rachel Parker, InfoWorld, May 23, 1988, p. 5. Computers & Security, 7(4):423, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905871.

Highland:1988:MPW

[Hig88-72] Esther H. Highland. Monitoring points out weak spots: Gardner Terrell, Government Computer News, April 29, 1988, pp. 1, 88. Computers & Security, 7(4):429, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906268.

Highland:1988:NBE

[Hig88-73] Esther H. Highland. A nasty bug: Editorial Computerworld, April 4, 1988, p. 18. Computers & Security, 7(4):425, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906049.

Highland:1988:NEG

[Hig88-74] Esther H. Highland. New emphasis given to UNIX security problems. Computers & Security, 7(3):326, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900818.

Highland:1988:NSL

[Hig88-75] Esther H. Highland. New security law compels federal agencies to act. Computers & Security, 7(3):321, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900466.

Highland:1988:NRH

[Hig88-76] Esther H. Highland. NSA regs hinder inexpensive network security: Barbara Bates, Government Computer News, April 29, 1988, pp. 50, 54. Computers & Security, 7(4):431, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906396.

Highland:1988:NSG

[Hig88-77] Esther H. Highland. NSA suppresses GAO report. Computers & Security, 7(3):323–324, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900673.

Highland:1988:OAC

[Hig88-78] Esther H. Highland. OMB advocates control model for systems security: Kevin Power, Government Computer News, March 18, 1988, p. 21. Computers & Security, 7(4):427–428, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888906189.

Highland:1988:OSW

[Hig88-79] Esther H. Highland. Operating system is weak link in MVS security: Brian Currah, Government Computer News, April 29, 1988, p. 44. Computers & Security, 7(4):430–431, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906359.

Highland:1988:PAV

[Hig88-80] Esther H. Highland. Pentagon attacks viruses. Computers & Security, 7(3):323, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900648.

Highland:1988:PCN

[Hig88-81] Esther H. Highland. Personal but not confidential: a new debate over privacy. Computers & Security, 7(3):321, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900508.

Highland:1988:PEM

[Hig88-82] Esther H. Highland. Picking encryption method is no time for secrets: Harold Joseph Highland, Government Computer News, April 29, 1988, p. 39. Computers & Security, 7(4):431, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888906360.

Highland:1988:PCB

[Hig88-83] Esther H. Highland. Planning, consulting boost biometric bill. Computers & Security, 7(3):327, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900892.

Highland:1988:PQI

[Hig88-84] Esther H. Highland. Power quality issues smack micro market. Computers & Security, 7(3):328, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900946.

Highland:1988:PME

[Hig88-85] Esther H. Highland. Preventive measures to ensure PC data security. Computers & Security, 7(3):328–329, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900995.

Highland:1988:PWS

[Hig88-86] Esther H. Highland. Programmer workstation security. Computers & Security, 7(3):329, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888901058.

Highland:1988:RRM

[Hig88-87] Esther H. Highland. Recent rulings may give states copyright immunity. Computers & Security, 7(3):320, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900429.

Highland:1988:RBI

[Hig88-88] Esther H. Highland. Red book includes a metaphysical look at security. Computers & Security, 7(3):324–325, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900739.

Highland:1988:RRC

[Hig88-89] Esther H. Highland. A redundancy reducing cipher: Peter Wayner, Cryptologia, April 1988, pp. 107–112. Computers & Security, 7(4): 431–432, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906414.

Highland:1988:RPB

[Hig88-90] Esther H. Highland. Reliable power: Beyond the big box. Computers & Security, 7(3):329, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888901022.

Highland:1988:RBR

[Hig88-91] Esther H. Highland. Review: Breach reported in U.S. computers: John Markoff, The New York Times, April 18, 1988, pp. A1, A19. Computers & Security, 7(4):428–429, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906256.

Highland:1988:RCG

[Hig88-92] Esther H. Highland. Review: Confronting the growing threat of harmful computer software viruses: Jim Seymour and Jonathan Matzkin PC Magazine, June 28, 1988, pp. 33–35. Computers & Security, 7(4):427, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888906153.

Highland:1988:RDM

[Hig88-93] Esther H. Highland. Review: Dispute over Mac user interface longstanding: Stuart J. Johnston, InfoWorld, April 25, 1988, p. 33. Computers & Security, 7(4):423, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905858.

Highland:1988:RLM

[Hig88-94] Esther H. Highland. Review: Loose MIPS sink ships: W. Mark Goode, Government Data Systems, April 1988, pp. 12–13. Computers & Security, 7(4):430, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888906335.

Highland:1988:RLR

[Hig88-95] Esther H. Highland. Review: Lotus removes 1–2–3 protection: Douglas Barney, Computerworld, May 16, 1988, p. 5. Computers & Security, 7(4):432, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488890644X.

Highland:1988:RMC

[Hig88-96] Esther H. Highland. Review: Microsoft charges Apple with breach of contract: Stuart J. Johnston and Rachel Parker InfoWorld, April 11, 1988, pp. 1, 81. Computers & Security, 7(4):423, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888905846.

Highland:1988:RNV

[Hig88-97] Esther H. Highland. Review: New virus damaging Mac data, group says: Scott Mace InfoWorld, April 11, 1988, p. 29. Computers & Security, 7(4):425, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888906037.

Highland:1988:RPAb

[Hig88-98] Esther H. Highland. Review: Planning averts disaster at L.A. Bond trading firm: James A. Martin, Computerworld, May 9, 1988, p. 4. Computers & Security, 7(4):432, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906463.

Highland:1988:RPAa

[Hig88-99] Esther H. Highland. Review: Problems in auditing expert system development: Wayne J. Socha, Edpacs, March 1988, pp. 1–6. Computers & Security, 7(4):428, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888906190.

Highland:1988:RPP

[Hig88-100] Esther H. Highland. Review: Proper procedures are final security safeguard: Christine E. Howkins, Government Computer News, April 29, 1988, p. 40. Computers & Security, 7(4):431, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888906372.

Highland:1988:RRO

[Hig88-101] Esther H. Highland. Review: Reliance on optical fibres to increase telecommunication security: Alan H. Friedberg, Robert M. Harper and Michael J. Cerullo, Information Age, April 1988, pp. 73–78. Computers & Security, 7(4):424, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905937.

Highland:1988:RSC

[Hig88-102] Esther H. Highland. Review: Security concerns in a local area network environment: Michael I. Sobol, Telecommunications, March 1988, p. 96 ff. Computers & Security, 7(4):428, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906207.

Highland:1988:REG

[Hig88-103] Esther H. Highland. Review: The executives guide to workstation management and data security: Mark Goode Computer Control Quarterly, Spring 1988, pp. 10–17. Computers & Security, 7(4): 424–425, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888905986.

Highland:1988:RUU

[Hig88-104] Esther H. Highland. Review: UFS: User-friendly security: Alan R. Krull Computer Control Quarterly, Spring 1988, pp. 1–9. Computers & Security, 7(4):424, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905974.

Highland:1988:RUS

[Hig88-105] Esther H. Highland. Review: Unix system security issues: Michael J. Knox and Edward D. Bowden, Information Age, April, 1988, pp. 67–72. Computers & Security, 7(4):424, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905925.

Highland:1988:RTF

[Hig88-106] Esther H. Highland. Risk! The third factor: Paul Tate, Datamation, April 15, 1988, p. 58 ff. Computers & Security, 7(4): 429, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906293.

Highland:1988:SCC

[Hig88-107] Esther H. Highland. “Safe computing” catching. Computers & Security, 7(3):326, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900831.

Highland:1988:SCD

[Hig88-108] Esther H. Highland. Satellite communications for disaster recovery. Computers & Security, 7(3):326, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900843.

Highland:1988:SAC

[Hig88-109] Esther H. Highland. Secure access control and MAC-based signatures. Computers & Security, 7(3):323, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900600.

Highland:1988:SAS

[Hig88-110] Esther H. Highland. Security act is signed: Copyright bill backed. Computers & Security, 7(3):320, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900430.

Highland:1988:SSY

[Hig88-111] Esther H. Highland. Security and software: Are you fully insured? Computers & Security, 7(3):325, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900740.

Highland:1988:SEG

[Hig88-112] Esther H. Highland. Security expert gives advice on computer viruses: Tim McGraw, Government Computer News, April 15, 1988, p. 17. Computers & Security, 7(4):425–426, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888906050.

Highland:1988:SHB

[Hig88-113] Esther H. Highland. Security habits build healthy software: David J. Stang, Government Computer News, May 27, 1988, p.

53. Computers & Security, 7(4):426–427, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888906116.

Highland:1988:SOD

[Hig88-114] Esther H. Highland. Security, OSI and distributed systems: T. Knowles Information Age, April 1988, pp. 79–84. Computers & Security, 7(4):424, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905949.

Highland:1988:SPG

[Hig88-115] Esther H. Highland. Security problems greater than ever: Barbara Bates, Information WEEK, April 18, 1988, p. 48. Computers & Security, 7(4):429, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888906281.

Highland:1988:SSR

[Hig88-116] Esther H. Highland. Security systems rescue data from snoopers, user blunders. Computers & Security, 7(3):327, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900867.

Highland:1988:SAP

[Hig88-117] Esther H. Highland. A sensible approach to protecting your data. Computers & Security, 7(3):328, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900983.

Highland:1988:SCN

[Hig88-118] Esther H. Highland. Smart cards: a new tool for identification and access protection: Rola Krayem, Information Age, April 1988, pp. 85–88. Computers & Security, 7(4):424, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905950.

Highland:1988:SHV

[Hig88-119] Esther H. Highland. Software, hardware “vaccines” stem threat of deadly computer viruses. Computers & Security, 7(3):325, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888900776.

Highland:1988:SCP

[Hig88-120] Esther H. Highland. SPA cracks down on piracy, raids computer retailer: Nick Arnett InfoWorld, April 4, 1988, p. 6. Computers & Security, 7(4):428, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888906220.

Highland:1988:SSE

[Hig88-121] Esther H. Highland. Speakers say employee attitude is key to security: Kevin Power, Computerworld, May 13, 1988, p. 14. Computers & Security, 7(4):432, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888906426.

Highland:1988:SWH

[Hig88-122] Esther H. Highland. Stalking the wily hacker: Clifford Stoll, Communications of the ACM, May 1988, pp. 484–497. Computers & Security, 7(4):429, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488890627X.

Highland:1988:SDS

[Hig88-123] Esther H. Highland. Structing the data security program. Computers & Security, 7(3):327, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900910.

Highland:1988:TIE

[Hig88-124] Esther H. Highland. Taming the indoor elements: Nina Frank, Digital Review, May 16, 1988, pp. 91–93. Computers & Security, 7 (4):432, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906475.

Highland:1988:TL

[Hig88-125] Esther H. Highland. Tangible losses. Computers & Security, 7 (3):324, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888900715.

Highland:1988:TMS

[Hig88-126] Esther H. Highland. Technology makes security of voice and data against sabotage both effective and affordable: Paul Kirvan, Communications News, May 1988, pp. 82–83. Computers & Security, 7 (4):432, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906438.

Highland:1988:TM

[Hig88-127] Esther H. Highland. Tempest micros. Computers & Security, 7 (3):324, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888900685.

Highland:1988:TSC

[Hig88-128] Esther H. Highland. Top secret — concepts and implementation. Computers & Security, 7(3):329, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888901034.

Highland:1988:TSV

[Hig88-129] Esther H. Highland. Top secret, and vulnerable: John Markoff, The New York Times, April 25, 1988, pp. D1, D4. Computers & Security, 7(4):430, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888906311.

Highland:1988:TCT

[Hig88-130] Esther H. Highland. Tough countermeasures thwart hackers who attack vital data: Ram Pemmaraju, Computer Technology Review, April 1988, p. 8 ff. Computers & Security, 7(4):429–430, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488890630X.

Highland:1988:TSA

[Hig88-131] Esther H. Highland. Typical system access control problems and solutions. Computers & Security, 7(3):323, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900612.

Highland:1988:UNR

[Hig88-132] Esther H. Highland. Under new rule, auditors must search out fraud. Computers & Security, 7(3):326, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890082X.

Highland:1988:UST

[Hig88-133] Esther H. Highland. Unprotected symphony tunes up. Computers & Security, 7(3):324, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900703.

Highland:1988:VGB

[Hig88-134] Esther H. Highland. A virus gives business a chill. Computers & Security, 7(3):326, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488890079X.

Highland:1988:VPU

[Hig88-135] Esther H. Highland. Virus-plagued user warned of snake-oil pitch: Robert L. Scheirer, PC Week, May 10, 1988, p. 170. Computers & Security, 7(4):427, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888906141.

Highland:1988:VCK

[Hig88-136] Esther H. Highland. Virus that could kill. Computers & Security, 7(3):323, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488890065X.

Highland:1988:VIC

[Hig88-137] Esther H. Highland. Viruses infect corporate MIS: Alan J. Ryan, Computerworld, May 30, 1988, pp. 29, 31. Computers & Security, 7 (4):427, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888906128.

Highland:1988:VEV

[Hig88-138] Esther H. Highland. VMSecure: Enhancing VM security and control. Computers & Security, 7(3):327–328, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900922.

Highland:1988:VFR

[Hig88-139] Esther H. Highland. Voice, finger and retina scans: Can biometrics secure your shop? Computers & Security, 7(3):327, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900880.

Highland:1988:WAU

[Hig88-140] Esther H. Highland. What about users should know about PC security. Computers & Security, 7(3):328, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900934.

Highland:1988:WTS

[Hig88-141] Esther H. Highland. When trouble strikes, a UPS can save the data. Computers & Security, 7(3):329, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888901010.

Highland:1988:AD

[Hig88-142] Harold Joseph Highland. The anatomy of a DOS disk. Computers & Security, 7(4):357–360, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905743.

Highland:1988:AVA

[Hig88-143] Harold Joseph Highland. Anatomy of a virus attack. Computers & Security, 7(2):145–150, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903288.

Highland:1988:VAS

[Hig88-144] Harold Joseph Highland. Are virus attack statistics reliable? Computers & Security, 7(4):338, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905640.

Highland:1988:BVF

[Hig88-145] Harold Joseph Highland. The brain virus: Fact and fantasy. Computers & Security, 7(4):367–370, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905767.

Highland:1988:CHV

[Hig88-146] Harold Joseph Highland. Case history of a virus attack. Computers & Security, 7(1):3–5, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888904889.

Highland:1988:CCC

[Hig88-147] Harold Joseph Highland. Chernobyl computer communications. Computers & Security, 7(1):7–9, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888904919.

Highland:1988:CVM

[Hig88-148] Harold Joseph Highland. Computer viruses: Media hyperbole, errors and ignorance. Computers & Security, 7(5):442–450, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888901940.

Highland:1988:EEM

[Hig88-149] Harold Joseph Highland. Electromagnetic eavesdropping machines for Christmas? Computers & Security, 7(4):341, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905676.

Highland:1988:EI

[Hig88-150] Harold Joseph Highland. Electromagnetic interference. Computers & Security, 7(1):9–10, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888904920.

Highland:1988:EAE

[Hig88-151] Harold Joseph Highland. Encryption, attacks and ethics. Computers & Security, 7(1):5–6, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888904890.

Highland:1988:YI

[Hig88-152] Harold Joseph Highland. For your information. Computers & Security, 7(1):10–11, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888904944.

Highland:1988:Ea

[Hig88-153] Harold Joseph Highland. From the editor. Computers & Security, 7(2):114, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888903203.

Highland:1988:Eb

[Hig88-154] Harold Joseph Highland. From the editor. Computers & Security, 7 (4):334, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888905615.

Highland:1988:Ec

[Hig88-155] Harold Joseph Highland. From the editor. Computers & Security, 7 (5):438, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888901897.

Highland:1988:Ed

[Hig88-156] Harold Joseph Highland. From the editor. Computers & Security, 7(6):532, December 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900016.

Highland:1988:HR

[Hig88-157] Harold Joseph Highland. Hackers revisited. Computers & Security, 7(4):339, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888905652.

Highland:1988:HSF

[Hig88-158] Harold Joseph Highland. How secure are fiber optics communications? Computers & Security, 7(1):25–26, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890497X.

Highland:1988:HCC

[Hig88-159] Harold Joseph Highland. How to combat a computer virus. Computers & Security, 7(2):157, 161–163, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890332X.

Highland:1988:HOC

[Hig88-160] Harold Joseph Highland. How to obtain a computer virus. Computers & Security, 7(4):337, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905639.

Highland:1988:HPMa

[Hig88-161] Harold Joseph Highland. How to protect a microcomputer system. Computers & Security, 7(1):6–7, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888904907.

Highland:1988:IS

[Hig88-162] Harold Joseph Highland. IFIP/Sec ’88. Computers & Security, 7 (1):10, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888904932.

Highland:1988:IRV

[Hig88-163] Harold Joseph Highland. Investigating a reported virus attack. Computers & Security, 7(4):344–345, August 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890569X.

Highland:1988:NVT

[Hig88-164] Harold Joseph Highland. New video terminal law. Computers & Security, 7(4):340, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905664.

Highland:1988:OVP

[Hig88-165] Harold Joseph Highland. An overview of 18 virus protection products. Computers & Security, 7(2):157–161, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903318.

Highland:1988:PR

[Hig88-166] Harold Joseph Highland. Product reviews. Computers & Security, 7(1):13–18, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888904956.

Highland:1988:PAO

[Hig88-167] Harold Joseph Highland. Program aids for office automation security. Computers & Security, 7(1):37–43, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905007.

Highland:1988:RBB

[Hig88-168] Harold Joseph Highland. Random bits and bytes. Computers & Security, 7(6):533–541, December 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900028.

Highland:1988:SIT

[Hig88-169] Harold Joseph Highland. Secretdisk II — transparent automatic encryption. Computers & Security, 7(1):27–34, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888904981.

Highland:1988:TT

[Hig88-170] Harold Joseph Highland. Tips on terrorism. Computers & Security, 7(4):344, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888905688.

Highland:1988:UCC

[Hig88-171] Harold Joseph Highland. U.S. Congress computer virus report. Computers & Security, 7(5):439, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888901903.

Highland:1988:VFE

[Hig88-172] Harold Joseph Highland. Virus filter evaluation. Computers & Security, 7(5):440, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888901927.

Highland:1988:VFUa

[Hig88-173] Harold Joseph Highland. Virus filter update. Computers & Security, 7(4):345, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905706.

Highland:1988:VFUb

[Hig88-174] Harold Joseph Highland. Virus filter update. Computers & Security, 7(5):441–442, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888901939.

Highland:1988:WPT

[Hig88-175] Harold Joseph Highland. Write-protect tab warning. Computers & Security, 7(5):439, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888901915.

Highlands:1988:E

[Hig88-176] Dr. Harold Joseph Highlands. From the editor. Computers & Security, 7(3):228, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900168.

Highland:1989:ARAa

[Hig89a] Esther H. Highland. Abstracts of recent articles and literature. Computers & Security, 8(1):74–82, February 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900448.

Highland:1989:ARAb

[Hig89b] Esther H. Highland. Abstracts of recent articles and literature. Computers & Security, 8(2):163–172, April 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900722.

Highland:1989:ARAc

[Hig89c] Esther H. Highland. Abstracts of recent articles and literature. Computers & Security, 8(3):259–269, May 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889901090.

Highland:1989:ARAd

[Hig89d] Esther H. Highland. Abstracts of recent articles and literature. Computers & Security, 8(4):353–361, June 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900977.

Highland:1989:ARAe

[Hig89e] Esther H. Highland. Abstracts of recent articles and literature. Computers & Security, 8(5):445–455, August 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900278.

Highland:1989:ARAf

[Hig89f] Esther H. Highland. Abstracts of recent articles and literature. Computers & Security, 8(6):537–546, October 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900837.

Highland:1989:DVN

[Hig89g] Harold Joseph Highland. Datacrime virus and new anti-virus products. Computers & Security, 8(8):659–661, December 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900035.

Highland:1989:Ea

[Hig89h] Harold Joseph Highland. From the editor. Computers & Security, 8(1):2, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889900308.

Highland:1989:Eb

[Hig89i] Harold Joseph Highland. From the editor. Computers & Security, 8(2):86, April 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488990059X.

Highland:1989:Ec

[Hig89j] Harold Joseph Highland. From the editor. Computers & Security, 8(3):176–177, May 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900990.

Highland:1989:Ed

[Hig89k] Harold Joseph Highland. From the editor. Computers & Security, 8(4):272–273, June 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900850.

Highland:1989:Ee

[Hig89l] Harold Joseph Highland. From the editor. Computers & Security, 8(5):366–367, August 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900163.

Highland:1989:Ef

[Hig89m] Harold Joseph Highland. From the editor. Computers & Security, 8(8):644–646, December 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900011.

Highland:1989:RBBa

[Hig89n] Harold Joseph Highland. Random bits and bytes. Computers & Security, 8(1):3–13, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488990031X.

Highland:1989:RBBb

[Hig89o] Harold Joseph Highland. Random bits and bytes. Computers & Security, 8(2):91–100, April 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900618.

Highland:1989:RBBc

[Hig89p] Harold Joseph Highland. Random bits and bytes. Computers & Security, 8(3):178–188, May 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889901004.

Highland:1989:RBBd

[Hig89q] Harold Joseph Highland. Random bits and bytes. Computers & Security, 8(4):274–282, June 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900862.

Highland:1989:RBBe

[Hig89r] Harold Joseph Highland. Random bits and bytes. Computers & Security, 8(5):369–380, August 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900175.

Highland:1989:RBBf

[Hig89s] Harold Joseph Highland. Random bits and bytes. Computers & Security, 8(6):460–478, October 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900758.

Highland:1989:RBBg

[Hig89t] Harold Joseph Highland. Random bits and bytes. Computers & Security, 8(7):553–561, November 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900473.

Highland:1989:RBBh

[Hig89u] Harold Joseph Highland. Random bits and bytes. Computers & Security, 8(8):647–657, December 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900023.

Highland:1989:RV

[Hig89v] Harold Joseph Highland. Reports from the victims. Computers & Security, 8(2):101, April 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488990062X.

Highland:1989:SDI

[Hig89w] Harold Joseph Highland. Secret disk II — administrator. Computers & Security, 8(7):563–568, November 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900485.

Highland:1989:VVS

[Hig89x] Harold Joseph Highland. VCHECKER — a virus search program. Computers & Security, 8(8):669–674, December 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900059.

Highlands:1989:E

[Hig89y] Harold Joseph Highlands. From the editor. Computers & Security, 8(6):458–459, October 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900746.

HJH:1985:BRC

[HJH85a] HJH. Book review: Communication regulation and international business: Juan F. Rada and G. Russell Pipe, editors Amsterdam: North-Holland, 1984. 266 + x pages, Dfl. 120.00/US $46.25. Computers & Security, 4(3):249, September 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900409.

HJH:1985:BRK

[HJH85b] HJH. Book review: Kahn on codes: Secrets of the new cryptology: David Kahn New York: Macmillan Publishing Company, 1984. 344 + viii pages, $19.95. Computers & Security, 4(3):247, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404885900379.

HJH:1985:BRM

[HJH85c] HJH. Book review: Managing risk: a computer-aided strategy: John M. Carroll Boston: Butterworths, 1984. 277 + xvi pages, $18.95. Computers & Security, 4(3):247–248, September 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900380.

HJH:1985:UCC

[HJH85d] HJH. U.S. computer crime legislation highlighted in special report. Computers & Security, 4(3):245–246, September 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900367.

Hoffman:1986:SVC

[HM86] Lance J. Hoffman and Lucy M. Moran. Societal vulnerability to computer system failures. Computers & Security, 5(3):211–217, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886900131.

Hoebeke:1986:CSP

[Hoe86] L. Hoebeke. Computer security — prevention: Lessons from the operation of a nuclear power plant. Computers & Security, 5(2): 122–127, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901343.

Hoffman:1989:RAC

[Hof89a] Lance J. Hoffman. Risk analysis and computer security: Towards a theory at last. Computers & Security, 8(1):23–24, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889900333.

Hoffman:1989:SBA

[Hof89b] Lance J. Hoffman. Smoking out the bad actors: Risk analysis in the age of the microcomputer. Computers & Security, 8(4): 325–344, June 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889900941.

Holt:1988:SEE

[Hol88] H. M. Holt. Software error experiment. Computers & Security, 7 (5):517, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902520.

Horwitt:1988:OSS

[Hor88] Elisabeth Horwitt. OSI security system revealed. Computers & Security, 7(1):105, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905378.

Hoving:1983:IAC

[Hov83] Per L. Hoving. To install an access control system, activities and checklists. Computers & Security, 2(2):163–170, June 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900548.

Hoving:1986:SST

[Hov86] Per Hoving. System security: the technical challenge. Computers & Security, 5(1):74–79, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901215.

Herschberg:1984:PTC

[HP84] I. S. Herschberg and R. Paans. The programmer’s threat: Cases and causes. Computers & Security, 3(4):263–272, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488490004X.

Hurford:1988:UUA

[HR88] J. M. Hurford and C. A. Ramsey. Understanding user attitudes: the cornerstone of computer documentation. Computers & Security, 7(5):516–517, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902519.

Hughes:1988:ARA

[Hug88] Pamela L. Hughes. Abstract of recent articles and literature: Radio frequency interfaces — myths, realities and a practical approach. Computers & Security, 7(2):214, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903471.

Hunter:1988:CTM

[Hun88a] Philip Hunter. Computer trouble makes British skies unfriendly. Computers & Security, 7(1):105, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890538X.

Hunter:1988:DRN

[Hun88b] Philip Hunter. Disaster recovery needed. Computers & Security, 7(1):104, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905305.

Hunter:1988:LSC

[Hun88c] Philip Hunter. In London, stocks up and computers down. Computers & Security, 7(1):107, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905536.

Jansen:1986:KSR

[Jan86] C. J. A. Jansen. On the key storage requirements for secure terminals. Computers & Security, 5(2):145–149, June 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886901380.

Jayant:1982:ASS

[Jay82] N. S. Jayant. Analog scramblers for speech privacy. Computers & Security, 1(3):275–289, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404882900475.

Jamieson:1989:SCI

[JL89] Rodger Jamieson and Graham Low. Security and control issues in local area network design. Computers & Security, 8(4):283–290, June 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889900874.

Jobusch:1989:SPMa

[JO89a] David L. Jobusch and Arthur E. Oldehoeft. A survey of password mechanisms: Weaknesses and potential improvements.Part 1. Computers & Security, 8(7):587–604, November 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900515.

Jobusch:1989:SPMb

[JO89b] David L. Jobusch and Arthur E. Oldehoeft. A survey of password mechanisms: Weaknesses and potential improvements.Part 2. Computers & Security, 8(8):675–689, December 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900060.

Johnstone:1987:KUF

[Joh87] Richard J. Johnstone. Key updating flags in EFT-POS security systems. Computers & Security, 6(3):252–260, June 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887901064.

Jones:1988:MHP

[Jon88] Stephen Jones. Microsoft, H-P shrug off Apple suit. Computers & Security, 7(5):521, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902805.

Jones:1988:SR

[JS88] Stephen Jones and Senior Editors. LA sites rocked. Computers & Security, 7(1):105, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905391.

Jung:1987:IRC

[Jun87] Achim Jung. Implementing the RSA cryptosystem. Computers & Security, 6(4):342–350, August 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900708.

Jung:1988:IRC

[Jun88] A. Jung. Implementing the RSA cryptosystem. Computers & Security, 7(5):510–511, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902131.

Kahn:1982:GLC

[Kah82] David Kahn. The grand lines of cryptology’s development. Computers & Security, 1(3):245–248, November 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900426.

Kaperonis:1984:IE

[Kap84] Ilias Kaperonis. Industrial espionage. Computers & Security, 3(2): 117–121, May 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900531.

Karger:1986:ADA

[Kar86] Paul A. Karger. Authentication and discretionary access control in computer networks. Computers & Security, 5(4):314–324, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404886900520.

Kinnon:1986:ASI

[KD86] A. Kinnon and R. H. Davis. Audit and security implications of electronic funds transfer. Computers & Security, 5(1):17–23, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404886901136.

Keenan:1989:EVO

[Kee89] Thomas P. Keenan. Emerging vulnerabilities in office automation security. Computers & Security, 8(3):223–227, May 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889901053.

Kingston:1986:SSB

[KG86] Peter P. C. H. Kingston and Martin Goulbourn. Security for small business computers. Computers & Security, 5(2):168, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404886901525.

Kingston:1987:SMU

[KG87] Peter P. C. H. Kingston and Martin Goulbourn. Security for microcomputers used in small business. Computers & Security, 6(1): 68–72, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887901271.

Kiefhaber:1988:NSW

[Kie88] A. Kiefhaber. National survey of worksite health promotion programs, 1985. Computers & Security, 7(5):517, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902532.

Kimball:1988:CCF

[Kim88] Mary C. Kimball. Central computer facility software directory. Computers & Security, 7(5):515, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902441.

Ko:1988:JMW

[Ko88] Dorothy Ko. In Japan, MIS weathers the storm easily. Computers & Security, 7(1):107, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905524.

Koch:1983:IIT

[Koc83] Harvey S. Koch. Implementing an integrated test facility audit using differential files. Computers & Security, 2(3):242–247, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488390007X.

Koch:1984:ALS

[Koc84] Harvey S. Koch. Auditing on-line systems: an evaluation of parallel versus continuous and intermittent simulation. Computers & Security, 3(1):9–19, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488490021X.

Kolodner:1988:RUV

[Kol88] E. K. Kolodner. Recovery using virtual memory. Computers & Security, 7(5):517, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902544.

Krause:1984:DEI

[Kra84] Lothar Krause. Data encryption in ISO, the international organization for standardization. Computers & Security, 3(3):234– 236, August 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900452.

Kruys:1989:SOS

[Kru89] Jan P. Kruys. Security of open systems. Computers & Security, 8(2):139–147, April 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900692.

Keefe:1989:SSO

[KTT89] T. F. Keefe, W. T. Tsai, and M. B. Thuraisingham. SODA: a secure object-oriented database system. Computers & Security, 8(6):517– 533, October 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889900813.

Kugler:1986:CPI

[Kug86] Han-Jürgen Kugler. Concepts of protection in the IFIP WG 2.7 reference model for command and response languages. Computers & Security, 5(2):168, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901471.

Kurzban:1982:SBA

[Kur82] Stanley A. Kurzban. A selective, briefly annotated bibliography of works on data security and privacy. Computers & Security, 1(1): 57–64, January 1982. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404882900268.

Lapid:1986:AHT

[LAN86] Yeheskel Lapid, Niv Ahituv, and Seev Neumann. Approaches to handling “Trojan Horse” threats. Computers & Security, 5(3):251– 256, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886900167.

Linden:1982:SEM

[LB82] Christer Lindén and Hans Block. Sealing electronic money in Sweden. Computers & Security, 1(3):226–230, November 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900372.

Lunt:1988:ELC

[LDS^{+}88] Teresa F. Lunt, Dorothy E. Denning, Roger R. Schell, Mark Heckman, and William R. Shockley. Element-level classification with a1 assurance. Computers & Security, 7(1):73–82, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905068.

Leahy:1986:PP

[Lea86] Patrick Leahy. Privacy and progress. Computers & Security, 5(4):347–349, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886900581.

Lewis:1988:CIB

[Lew88a] Brenda D. Lewis. Copyright issue blurs. Computers & Security, 7 (5):522, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902854.

Lewis:1988:WPP

[Lew88b] Peter H. Lewis. When the password is a passkey. Computers & Security, 7(1):103, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905287.

Louwerse:1984:DPA

[LK84] C. P. Louwerse and J. M. L. Kouwenberg. Data protection aspects in an integrated hospital information system. Computers & Security, 3(4):286–294, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900075.

Linn:1988:TOO

[LN88] R. J. Linn and J. S. Nightingale. Testing OSI (Open Systems Interconnection) protocols at the National Bureau of Standards. Computers & Security, 7(5):514, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902349.

Lobel:1983:SAC

[Lob83] Jerome Lobel. The state-of-the-art in computer security. Computers & Security, 2(3):218–222, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900032.

Longley:1987:ESA

[Lon87] Dennis Longley. Expert systems applied to the analysis of key management schemes. Computers & Security, 6(1):54–67, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488790126X.

Lee:1987:PAC

[LS87] J. A. N. Lee and Gerald Segal. Positive alternatives to computer misuse: a report of the proceedings of an ACM panel on hacking. Computers & Security, 6(2):177–179, April 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900915.

Lundberg:1988:CCA

[Lun88] B. G. Lundberg. CONST — a constructive approach to information modelling. Computers & Security, 7(5):514, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888902362.

Lunt:1989:ACP

[Lun89] Teresa F. Lunt. Access control policies: Some unanswered questions. Computers & Security, 8(1):43–54, February 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900382.

Madsen:1988:WMT

[Mad88] C. Wayne Madsen. The world meganetwork and terrorism. Computers & Security, 7(4):347–352, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890572X.

Malkoff:1988:RTF

[Mal88] D. B. Malkoff. Real-time fault detection and diagnosis: the use of learning expert systems to handle the timing of events. Computers & Security, 7(5):513–514, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902337.

Mueller-Boeling:1988:ISB

[MB88] D. Mueller-Boeling. Informationstechnik aus der Sicht der Benutzer im Wandel des letzten Jahrezehnts (Information technology from the point of the user during the last decade). Computers & Security, 7(5):516, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902507.

Moulton:1989:POP

[MB89] Rolf Moulton and Robert P. Bigelow. Protecting ownership of proprietary information. Computers & Security, 8(1):15–21, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889900321.

McDonald:1988:NDU

[McD88] Evelyn C. McDonald. North didn’t understand delete function. Computers & Security, 7(1):104, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905317.

Menkus:1983:NTD

[Men83] Belden Menkus. Notes on terrorism and data processing. Computers & Security, 2(1):11–15, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900299.

Menkus:1985:EAR

[Men85a] Belden Menkus. The EDP auditor’s role in computer security. Computers & Security, 4(2):135–138, June 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900173.

Menkus:1985:IBS

[Men85b] Belden Menkus. The impact of the Bell system breakup on DP disaster recovery planning. Computers & Security, 4(3):219–222, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404885900318.

Menkus:1988:B

[Men88a] Belden Menkus. Breakpoints. Computers & Security, 7(3):261– 268, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888900314.

Menkus:1988:GSD

[Men88b] Belden Menkus. Getting started in data processing disaster recovery. Computers & Security, 7(1):47–52, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905020.

Menkus:1988:UPC

[Men88c] Belden Menkus. Understanding password compromise. Computers & Security, 7(6):549–552, December 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900041.

Menkus:1988:UUP

[Men88d] Belden Menkus. Understanding the use of passwords. Computers & Security, 7(2):132–136, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888903252.

Menkus:1988:UGA

[Men88e] Belden Menkus. U.S. Government agencies belatedly address information system security issues. Computers & Security, 7(4):361– 366, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888905755.

Menkus:1989:CRF

[Men89a] Belden Menkus. Computer-related fire problems revisited. Computers & Security, 8(7):581–586, November 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900503.

Menkus:1989:CVS

[Men89b] Belden Menkus. The computer virus situation is not encouraging. Computers & Security, 8(2):115–119, April 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900667.

Menkus:1989:ERP

[Men89c] Belden Menkus. The employee’s role in protecting information assets. Computers & Security, 8(6):487–492, October 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900771.

Menkus:1989:TRD

[Men89d] Belden Menkus. It’s time to rethink data processing fire protection. Computers & Security, 8(5):389–394, August 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900199.

Menkus:1989:PSS

[Men89e] Belden Menkus. Physical security: Selecting an access control system. Computers & Security, 8(3):201–205, May 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889901028.

Miller:1985:TCS

[Mil85] James E. Miller. Transaction controls: Steps to take after the equipment has been secured. Computers & Security, 4(2):139–141, June 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404885900185.

Miles:1988:PPP

[Mil88a] J. B. Miles. Public-private partners work on secure protocols. Computers & Security, 7(1):108, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905585.

Miles:1988:USC

[Mil88b] J. B. Miles. Unclassified security calls for a ‘friendly’ approach. Computers & Security, 7(1):108, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905573.

Miller:1988:GPS

[Mil88c] D. F. Miller. Guidelines for preparing software user documentation. Computers & Security, 7(5):517, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902568.

Milor:1988:CSA

[Mil88d] G. Will Milor. The computer security act of 1987. Computers & Security, 7(3):251–253, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900284.

Moulton:1982:MAI

[Mou82] Rolf Moulton. A multidisciplinary approach to the investigation of computer abuse incidents. Computers & Security, 1(2):146–151, June 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404882900074.

Moulton:1984:DSM

[Mou84a] Rolf Moulton. Data security is a management responsibility. Computers & Security, 3(1):3–7, February 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900208.

Moulton:1984:PAS

[Mou84b] Rolf T. Moulton. A practical approach to system security devices. Computers & Security, 3(2):93–99, May 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900518.

Moulton:1987:HSC

[Mou87] Rolf T. Moulton. How to supervise and control I/S security officers and auditors. Computers & Security, 6(1):17–21, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887901210.

Mitchell:1987:CRK

[MP87] Chris Mitchell and Fred Piper. The cost of reducing key-storage requirements in secure networks. Computers & Security, 6(4):339– 341, August 1987. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404887900691.

Mitchell:1989:RHF

[MRW89] Chris Mitchell, Dave Rush, and Michael Walker. A remark on hash functions for message authentication. Computers & Security, 8(1): 55–58, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889900394.

Mollema:1989:PDC

[MS89] Kor IJ. Mollema and Ad P. M. Smulders. Prerequisites for data control. Computers & Security, 8(4):291–296, June 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900886.

McHugh:1988:MSI

[MT88] John McHugh and Bhavani M. Thuraisingham. Multilevel security issues in distributed database management systems. Computers & Security, 7(4):387–396, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905792.

Mullin:1984:NMP

[Mul84] Albert A. Mullin. A note on the mathematics of public-key cryptosystems. Computers & Security, 3(1):45–47, February 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900269.

Munro:1988:ETD

[Mun88a] Neil Munro. Expert tells DoD it needs training, not more tempest. Computers & Security, 7(1):108, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905561.

Munro:1988:NRM

[Mun88b] Neil Munro. New risk management lab will tackle security needs. Computers & Security, 7(1):105, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890541X.

Murray:1982:SPP

[Mur82] William Hugh Murray. Security procedures for program libraries. Computers & Security, 1(3):201–209, November 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900347.

Murray:1983:CSO

[Mur83] William H. Murray. Computer security: Observations on the state of the technology. Computers & Security, 2(1):16–23, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404883900305.

Murine:1986:PRS

[Mur86a] Gerald E. Murine. Producing reliable, secure software. Computers & Security, 5(2):169, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901549.

Murine:1986:SSI

[Mur86b] Gerald E. Murine. Secure software’s impact on reliability. Computers & Security, 5(4):336–339, December 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886900556.

Murray:1988:AEC

[Mur88] W. H. Murray. The application of epidemiology to computer viruses. Computers & Security, 7(2):139–145, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903276.

Mitchell:1988:SMS

[MW88] Chris Mitchell and Michael Walker. Solutions to the multidestination secure electronic mail problem. Computers & Security, 7(5):483–488, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902015.

Nicholson:1988:MDA

[NCCW88] W. L. Nicholson, P. J. Cowley, D. B. Carr, and M. A. Whiting. Managing the data analysis process. Computers & Security, 7(5): 516, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902477.

Neugent:1982:TCS

[Neu82] William Neugent. Teaching computer security: a course outline. Computers & Security, 1(2):152–163, June 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900086.

Nestman:1984:TTS

[NWH84] Chadwick H. Nestman, John C. Windsor, and Mary C. Hinson. Tutorial on telecommunications and security. Computers & Security, 3(3):215–224, August 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900427.

Oldehoeft:1984:SSU

[OM84] Arthur E. Oldehoeft and Robert McDonald. A software scheme for user-controlled file encryption. Computers & Security, 3(1):35– 41, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900245.

Osborne:1988:ABS

[OR88] W. M. Osborne and R. Raigrodski. Annotated bibliography on software maintenance. Computers & Security, 7(5):516, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888902489.

Owen:1986:CSI

[Owe86] Kenneth Owen. Computer security: IFIP addresses the practical issues. Computers & Security, 5(1):68–72, March 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886901197.

Paans:1984:CSE

[Paa84a] Ronald Paans. Computer security and EDP-auditing: the twin guards. Computers & Security, 3(4):259–261, November 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900038.

Paans:1984:EEC

[Paa84b] Ronald Paans. ECOMA-12: European Computer Measurement Association Conference: Special interest group on performance aspects of computer security. Computers & Security, 3(4):323– 324, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900117.

Paans:1987:E

[Paa87a] Ronald Paans. ECOMA-14. Computers & Security, 6(2):172, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900897.

Paans:1987:FEC

[Paa87b] Ronald Paans. First European conference on computer audit, control and security. Computers & Security, 6(2):173–176, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900903.

Paans:1987:ISI

[Paa87c] Ronald Paans. IFIP/Sec’86: Information security: the challenge. Computers & Security, 6(2):167–171, April 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900885.

Paans:1988:SEC

[Paa88] Dr. Ronald Paans. The second European conference on computer audit, control and security. Computers & Security, 7(3):312–315, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888900405.

Parker:1984:SSP

[Par84] Donn B. Parker. Safeguards selection principles. Computers & Security, 3(2):81–91, May 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900506.

Parker:1988:IPC

[Par88a] Rachel Parker. Industry praises copyright decision. Computers & Security, 7(5):521, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902787.

Parker:1988:SML

[Par88b] Rachel Parker. Scent of money lures firms to jump on antiviral bandwagon. Computers & Security, 7(5):523, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890288X.

Paans:1983:SSV

[PB83] Ronal Paans and Guus Bonnes. Surreptitious security violation in the MVS operating system. Computers & Security, 2(2):144–152, June 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404883900524.

Perry:1982:DCS

[Per82] William E. Perry. Developing a computer security and control strategy. Computers & Security, 1(1):17–26, January 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900219.

Perry:1988:SSA

[Per88] William E. Perry. Software still is arriving with those annoying bugs. Computers & Security, 7(1):108, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890555X.

Peercy:1988:RAM

[PES88] D. E. Peercy, M. D. Estill, and K. L. Shaw. Risk assessment methodology for software supportability (RAMSS): User’s handbook. Computers & Security, 7(5):510, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902118.

Pozzo:1987:ACC

[PG87] Maria M. Pozzo and Terence E. Gray. An approach to containing computer viruses. Computers & Security, 6(4):321–331, August 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900678.

Paans:1984:TSM

[PH84] R. Paans and I. S. Herschberg. A topology for secure MVS systems. Computers & Security, 3(4):278–285, November 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900063.

Paans:1986:HCM

[PH86] R. Paans and I. S. Herschberg. How to control MVS user supervisor calls. Computers & Security, 5(1):46–54, March 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886901173.

Paans:1987:CSL

[PH87] R. Paans and I. S. Herschberg. Computer security: the long road ahead. Computers & Security, 6(5):403–416, October 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900137.

Peercy:1988:SSR

[PHEW88] D. E. Peercy, W. Huebner, M. D. Estill, and J. C. Wu. Software supportability risk assessment in OT&E (operational test and evaluation): Historical baselines for risk profiles. volumes 1 and 2. Computers & Security, 7(5):510, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902106.

Pluimakers:1988:SNA

[Plu88] G. M. J. Pluimakers. Some notes on authorization and transaction management in distributed database systems. Computers & Security, 7(3):287–298, June 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888900351.

Polis:1984:ISR

[Pol84] Richard I. Polis. Information security: Reality and fiction. Computers & Security, 3(3):225–228, August 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900439.

Porter:1982:PEI

[Por82] Sigmund N. Porter. A password extension for improved human factors. Computers & Security, 1(1):54–56, January 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900256.

Porter:1984:CNS

[Por84] Sig Porter. Cryptology and number sequences: Pseudorandom, random, and perfectly random. Computers & Security, 3(1):43– 44, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900257.

Post:1987:IOS

[Pos87] Gerald V. Post. Improving operating system security. Computers & Security, 6(5):417–425, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900149.

Pfleeger:1988:TFA

[PP88] Charles P. Pfleeger and Shari Lawrence Pfleeger. A transaction flow approach to software security certification for document handling systems. Computers & Security, 7(5):495–502, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902039.

Pfleeger:1989:MPT

[PPT89] Charles P. Pfleeger, Shari Lawrence Pfleeger, and Mary Frances Theofanos. A methodology for penetration testing. Computers & Security, 8(7):613–620, November 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900540.

Pieprzyk:1985:DPK

[PR85a] Józef P. Pieprzyk and Dominik A. Rutkowski. Design of public key cryptosystems using idempotent elements. Computers & Security, 4(4):297–308, December 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900483.

Pieprzyk:1985:MDI

[PR85b] Józef P. Pieprzyk and Dominik A. Rutkowski. Modular design of information encipherment for computer systems. Computers & Security, 4(3):211–218, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900306.

Preston:1989:RV

[Pre89] Charles M. Preston. Review of VIRUSCAN. Computers & Security, 8(8):663–668, December 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900047.

Pluimakers:1986:ACS

[PvL86] G. M. J. Pluimakers and J. van Leeuwen. Authentication: a concise survey. Computers & Security, 5(3):243–250, September 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886900155.

Power:1986:AHE

[PW86] June M. Power and Steve R. Wilbur. Authentication in a heterogeneous environment. Computers & Security, 5(2):167, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404886901458.

Pfitzmann:1987:NUO

[PW87a] Andreas Pfitzmann and Michael Waidner. Networks without user observability. Computers & Security, 6(2):158–166, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900873.

Power:1987:AHE

[PW87b] June M. Power and Steve R. Wilbur. Authentication in a heterogeneous environment. Computers & Security, 6(1):41–48, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887901246.

Raab:1988:PCM

[Raa88] Selwyn Raab. Police chief of metro-North indicted in computer misuse. Computers & Security, 7(5):522, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902842.

Radai:1989:IPV

[Rad89] Yisrael Radai. The Israeli PC virus. Computers & Security, 8(2): 111–113, April 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889900655.

Ramaswamy:1989:PDI

[Ram89a] Raju Ramaswamy. Placement of data integrity security services in open systems interconnection architecture. Computers & Security, 8(6):507–516, October 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900801.

Ramaswamy:1989:SAD

[Ram89b] Raju Ramaswamy. Security architecture for data transfer through TCP/IP protocols. Computers & Security, 8(8):709–720, December 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889900096.

Reed:1982:RTY

[RB82] Susan K. Reed and Dennis K. Branstad. Reflections on ten years of computer security. Computers & Security, 1(3):231–232, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404882900384.

Requa:1988:WMC

[Req88] J. Requa. Workstation/mainframe connection. Computers & Security, 7(5):517, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902556.

RFB:1982:BRMa

[RFB82a] RFB. Book review: Management guide to computer security: J. P. Talbot, New York: John Wiley and Sons, 1981. xi + 180 pp. $29.95. Computers & Security, 1(3):290–291, November 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900487.

RFB:1982:BRMb

[RFB82b] RFB. Book review: Multi-national computer systems: an introduction to transnational data flow and data regulation: Harry S. Katzan, Jr.: New York: Van Nostrand Reinhold Co., 1980. xii + 198 pp. $16.95. Computers & Security, 1(3):291–292, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404882900499.

Rutledge:1986:SIC

[RH86] Linda S. Rutledge and Lance J. Hoffman. A survey of issues in computer network security. Computers & Security, 5(4):296– 308, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886900507.

Richards:1986:ICR

[Ric86] Roy Martin Richards. Insuring computer risks. Computers & Security, 5(3):207–210, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488690012X.

Richman:1988:APR

[Ric88] Sheldon L. Richman. Apple property rights suit can backfire. Computers & Security, 7(5):521–522, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902829.

Rihaczek:1987:TOC

[Rih87] Karl Rihaczek. TeleTrusT-OSIS and communication security. Computers & Security, 6(3):206–218, June 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887901027.

Riddle:1989:PUU

[RMS89] Bruce L. Riddle, Murray S. Miron, and Judith A. Semo. Passwords in use in a university timesharing environment. Computers & Security, 8(7):569–579, November 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900497.

Richards:1986:RMK

[RY86] Roy Martin Richards and Jan Yestingsmeier. Risk management — a key to security in the electronic funds transfer system. Computers & Security, 5(2):135–140, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901367.

Saari:1987:CCN

[Saa87] Juhani Saari. Computer crime — numbers lie. Computers & Security, 6(2):111–117, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900812.

Salomaa:1988:PKC

[Sal88] Arto Salomaa. A public-key cryptosystem based on language theory. Computers & Security, 7(1):83–87, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488890507X.

Sandhu:1988:NCA

[San88] Ravinderpal Singh Sandhu. Nested categories for access control. Computers & Security, 7(6):599–605, December 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900107.

Sorkin:1984:MCC

[SB84] Arthur Sorkin and C. James Buchanan. Measurement of cryptographic capability protection algorithms. Computers & Security, 3(2):101–116, May 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488490052X.

Schanning:1982:APK

[Sch82] Brian P. Schanning. Applying public key distribution to local area networks. Computers & Security, 1(3):268–274, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404882900463.

Schweitzer:1984:PWA

[Sch84] James A. Schweitzer. Personal workstation automation security vulnerabilities. Computers & Security, 3(1):21–28, February 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404884900221.

Schweitzer:1985:MVC

[Sch85] James A. Schweitzer. A management view: computer security as a discretionary decision. Computers & Security, 4(1):13–22, March 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404885900057.

Schweitzer:1987:HSF

[Sch87] James A. Schweitzer. How security fits in — a management view: Security is an essential for quality information. Computers & Security, 6(2):129–132, April 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900836.

Schweitzer:1988:SNB

[Sch88] Jim Schweitzer. Security news briefs. Computers & Security, 7(5):467–469, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888901976.

Seif:1984:CPB

[Sei84] Ruth A. Seif. Contingency planning in the banking community. Computers & Security, 3(1):29–34, February 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900233.

Serpell:1985:CES

[Ser85] Stephen C. Serpell. Cryptographic equipment security: a code of practice. Computers & Security, 4(1):47–64, March 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900082.

Shain:1989:SEF

[Sha89] M. Shain. Security in electronic funds transfer: Message integrity in money transfer and bond settlements through GE information services’ global network. Computers & Security, 8(3):209–221, May 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889901041.

Sherman:1988:CVV

[She88] A. T. Sherman. Cryptology and VLSI (Very Large Scale Integration). I. Detecting and exploiting algebraic weaknesses in cryptosystems. II. Algorithms for placing modules on a custom VLSI chip. Computers & Security, 7(5):512, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902234.

Shaw:1982:EGA

[SK82] James K. Shaw and Stuart W. Katzke. An executive guide to ADP contingency planning. Computers & Security, 1(3):210–215, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404882900359.

Summers:1988:PAK

[SK88] Rita C. Summers and Stanley A. Kurzban. Potential applications of knowledge-based methods to computer security. Computers & Security, 7(4):373–385, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905780.

Smith:1987:AUW

[Smi87] Sidney L. Smith. Authenticating users by word association. Computers & Security, 6(6):464–470, December 1987. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900277.

Snyder:1988:DQS

[Sny88] C. E. Snyder. Data quality: a strategic perspective. Computers & Security, 7(5):514, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902350.

Solarz:1987:CRE

[Sol87] Artur Solarz. Computer-related embezzlement. Computers & Security, 6(1):49–53, February 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887901258.

Sopko:1988:FHD

[Sop88] Sandra Sopko. Fire and humidity: Data safes protect against both. Computers & Security, 7(1):106, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905445.

Sorkin:1983:RST

[Sor83] Arthur Sorkin. Requirements for a secure terminal switch. Computers & Security, 2(3):268–274, November 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900111.

Spender:1987:ICU

[Spe87] J-C. Spender. Identifying computer users with authentication devices (tokens). Computers & Security, 6(5):385–395, October 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900113.

Stanley:1986:CCI

[Sta86a] Philip M. Stanley. Computer crime investigation and investigators. Computers & Security, 5(4):309–313, December 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886900519.

Stanley:1986:ECC

[Sta86b] Philip M. Stanley. Educating computer crime investigators. Computers & Security, 5(2):168–169, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886901537.

Stoll:1988:HSC

[Sto88] Cliff Stoll. How secure are computers in the U.S.A.?: an analysis of a series of attacks on Milnet computers. Computers & Security, 7(6):543–547, December 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488890003X.

Straub:1988:OSC

[Str88] Detmar W. Straub. Organizational structuring of the computer security function. Computers & Security, 7(2):185–195, April 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404888903355.

Tanca:1988:EES

[TC88] L. Tanca and S. Ceri. ELAND: an expert system for the configuration of local area networks applications. Computers & Security, 7 (5):514, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902374.

Tener:1986:DSS

[Ten86a] William T. Tener. Decision support systems — a necessity for effective auditing of controls. Computers & Security, 5(2):168, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404886901501.

Tener:1986:DCD

[Ten86b] William T. Tener. Detection of control deterioration using decision support systems. Computers & Security, 5(4):290–295, December 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404886900490.

Thuraisingham:1987:SCR

[Thu87] M. B. Thuraisingham. Security checking in relational database management systems augmented with inference engines. Computers & Security, 6(6):479–492, December 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404887900290.

Thuraisingham:1989:FVM

[Thu89] M. B. Thuraisingham. A functional view of multilevel databases. Computers & Security, 8(8):721–730, December 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900102.

Toppen:1984:ICA

[Top84] R. Toppen. Infinite confidence: ‘the audit of data communication networks’. Computers & Security, 3(4):303–313, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404884900099.

Tompkins:1986:ISAa

[TR86a] Frederick G. Tompkins and Russel Rice. Intergrating security activities into the software development life cycle and the software quality assurance process. Computers & Security, 5(2):167, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404886901446.

Tompkins:1986:ISAb

[TR86b] Frederick G. Tompkins and Russell S. Rice. Integrating security activities into the software development life cycle and the software quality assurance process. Computers & Security, 5(3):218– 242, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886900143.

Tonn:1988:AIR

[TTA^{+}88] B. E. Tonn, C. B. Travis, L. Arrowood, R. Goeltz, and C. A. Mann. Analysis of individual risk belief structures. Computers & Security, 7(5):509–510, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902088.

Thomsen:1989:PEM

[TTT89] D. J. Thomsen, W. T. Tsai, and M. B. Thuraisingham. Prototyping to explore MLS/DBMS design. Computers & Security, 8(3): 229–245, May 1989. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404889901065.

VanderBank:1988:CFM

[VA88] Dirk Van der Bank and Edwin Anderssen. Cryptographic figures of merit. Computers & Security, 7(3):299–303, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900363.

Varadharajan:1989:VNS

[Var89] Vijay Varadharajan. Verification of network security protocols. Computers & Security, 8(8):693–708, December 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900084.

vanBiene-Hershey:1988:TEA

[vBH88] Margaret E. van Biene-Hershey. The technical EDP-audit function. Computers & Security, 7(3):271–277, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900338.

vandenAssem:1986:CPA

[vdAvE86] R. van den Assem and W. J. van Elk. A chosen-plaintext attack on the Microsoft BASIC protection. Computers & Security, 5(1):

36–45, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901161.

vandenBerg:1984:ATC

[vdBL84] Bram van den Berg and Hans Leenaars. Advanced topics of a computer center audit. Computers & Security, 3(3):171–185, August 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404884900397.

vandeGoor:1984:ETD

[vdG84] A. J. van de Goor. Effects of technical developments on system security. Computers & Security, 3(4):314–322, November 1984. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404884900105.

vanderLans:1986:DSR

[vdL86] R. F. van der Lans. Data security in a relational database environment. Computers & Security, 5(2):128–134, June 1986. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404886901355.

vanEck:1985:ERV

[vE85] Wim van Eck. Electromagnetic radiation from video display units: an eavesdropping risk? Computers & Security, 4(4):269–286, December 1985. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/016740488590046X.

vanHoboken:1984:BV

[vH84] W. R. C. van Hoboken. The burglar’s viewpoint. Computers & Security, 3(4):295–302, November 1984. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404884900087.

vanHeurck:1987:TBS

[vH87] Ph. van Heurck. TRASEC: Belgian security system for electronic funds transfers. Computers & Security, 6(3):261–268, June 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887901076.

Villarreal:1988:IW

[Vil88] J. Villarreal. Innovative workstation. Computers & Security, 7(5): 512, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888902209.

Voydock:1985:SMT

[VK85] Victor L. Voydock and Stephen T. Kent. Security mechanisms in a transport layer protocol. Computers & Security, 4(4):325– 341, December 1985. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404885900513.

Voigt:1988:DRT

[Voi88] R. J. Voigt. Design of a real time operating system for a fault tolerant microcomputer. Computers & Security, 7(5):518, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/016740488890260X.

vanWyk:1989:LV

[vW89] Kenneth R. van Wyk. The Lehigh virus. Computers & Security, 8(2):107–110, April 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900643.

Warren:1983:ERC

[War83] Andrew D. Warren. Evaluating the risks of computer fraud and error. Computers & Security, 2(2):132–143, June 1983. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404883900512.

Ware:1986:EPIa

[War86a] Willis H. Ware. Emerging privacy issues. Computers & Security, 5(2):101–113, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488690132X.

Ware:1986:EPIb

[War86b] Willis H. Ware. Emerging privacy issues. Computers & Security, 5(2):167, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901410.

Witten:1988:PAA

[WC88] Ian H. Witten and John G. Cleary. On the privacy afforded by adaptive text compression. Computers & Security, 7(4):397– 408, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404888905809.

Worthington:1986:IDS

[WCWG86] T. K. Worthington, J. J. Chainer, J. D. Willford, and S. C. Gunderson. IBM dynamic signature verification. Computers & Security, 5(2):167–168, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/016740488690146X.

Weber:1988:EAP

[Web88] Ron Weber. Has EDP audit progressed in the last 15 years? Computers & Security, 7(1):103, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888905275.

Weber:1989:CEF

[Web89a] Ron Weber. Controls in electronic funds transfer systems: a survey and synthesis. Computers & Security, 8(2):123–137, April 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889900680.

Webster:1989:UDP

[Web89b] Anne E. Webster. University of Delaware and the Pakistani computer virus. Computers & Security, 8(2):103–105, April 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900631.

Weingarten:1983:CCP

[Wei83] F. Weingarten. Controlling cryptographic publication. Computers & Security, 2(1):41–48, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900330.

White:1989:VWC

[Whi89] Clinton E. White. Viruses and worms: a campus under attack. Computers & Security, 8(4):303, June 1989. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900916.

Willett:1982:CON

[Wil82a] Michael Willett. Cryptography old and new. Computers & Security, 1(2):177–186, June 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404882900104.

Willett:1982:GE

[Wil82b] Michael Willett. Guest editorial. Computers & Security, 1(3):243– 244, November 1982. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404882900414.

Willett:1982:TPK

[Wil82c] Michael Willett. A tutorial on public key cryptography. Computers & Security, 1(1):72–79, January 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404882900281.

Wiseman:1989:PVC

[Wis89] Simon Wiseman. Preventing viruses in computer systems. Computers & Security, 8(5):427–432, August 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404889900242.

Wong:1985:CCR

[Won85] Ken Wong. Computer crime — risk management and computer security. Computers & Security, 4(4):287–295, December 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404885900471.

Wong:1988:SOA

[Won88a] Ken Wong. Security and office automation. Computers & Security, 7(1):106–107, February 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888905482.

Wong:1988:MOS

[Won88b] W. Wong. Management overview of software reuse. Computers & Security, 7(5):516, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404888902490.

Wood:1982:FAC

[Woo82a] Charles Cresson Wood. Future applications of cryptography. Computers & Security, 1(1):65–71, January 1982. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488290027X.

Wood:1982:PDC

[Woo82b] Charles Cresson Wood. Policies for deterring computer abuse. Computers & Security, 1(2):139–145, June 1982. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404882900062.

Wood:1983:EISa

[Woo83a] Charles Cresson Wood. Effective information system security with password controls. Computers & Security, 2(1):5–10, January 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404883900287.

Wood:1983:EISb

[Woo83b] Charles Cresson Wood. Enhancing information security with the information resource management approach. Computers & Security, 2(3):223–229, November 1983. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404883900044.

Wood:1984:S

[Woo84] Charles Cresson Wood. Securicom ’84. Computers & Security, 3(2): 140–142, May 1984. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404884900579.

Wood:1985:BRS

[Woo85a] Charles Cresson Wood. Book review: Security for computer networks: D. W. Davies and W. L. Price New York: John Wiley and Sons, 1984. 386 + xix pages, $19.50. Computers & Security, 4(3):248–249, September 1985. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404885900392.

Wood:1985:FDS

[Woo85b] Charles Cresson Wood. Floppy diskette security measures. Computers & Security, 4(3):223–228, September 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488590032X.

Wood:1986:EIT

[Woo86a] Charles Cresson Wood. Establishing internal technical systems security standards. Computers & Security, 5(3):193–200, September 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404886900106.

Wood:1986:ETS

[Woo86b] Charles Cresson Wood. Establishing technical systems security standards at a large multinational bank. Computers & Security, 5 (2):169, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901550.

Wood:1987:HIS

[Woo87a] Charles Cresson Wood. The human immune system as an information systems security reference model. Computers & Security, 6(6):511–516, December 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404887900320.

Wood:1987:ISS

[Woo87b] Charles Cresson Wood. Information systems security: Management success factors. Computers & Security, 6(4):314–320, August 1987. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404887900666.

Wood:1988:CIS

[Woo88] Charles Cresson Wood. A context for information systems security planning. Computers & Security, 7(5):455–465, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888901964.

Wood:1989:PMA

[Woo89] Charles Cresson Wood. Planning: a means to achieve data communications security. Computers & Security, 8(3):189–199, May 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0167404889901016.

Wood:1986:SMP

[WZ86] Charles Cresson Wood and Howard M. Zeidler. Security modules: Potent information security system components. Computers & Security, 5(2):114–121, June 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404886901331.

Yeo:1985:IAC

[Yeo85] Gee Kin Yeo. Incorporating access control in forms systems. Computers & Security, 4(2):109–122, June 1985. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488590015X.

Zajac:1986:WDW

[Zaj86] Bernard P. Zajac. What to do when you have reason to believe your computer has been compromised. Computers & Security, 5(1): 11–16, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.com/
science/article/pii/0167404886901124.

Zajac:1988:DCL

[Zaj88a] Bernard P. Zajac. Dial-up communication lines: Can they be secured? Computers & Security, 7(1):35–36, February 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888904993.

Zajac:1988:DDP

[Zaj88b] Bernard P. Zajac. Distributed data processing: New problems. Computers & Security, 7(3):249–250, June 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888900272.

Zajac:1988:POH

[Zaj88c] Bernard P. Zajac. Personnel: the other half of data security. Computers & Security, 7(2):131–132, April 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888903240.

Zajac:1988:VSW

[Zaj88d] Bernard P. Zajac. Viruses: Should we quit talking about them? Computers & Security, 7(5):471–472, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888901988.

Zajac:1989:DRY

[Zaj89a] Bernard P. Zajac. Disaster recovery — are you really ready? Computers & Security, 8(4):317–323, June 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
016740488990093X.

Zajac:1989:LOC

[Zaj89b] Bernard P. Zajac. Legal options to computer viruses. Computers & Security, 8(1):25–27, February 1989. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0167404889900345.

Zaritsky:1988:MSI

[Zar88] I. S. Zaritsky. Microcomputer security issues in an office environment. Computers & Security, 7(5):509, October 1988. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/pii/
0167404888902076.